Syllabus of Subjects

Graduate School of Medical Sciences The University of Tokushima

Table of Contents

Chapter 1 Master’s Degree 1

Master Course of Medical Sciences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Chapter 2 Doctor’s Degree 13

Course of Medicine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Course of Proteomics Medical Sciences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

Chapter 1

Master’s Degree

Graduate School of Medical Sciences (2011) ⟩ Master’s Degree ⟩ Master Course of Medical Sciences

Master Course of Medical Sciences Syllabus of Subjects

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ....3 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) .................. 4

Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka

・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) ........................................4

Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・

Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester) ...............................................5 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao

・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ....................................................... 5 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・

Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga ......................................6

• 各専攻系間共通カリキュラム科目

Human Science (Basic human science from structure to function) … Rokutan ・ Kawai ・ Masuda / each-year(1st

semester) ........................................................................................... 6 Introduction to Clinical Pharmacology … Tamaki ・ Takiguchi ・ Kiwada ・ Tsuchiya ・ Yamauchi ・ Tomita / each-year(2nd

semester) ........................................................................................... 7 Introduction to Proteomics … Taniguchi ・ Kido ・ Ebina ・ Noma ・ Miyamoto / each-year(1st semester) .............7 Genomic Drug Discovery … Itou ・ Itakura ・ Tamaki ・ Sano ・ Part-time Lecturer / each-year(1st semester) .......... 8

• 専門科目

Human Science(Pathology) … Sakashita ・ Izumi ・ Kuwahara ・ Adachi ・ Yasutomo / 1st-year(1st semester, 2nd semester)8 General Medicine … Kaji ・ Ito ・ Soeki ・ Doi ・ Bando ・ Nakajo ・ Kawano ・ Takeda ・ Saito .........................9 Summary of Medical Genomic … Ebina ・ Itakura ・ Fukui / each-year(1st semester)............................9 Basic practice for medical science … Matsumoto ・ Fukui ・ Kido ・ Yano ・ Ishimura ・ Fukui ・ Sano ・ Izumi ・ Sasaki ・ Ebina

・ Yasutomo ・ Tamaki ・ Morita ・ Yoshizaki ・ Taniguchi ・ Itakura ・ Matsumoto ・ Takahama ・ Adachi / 1st-year(1st semester)9 Proteomics in Medicine … Siomi ・ Sakaguchi / 1st-year(1st semester) .......................................10

Applied Molecular Enzymology and Pathogenesis … Kido ・ Yano / 1st-year(1st semester)..................10

Medical Electronics … Yoshizaki ・ Sei ・ Moriguchi ・ Kinouchi / each-year(1st semester) .........................10 Medical Application of Nanotechnology … Doi ・ Miyoshi ・ Ishida ・ Ishikawa / 1st-year(1st semester) .......... 11 Clinical Genetics … Nakahori ・ Sato / 1st-year(1st semester) ................................................11

Development, Differentiation, and Regeneration … Takahama ・ Matsumoto ・ Noji ・上野・ Hibi / each-year(1st

semester) .......................................................................................... 11 Project Search Study Course … Fukui ・ Sone ・ Ishimura ・ Ebina ・ Yoshizaki ・ Itakura ・ Tamaki ・ Siomi ・ Fukui ・ Kido ・

Matsumoto ・ Taniguchi ・ Sasaki ・ Takahama ・ Sano ・ Izumi ・ Yasutomo ・ Adachi ・ Nakahori ・ Kubo ・ Moriguchi ・ Matsumoto ・ Doi ・ Nagahiro ・ Morita ・ Yano ・ Matsumoto ・ Tokunaga ・ Kagami ・ Rokutan ・ Ohmori ・ Takeda ・ Sei / 1st-year(whole year), 2nd-year(whole year) ................................................................................12

メディカルサイエンス特別研究 … 修士課程全教員 / 2nd-year .............................................12

Introduction to Biological Ethics 2 units (compulsory) each-year(1st semester)

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

, Seiichiro Kitamura · Professor / Institute of Health Biosciences

Junji Terao · Professor / Institute of Health Biosciences , Nobuyoshi Nakajo · Professor

, Kazuo Minakuchi · Professor / Institute of Health Biosciences

Target⟩ バイオサイエンスおよび医療に従事する者は，人権，生命倫理に十分な配慮を行い，個人情報保護，実験動物愛護にも同じく目を向けなければならない．本授業は生命倫理に関わる基本的知識を修得することを目的としている．

Outline⟩ 生命倫理学，臨床倫理学，社会倫理，個人情報保護，実験動物愛護などの問題に日頃接することの多

い講師が，経験に基づいた講義をオムニバス方式で行う． Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 「ヒトゲノム · 遺伝子解析に関する倫理指針」について板倉光夫
2. 「実験動物管理と情報処理にまつわる倫理学」太田房雄
3. 「キャンパスハラスメントを通して「医療従事者の倫理」を考える」 北村清一郎
4. 「法規や制度に関する基本的考え方—法医学の立場から」久保真一
5. 「大学における人権問題」佐野壽昭
6. 「組織検体取り扱い上の倫理的問題」〃
7. 「食品の機能性 · 安全性の評価と社会倫理」寺尾純二
8. 「尊厳死や安楽死など生命倫理について」中條信義
9. 「遺伝医学の視点」中堀豊
10. 「先天異常， (or 神経疾患，家族性腫瘍) の遺伝カウンセリング」〃
11. 「動物実験倫理」松本耕三
12. 「臨床治験に関する倫理」水口和生
13. 「ヒト胚の倫理学的な地位」山野修司
14. 「臓器移植の倫理的問題 (仮題)」井藤久雄 · 予定
15. (未定)

Evaluation Criteria⟩ 出席状況 (厳格にする予定) を踏まえ，さらに小テストないし面接試験で評価することを

検討している． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218234 Contact⟩ ⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp) ⇒ Kitamura (口腔顎顔面形態学分野教授室, +81-88-633-7319, kitamura@dent.tokushima-u.ac.jp) (Office

Hour: 前後期:金曜日 12 時 ∼ 13 時) ⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp) ⇒ Nakajo (+81-88-633-7366, nakajo@dentclin.tokushima-u.ac.jp) ⇒ Minakuchi (+81-88-633-7212, minakuti@clin.med.tokushima-u.ac.jp)

Clinical Psychology 2 units (selection) 1st-year(1st semester)

Kenji Sato · Professor / Institute of Socio-Arts and Sciences

, Mayumi YAMAMOTO · Professor / Institute of Socio-Arts and Sciences

Motohiro Sakai · Associate Professor / Institute of Socio-Arts and Sciences , Takaki Fukumori · Associate Professor / Institute of Socio-Arts and Sciences

Masato Mitome · Professor / Institute of Health Biosciences

Target⟩ 臨床心理学の基礎的理論・技法および今日的課題の概説 Outline⟩ 心と身体は密接につながっている．したがって「心の問題」の理解と制御を扱う臨床心理学の基礎

の習得は，精神医学・心身医学のみならず，ヘルスバイオサイエンスを基盤とする医学，歯学，薬学，栄養学，保健学領域において重要である．そこで，本講義では，臨床心理学の定義，対象，方法 (代表的な心理検査，心理療法) について，初学者を考慮して，その基礎と今日的課題を概説する．心理療法に関しては，医学領域などでエビデンスを示す認知行動療法について詳述する．

Notice⟩ e-learning 対応． Goal⟩ 臨床心理学の基礎的理論・技法および今日的課題を説明できる Schedule⟩

大項目中項目
1. 臨床心理学の定義と対象 (佐藤)
2. 臨床心理学の測定方法 (1)  パーソナリティの査定法 (福森)
3. 心理療法・カウンセリングとは (福森)
4. 心理療法の基礎 (1)  精神分析 (福森)
5. 臨床心理学の測定方法 (2)  知能，発達の査定法 (山本)
6. 心理療法の基礎 (2)  分析心理学，クライエント中心療法 (山本)
7. 発達障害と特別支援教育の現状と課題 (山本)
8. 行動理論と行動療法 (境)
9. 認知療法うつ病 (境)
10. 認知行動療法不安障害 (境)
11. 臨床心理的地域援助の基礎と実際ひきこもりに焦点を当てて (境)
12. 外傷後ストレス障害の認知行動療法 (佐藤)
13. 摂食障害の認知行動療法 (佐藤)
14. 体重減量・糖尿病の認知行動療法 (佐藤)

Evaluation Criteria⟩ 対面講義， e-learning とも受講とレポート提出を以て出席扱い．対面講義の場合，授業内の 15 分間で書く「出席・質問票・ミニレポート」の提出を以て， e-learning の場合，視聴記録 (受講) と MLS オンラインのレポート提出を以て出席とみなす．どちらのレポート課題とも「講義内容を A4 一枚でまとめる」．出席が三分の二以上の受講生が評価対象者

Re-evaluation⟩ 再試験無し Textbook⟩ 教科書は使用しない．参考書などは，適宜，授業時に紹介する Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218235 Contact⟩

⇒ Mitome (mitome@dent.tokushima-u.ac.jp) (Office Hour: 三留:水曜日の12:00∼ 13:00(e-mail により調整可能))

Introduction to social medicine, epidemiology and biostatistics 2 units (selection)

each-year(2nd semester)

Kokichi Arisawa · Professor / Course of Social and Environmental Medicine, Medical Science

, Hiro-O Ito · Professor / Institute of Health Biosciences

Hirokazu Uemura · Associate Professor / Course of Social and Environmental Medicine, Medical Science

, Aiko Yamauchi · Professor / Institute of Health Biosciences

Tetsuya Tanioka · Professor / Institute of Health Biosciences

, Akira Tokumura · Professor / Institute of Health Biosciences

Daisuke Hinode · Professor / Institute of Health Biosciences , Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Hiroaki Yanagawa · Associate Professor / University Hospital

Target⟩ 1, Clinical trial as a tool to establish clinical evidence 2. The role of the IT in the medical treatment and the hospital management are outlined. 3. Outline analytical skills for clarifying the phenomenon about Mental Health and Welfare. 4. Outline epidemiology and analytical methods for oral health promotion and/or oral disease prevention. 6. Outline damages from medicines and its etiology and prevention. 7. Outline designing of surveys and statistical analysis of data for writing a thesis.

Outline⟩ 1) Principles and practice of clinical trial from scientific and ethical viewpoint will be reviewed. 2) The usage of IT and the method of the management in the university hospital are outlined. 3) We will explain the statistical method for clarifying the feature of phenomenon based on previous work in Mental Health and Welfare. 4) We will explain the etiology of oral disease and its prevention based on the evidence of oral sciences. 6) Instruct background and etiology of several examples of damages from medicines and discuss possible prevention of the damages. 7) Basic aspects of statistical analysis, such as type of data and appropriate treatment of data, and meaning of hypothesis testing, will be summarized. In addition, practice of data analysis will be done using SPSS.

Notice⟩ It is possible to take a lesson by e-learning. Goal⟩ The goal is to deepen one’s understanding of contents presented during lectures, regarding social

medicine, pharmaceutical science and dentistry.

Schedule⟩

大項目担当

1.

Basic handling and statistical analysis of data for article makingHirokazu Uemura

2.  〃〃

3.

Practice of epidemiology and statistical data analysis for writing a thesis

Kokichi Arisawa

4.  〃〃

5.

Principles and practice of clinical trial Hiroaki Yanagawa

6.  〃〃

7.

The role of IT in a hospital Hiroki Moriguchi

8.  〃〃

9.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (1)

Shu-ichi Ueno

10.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (2)

Tetsuya Tanioka

11.

Epidemiology of dental caries and periodontal disease, and the relationship between oral disease and systemic disease

Hiro-o Ito

12.

〃

Daisuke Hinode

13. 医薬品の情報管理，医薬品の安全対策木原勝
14.  〃〃

15.

What is damages from medicines? Why did the damages occurred? How to prevent the damages?

Akira Tokumura

16.

What is the damages from medicines? Why did the damages occurred? How to prevent the damages?

〃 Evaluation Criteria⟩ Reports should be submitted on questions presented during lectures. Grades are

determined by summing points from all instructors. Re-evaluation⟩ In principle, there is no reexaminations. Textbook⟩ Teaching materials designated by each instructor wil be used.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218231 Contact⟩

⇒ Arisawa (+81-88-633-7071, arisawa@basic.med.tokushima-u.ac.jp) (Office Hour: Please contact eac

h instructor by an e-mail.)

Note⟩ Different classrooms may be used according instructors.

Introduction to How to Write Up Scientific Manuscript in English 2 units (selection)

each-year(2nd semester)

Shuhei Tomita · Professor / Course of Biological Medicine, Medical Science

, Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Yutaka Nakaya · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Hirofumi Nagashino · Professor / Institute of Health Biosciences

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Bukasa Kalubi · Assistant Professor / Medical Science

Target⟩ 21 世紀に医学，歯学，薬学，栄養学，保健学の各分野で活躍する人材は英語が堪能であることが要求される．本授業ではこれらの領域で用いられる独特の英語表現法に関わる基本的知識を修得することを目的とする．

Outline⟩ 研究成果を国際的な学術雑誌に発表したり，海外の学会やシンポジウムで発表や講演をしたりすることは，研究者養成を目的とする大学院教育において必須の履修目標である．そこで，本授業では医学英語論文，用紙のの作成方法について系統的な講義をビデオ並びにマルチメディア教材等を積極的に活用しながら行う．更に，医科学用英語の聴き取り及び英語による討論の訓練を行い，発表技術の向上を目指す．

Notice⟩ 1) 授業は後期のみに開講される．時間帯は原則的に火曜日の午後であるが，講師の事情等により変更されることがある． 2) 講義の一部は e-learning 化されているので， e-learning 学習も出席として取り扱う． 3)20 分以上の遅刻は出席と見なさない．

Schedule⟩

大項目担当

1.

生物医学雑誌投稿に要求される条件 (I)

Fukui Kiyoshi

2. 生物医学雑誌投稿に要求される条件について (II)  〃
3. 研究，論文，学会発表の進め方 (I)  市原明

4.

研究，論文，学会発表の進め方 (II)

Ichihara Akira

5. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (I) 長篠博文
6. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (II)  〃
7. 学会口演要旨作成上の留意点について実例をまじえて冨田修平
8. 論文作成に役立つコンピュータの利用法ー実例を示した結果，考察の書き方ー 中屋豊

9.

英語による PowerPoint presentation についての要点

梶龍兒

10. 英語論文の特性と日本語論文との比較林良夫
11. 口頭による英語発表と短報の書き方に関する基本を講義する I  岡崎拓
12.  〃〃
13. 研究の区切りとしての論文作成 I  福井裕行
14. II

Evaluation Criteria⟩ 講義への出席，随時の試験，受講態度を総合して行う． 2/3 以上の出席がなければ不合

格とする．再試験はない． Textbook⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218236

Psychosomatic health and environmental stress 2 units (selection)

Eiji Takeda · Professor / Institute of Health Biosciences

, Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Junji Terao · Professor / Institute of Health Biosciences

, Ken-ichi Miyamoto · Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Rintaro Yamanishi · Associate Professor / Institute of Health Biosciences

Hiroshi Sakaue · Associate Professor / Institute of Health Biosciences

, Toshio Hosaka · Assistant Professor / Institute of Health Biosciences

Target⟩ The aim of this program is to study on psychosomatic stress in daily life. Contents of this program

include gene expression profile under stress, stress response in health and diseases. Outline⟩ By the systems biology, clinical neurology, nutritional science techniques, the effects of stress on

development and growth in child will be assessed and discussed. �������⟩ psychosomatic health, environment, stress Goal⟩ To understand stress assessment technology

Schedule⟩

大項目担当

1.

Anxiety and depression (Tetsurou Ohmori) Noriaki Okazaki

2.

The development of the immune system (Yousuke Takahama)

〃

3.

Mechanism of sleep regulation (Hiroyoshi Sei)

Ryuji Kaji

4.

Endoplasmic reticulum stress in health and diseases (Seiichi Oyadomari)

Kazuhito Rokutan

5.

Stress genomics (Kazuhito Rokutan)

〃

6.

New concept and innovation of functional foods in Japan (Junji Terao)

Hiroyoshi Sei

7.

Why do we need ”stress control” through foods and food style? (Junji Terao)

〃

8.

Aging and mineral nutrition (ken-ichi Miyamoto)

Megumi Goto

9.

Caloric restriction and bone metabolism (Ken-ichi Miyamoto)

〃

10.

Breakfast and chewing for stress (Eiji Takeda) Hiroshi Chuman

11.

Food and stress related hormone (Eiji Takeda) Suehiro Sakaguchi

12.

Physical impact caused by stress (Yutaka Nakaya)

Eiji Takeda

13.

Gut immunity and stress (Yutaka Nakaya) Tetsurou Ohmori

14.

Oxidative stress and immunity (rintaro Yamanishi)

Gen Tamiya

15.

The role of adipokines on metabolic stress in obesity and obesity-related diseases (hiroshi Sakaue)

〃

16.

Diabetes mellitus and stress (Toshio Hosaka)

Eiji Takeda

Evaluation Criteria⟩ Reports on more than two third of the studies will be required for evaluation. Re-evaluation⟩ 再試験無し Textbook⟩ none Reference⟩ 授業時に適宜，紹介する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218232 Contact⟩

⇒ Takeda (+81-88-633-7093, takeda@nutr.med.tokushima-u.ac.jp) ⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour:

8:30-17:30) ⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp) ⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp) ⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp)

⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp) ⇒ Miyamoto (Molecular Nutrition, +81-88-633-7081, miyamoto@nutr.med.tokushima-u.ac.jp) ⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Yamanishi (+81-88-633-9366, rintaro@nutr.med.tokushima-u.ac.jp) ⇒ Sakaue (303, +81-88-633-7091, hsakaue@nutr.med.tokushima-u.ac.jp) ⇒ Hosaka .

Basic methods in life sciences 2 units (selection)

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Hiroyuki Mizuguchi · Associate Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Tatsuji Haneji · Professor / Institute of Health Biosciences

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Kazuko Seta · Part-time Lecturer

Kouji Itou · Professor / Institute of Health Biosciences

, Takahiro Matsumoto · Associate Professor / Institute of Health Biosciences

Keiko Kataoka · Associate Professor / Course of Microbiology and Immunology, Medical Science

, Tetsuo Yamazaki · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Takeshi Nikawa · Professor / Institute of Health Biosciences

Hironori Yamamoto · Assistant Professor / Institute of Health Biosciences

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics , Toshihiko Nagata · Professor / Institute of Health Biosciences Tetsuya Yoshinaga · Professor / Institute of Health Biosciences

Target⟩ 大学院に進学した直後の院生に生命科学に関する基礎的な実験方法を理解させる． Outline⟩ 実験動物の取り扱い方，細胞と器官の培養方法とその応用，蛋白質の取り扱い方とその解析方法，遺

伝子解析の方法とその応用，抗体を用いた研究方法とその蛋白質，免疫組織細胞化学に対する応用，数理モデルによる生体機能発現機構の解明等生命科学の基礎的な研究技法を講義する．

�������⟩ 生命科学, 実験手法 Notice⟩ 受講生は必ず出席すること． Goal⟩ 現在行われている生命科学研究方法の基礎を大学院進学直後に理解し，研究生活にスムーズに入れるよ

うにする． Schedule⟩

大項目

1. {大学院に入学直後に現在用いられている研究手法を集中的に講義する．第 1 回目は授業ガイダンスと生命科学の研究手法総 \atop 論 (羽地，研究部長)}
2. 実験動物を利用した研究 (松本)
3. 細胞と器官の培養 (羽地)
4. 細胞培養の応用 (宮本)
5. 蛋白研究 1(福井清)
6. 蛋白研究 2(藤原)
7. 抗体を用いた生化学的研究方法 (二川，山本)
8. 抗体を用いた研究の臨床応用 (中屋)
9. 免疫学研究の基礎 (安友)
10. 免疫学研究の最近の進歩 (安友)
11. 遺伝子解析 1(福井裕)
12. 遺伝子解析 2(伊藤)
13. 遺伝子解析 3(高濱)
14. 遺伝子解析 4(高濱)
15. 数理モデルによる生命機能の解析 (吉永)

Evaluation Criteria⟩ Web によるレポート，出席状況，受講態度等により総合的に判定する． Re-evaluation⟩ しない Textbook⟩ 指定しないが，講義の都度プリント等資料を配布する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218233 Contact⟩

⇒ Fukui (hfukui@ph.tokushima-u.ac.jp) ⇒ Mizuguchi (+81-88-633-7264, guchi003@ph.tokushima-u.ac.jp) ⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Haneji (+81-88-633-7321, tat-hane@dent.tokushima-u.ac.jp) ⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 8:30 to 17:30) ⇒ Seta . ⇒ Itou (Institute for Medicinal Resources, +81-88-633-7290, kitoh@ph.tokushima-u.ac.jp) (Office Hour:

Monday, 17:0019:00) ⇒ Matsumoto . ⇒ Kataoka (kataoka@basic.med.tokushima-u.ac.jp) ⇒ Yamazaki (医学臨床 B 棟 5 階医薬品病態生化学研究室, +81-88-633-7886, yamazakt@ph.tokushima-u

.ac.jp) (Office Hour: 特に定めない) ⇒ Okazaki (414, +81-88-633-9158, tokazaki@genome.tokushima-u.ac.jp) ⇒ Nikawa (nikawa@nutr.med.tokushima-u.ac.jp) ⇒ Yamamoto (+81-88-633-7235, yamamoto@nutr.med.tokushima-u.ac.jp) ⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp) ⇒ Nagata (+81-88-633-7343, nagata@dent.tokushima-u.ac.jp) ⇒ Yoshinaga (保健学 B 棟教員研究室 (吉永), +81-88-633-9050, yosinaga@medsci.tokushima-u.ac.jp)

(Office Hour: 18:00-19:00 every Friday)

Human Science (Basic human science from structure to function) 2 units (selection)

each-year(1st semester)

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Tomoko Kawai · Associate Professor Kiyoshi Masuda · Assistant Professor / Institute of Health Biosciences

Target⟩ 生体を構成する組織や器官の構造と働きを理解するため，基本的な生体物質と細胞についての講義を

し，細胞レベルから見たからだの働きについて理解させる． Outline⟩ 細胞の基本構造と機能，遺伝子からタンパク質を合成する仕組み，細胞の情報伝達，細胞の増殖 · 分

化 · 死の分子機構，受精，発生について講義を行い，メデイカルサイエンス研究を行うための基礎細胞生物学についての講義を行う．

Schedule⟩

大項目

1.

Introduction

2.

DNA, Chromosomes, and Genomes

3.

Replication, Repair, and Recombination

4.

From DNA to Protein

5.

Control of gene expression

6.

Membrane Structure

7.

Membrane Transport

8.

Mitochondria

9.

Intracellular Trafic

10.

Signal Transduction

11.

Cytoskeleton

12.

Cell Cycle and Apoptosis 13.

Cell Division

14. Genetics

15.

Cellualer Stress Response

16. Summary

Evaluation Criteria⟩ 出席状況，筆記試験，又はレポート等により評価する． Textbook⟩ エッセンシャル細胞生物学 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217008 Contact⟩

⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 18 :00 (e-mail により時間調節を適宜おこないます))

Introduction to Clinical Pharmacology 2 units (selection) each-year(2nd semester)

Toshiaki Tamaki⟨Manager⟩ · Professor / Course of Biological Medicine, Medical Science

, Yoshiharu Takiguchi · Professor / Institute of Health Biosciences

Hiroshi Kiwada · Professor / Institute of Health Biosciences

, Koichiro Tsuchiya · Professor / Institute of Health Biosciences

Aiko Yamauchi · Professor / Institute of Health Biosciences

, Shuhei Tomita · Associate Professor / Course of Biological Medicine, Medical Science

Target⟩ 薬の効果に個人差が生じる要因を理解し，有効で安全な医薬品の適正な使用方法に関する基本事項を

身につける． Outline⟩ 人々の病気の治療・予防に多大な役割を担っている薬物をより有効に，かつ安全に使うために，薬物

の人体における作用と生体内動態に影響を及ぼす諸因子について解説し，モデル式を利用した科学的評価に基づく薬物療法について講義する．また，具体的な疾患に対する薬物療法の現状と問題点や薬害についても解説する．

Notice⟩ 講義の一部は e-learning 化する予定である． e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 授業ガイダンス+臨床薬理学序論玉置俊晃
2. 薬物動態の原理冨田修平
3. 薬物の体内動態〃
4. 薬物のバイオアベイラビリティー際田弘志
5. TDM の実践滝口祥令
6. P450 と薬物相互作用冨田修平
7. 薬物相互作用玉置俊晃
8. 腹痛と治療薬〃
9. 高血圧治療薬〃
10. かぜ薬〃
11. 抗酸化薬土屋浩一郎
12. 薬物と健康食品玉置俊晃
13. 薬害〃
14. 妊婦・授乳婦と医薬品情報山内あい子
15. 臨床試験と生命倫理玉置俊晃

Evaluation Criteria⟩ 出席状況と小テストまたはレポート等により評価する． 12 回以上の出席が必要． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=222226 Contact⟩ ⇒ Tamaki (+81-88-633-7061, tamaki@basic.med.tokushima-u.ac.jp) (Office Hour: 面談希望教員に e-m

ail にて時間調節の上，面談して下さい． ) ⇒ Takiguchi (+81-88-633-7466, takiguti@ph.tokushima-u.ac.jp) (Office Hour: 面談希望教員に e-mail に

て時間調節の上，面談して下さい． ) ⇒ Kiwada (+81-88-633-7259, hkiwada@ph.tokushima-u.ac.jp) (Office Hour: 面談希望教員に e-mail に

て時間調節の上，面談して下さい． ) ⇒ Tsuchiya (Department of Medical Pharmacology, +81-88-633-7250, tsuchiya@ph.tokushima-u.ac.jp)

(Office Hour: 面談希望教員に e-mail にて時間調節の上，面談して下さい． ) ⇒ Yamauchi (+81-88-633-7266, aiko@ph.tokushima-u.ac.jp) (Office Hour: 面談希望教員に e-mail にて

時間調節の上，面談して下さい． ) ⇒ Tomita (207, +81-88-633-6071, tomita@basic.med.tokushima-u.ac.jp)

Introduction to Proteomics 2 units (compulsory) each-year(1st semester)

Hisaaki Taniguchi · Professor / Course of Medicine for Biological Responses, Proteomics

, Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics

Yousuke Ebina · Professor / Course of Medicine for Biological Responses, Proteomics , Takafumi Noma · Professor / Institute of Health Biosciences Ken-ichi Miyamoto · Professor / Institute of Health Biosciences

Target⟩ ポストゲノム時代における医学，生物学に重要な位置を占めるプロテオミクスの基礎を習得する． Outline⟩ 遺伝子産物である蛋白質がどのように相互作用して生命活動を制御するかを理解する手段としての

プロテオミクスの理論，解析法，応用例を体系的に講義する． Notice⟩ 講義においてはプリントも配布するが，プロジェクターなどを使用しての講義なのでノートの取り方

は工夫すること． Schedule⟩

大項目担当
1. 授業ガイダンス+プロテオミクス序論谷口寿章
2. プロテオミクスを用いたインスリンシグナル伝達の解明 1  蛯名洋介
3. プロテオミクスを用いたインスリンシグナル伝達の解明 2  〃
4. プロテオミクスを用いたインスリンシグナル伝達の解明 3  〃
5. 健康長寿食品とプロテオミクス 1  宮本賢一
6. 健康長寿食品とプロテオミクス 2  〃
7. エネルギー代謝のプロテオミクス 1  野間隆文
8. エネルギー代謝のプロテオミクス 2  〃
9. プロテオミクス解析の疾患への応用【基礎】木戸博
10. インフルエンザ脳症のプロテオミクス解析【応用 1】〃
11. ゲノミクスとプロテオミクスを統合したインフルエンザ脳症の解析【応用 2】 〃
12. プロテオミクスの解析技術谷口寿章
13. 生体高分子の質量分析法〃
14. 翻訳後修飾の解析〃
15. プロテオミクスの生命科学への応用〃

Evaluation Criteria⟩ 出席状況，受講態度，レポートなどにより判断する．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217566

Contact⟩

⇒ Taniguchi (+81-88-633-7426, hisatan@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 18:00 (e-

mail により時間調節を適宜おこないます)) ⇒ 他の教員についても, e-mail にて時間調節の上,面談して下さい.

Genomic Drug Discovery 2 units (selection) each-year(1st semester)

Kouji Itou⟨Manager⟩ · Professor / Institute of Health Biosciences

, Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics

Toshiaki Tamaki · Professor / Course of Biological Medicine, Medical Science , Shigeki Sano · Professor / Institute of Health Biosciences

Part-time Lecturer

Target⟩ ゲノム創薬の考え方と実際について学ぶ． Outline⟩ マイクロサテライトと単一ヌクレオチド多型を含ゲノム塩基配列に関する多型情報を用い，単一遺

伝子の原因遺伝子と多遺伝子疾患の疾患感受性遺伝子とその多型を明らかに出来る．またゲノム情報と疾患メタボロームとの相関，および疾患治療の分子標的の構造と機能に関する情報に基づく「ゲノム創薬」と創薬の歴史や創薬の実際について講義する．

Notice⟩ 講義の一部は e-learning 化を予定しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 授業ガイダンス+創薬の歴史玉置俊晃
2. ゲノム情報の現状板倉光夫
3. 一遺伝子疾患の原因遺伝子の探索板倉光夫
4. 疾患モデル動物を用いた疾患感受性遺伝子の探索板倉光夫
5. 患者を対象とする解析とゲノム創薬〃
6. グライコサイエンスの基礎と現状伊藤孝司
7. グライコバイオロジーと疾患〃
8. グライコサイエンスと創薬〃
9. 感染症克服に向けた創薬櫻井純
10. 有機化学を基盤とする創薬研究 (1)  佐野茂樹
11. 有機化学を基盤とする創薬研究 (2)  長尾善光
12. ポストゲノム時代の創薬研究 (1)  佐野茂樹
13. ポストゲノム時代の創薬研究 (2)  〃
14. 新薬の開発玉置俊晃
15. 臨床試験〃

Evaluation Criteria⟩ 出席により評価する． 12 回以上出席すること．欠席する場合は，該当講義日担当教員が指定する課題についてレポートを提出することにより出席と認定する．なお， 15 回出席したものは優， 13∼ 14 回出席したものは良， 12 回出席したものは可と判定する．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217071 Contact⟩

⇒ Itou (Institute for Medicinal Resources, +81-88-633-7290, kitoh@ph.tokushima-u.ac.jp) (Office Hour:

E-mailで質問等を受付け，必要があれば面談する． ) ⇒ Itakura (306, +81-88-633-9454, itakura@genome.tokushima-u.ac.jp) (Office Hour: E-mailで質問等を

受付け，必要があれば面談する． ) ⇒ Tamaki (+81-88-633-7061, tamaki@basic.med.tokushima-u.ac.jp) (Office Hour: E-mailで質問等を受

付け，必要があれば面談する． ) ⇒ Sano (薬学部本館 6 階東, +81-88-633-7273, ssano@ph.tokushima-u.ac.jp) (Office Hour: E-mailで質

問等を受付け，必要があれば面談する． )

Human Science(Pathology) 2 units (compulsory) 1st-year(1st semester, 2nd semester)

Naomi Sakashita · Professor / Institute of Health Biosciences

, Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science Tomomi Kuwahara · Associate Professor / Course of Microbiology and Immunology, Medical Science

Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science , Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science Target⟩ 循環障害，腫瘍，炎症，感染症，免疫異常など病気の原因と発生機序について，また，細菌，ウイル

ス，寄生虫など微生物学の基礎知識について，分子レベルで理解できるように教育するとともに，細胞レベル，臓器レベル，個体レベルの変化と関連させることを目的とした講義を行う．

Outline⟩ Prof. Yasutomo: Basic concepts of immune systems;effector immune cells and escape from immune

system Manner⟩ オムニバス形式 Notice⟩ 遅刻して受講した学生は，欠席とみなすことがある．講義は e-learning 化していない． Schedule⟩

大項目中項目担当
1. 昼間の講義 (学部学生への講義を利用する場合)
2. 病理学 1  細胞の病的変化佐野教授 (M3 の 60 分授業)
3.  〃循環障害〃
4.  〃細胞の分化と増殖の異常〃
5.  〃障害の修復〃
6.  〃炎症とその周辺〃
7. 病理学 2  病理解剖泉教授 (M3 の 60 分授業)
8.  〃がんの原因〃
9.  〃がんの種類と悪性度〃
10.  〃前がん性病変〃
11.  〃発がん機構〃
12. 細菌学細菌学の歴史{桑原助教授 (M2， M3 の 60 分 \atop 授業)}
13.  〃細菌の構造と分類〃
14.  〃細菌の増殖〃
15.  〃細菌の代謝〃
16. ウイルス学ウイルス学概論足立教授 (M2 の 60 分授業)
17.  〃ウイルスの構造と分類〃
18.  〃ウイルスの複製〃
19. 免疫学免疫学の基礎安友教授 (M2 の 60 分授業)
20.  〃抗体の多様性の獲得機構〃
21.  〃抗体の機能〃

Evaluation Criteria⟩ 出席率，講義の内容の理解度など．再試験は行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217563 Contact⟩

⇒ Sakashita . ⇒ Kuwahara (+81-88-633-9229, tomomi@basic.med.tokushima-u.ac.jp) (Office Hour: 月曜日の17:00∼ 1

8:00 (e-mailにより適宜時間調節を行う． )) ⇒ Yasutomo (+81-88-633-7048, yasutomo@basic.med.tokushima-u.ac.jp) (Office Hour: 月曜日の11: 00

∼ 12: 00 (e-mailにより適宜時間調節を行う． )) ⇒ 他の教員についても,e-mailにて時間調節の上,面談する.

Note⟩ 出席率が 60%に満たない学生は，受講しなかった講義を，次学年に追加して受けること．また，出席率が 60%に満たない場合はレポートとすることがある．

General Medicine 2 units (compulsory)

Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

, Susumu Ito · Professor

Takeshi Soeki · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science

, Toshio Doi · Professor / Course of Biological Medicine, Medical Science

Eiichi Bando · Professor , Nobuyoshi Nakajo · Professor

, Fumiaki Kawano · Professor / Institute of Health Biosciences

Eiji Takeda · Professor / Institute of Health Biosciences

, Ken Saito · Professor / Institute of Health Biosciences

Target⟩ 臨床医学の基礎知識について概説する． Outline⟩ 講義 Goal⟩ 循環器，呼吸器，消化器，腎臓，神経 · 筋，内分泌 · 代謝，血液の各臨床領域における代表的な疾病に

つき，発生機序および原因となる遺伝子などの異常，そして各々の疾患の病態生理を理解させ，最新の診断および治療法の理論と実践を学ばせる．

Schedule⟩

大項目中項目
1. 消化器疾患の診断と治療 (担当伊東進)
2. 循環器疾患の診断と治療 (担当添木武)
3. 臨床医科学概論・神経系総論 1 ・ 2 ・ 3(担当 梶龍兒)
4. 腎臓病 1 ・ 2(担当土井俊夫)
5. 顎口腔機能入門 1 ・ 2(担当坂東永一)
6. 麻酔薬と精神機能 1 ・ 2(担当中條信義)
7. 顎顔面領域の手術後のリハビリテーション 1 ・ 2(担当 河野文昭)
8. 医食同源 (担当武田英二)
9. 食の病理 (担当〃
10. 循環器疾患の病態解析と時間医学 (担当齋藤憲)

Evaluation Criteria⟩ 出席状況と小テストまたはレポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217553 Contact⟩

⇒ Kaji (rkaji@clin.med.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00-18:00 (e-mail により時間調節を適宜おこないます: rkaji@clin.med.tokushima-u.ac.jp)．他の教員についても， e-mail にて時間調節の上，面談して下さい． )

Summary of Medical Genomic 2 units (selection) each-year(1st semester)

Yousuke Ebina · Professor / Course of Medicine for Biological Responses, Proteomics , Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ ゲノムを通し，医科学を学ぶ Outline⟩ ゲノムの構造と機能を基盤とした生命化学および医学について教授する Notice⟩ e-learning は取り扱っていない

Schedule⟩

大項目担当
1. 授業ガイダンス+ゲノムの基本構造と機能 (1)  蛯名
2. ゲノムの基本構造と機能 (2)  〃
3. ゲノム解析の方法論 (1)  〃
4. ゲノム解析の方法論 (2)  〃
5. ゲノムと疾患〃

6. DNA と染色体の構造と機能板倉
7. 遺伝と家系〃
8. マッピング遺伝子地図の作製〃
9. ヒト疾患遺伝子の同定〃
10. 複合遺伝性疾患〃
11. 分子病の発見と遺伝子診断福井
12. DNA の構造と複製〃
13. 遺伝情報とその発現〃
14. 核酸・ヌクレオチドの生合成と代謝〃
15. ヒト統合失調症疾患感受性遺伝子〃

Evaluation Criteria⟩ 受講状況と授業の理解の程度に応じて評価する．再試験は行わない．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217558

Contact⟩

⇒ Ebina (+81-88-633-7436, iden@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 17:30(e-mailに

より時間調節を行います． )) ⇒ 他の教員についても， e-mailにて時間調節の上，面談してください．

Basic practice for medical science 2 units (selection) 1st-year(1st semester)

Mitsuru Matsumoto⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Kazunori Ishimura · Professor / Course of Integrated Brain Sciences, Medical Science

, Yoshihiro Fukui · Professor / Course of Human Development, Medical Science

Toshiaki Sano ·, Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Takuya Sasaki · Professor / Molecular Medicine, Proteomics

, Yousuke Ebina · Professor / Course of Medicine for Biological Responses, Proteomics

Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science

, Toshiaki Tamaki · Professor / Course of Biological Medicine, Medical Science

Kyoji Morita ·, Kazuo Yoshizaki · Professor / Course of Preventive Medicine, Medical Science

Hisaaki Taniguchi · Professor / Course of Medicine for Biological Responses, Proteomics

Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics

, Kozo Matsumoto ·

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ メディカルサイエンスを学ぶために必要な基本的知識と手技を分子細胞生物学，生理・生化学分野，

およびゲノム科学について，実際の実験を通して学ぶ． Outline⟩ 指導教官が学生の習得している専門知識・技能や修士論文テーマ，進路希望等を勘案して，以下の 3

研究分野からそれぞれ 1 人の教官の実習を指定する． Manner⟩ オムニバス方式 Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目内容担当
1. 分子細胞生物学分野分子生物学の基本的な知識と手技について実習を行う．福井清
2.  〃蛋白質の構造と機能の解析法を中心に，蛋白質化学の実習を行う．木戸・矢野
3.  〃細胞レベルでの遺伝子解析法，ゲノム操作法について実習を行う．塩見
4.  〃形態学的手法を中心とした医学研究について実習を行う．石村・福井義・佐野・泉
5. 生理・生化学分野生化学の基本的な知識と手技について実習を行う．佐々木
6.  〃シグナル伝達に関わる分子機構の解析を中心とした実習を行う．蛯名
7.  〃細胞の増殖機構に関わるメカニズムの解析方法について実習を行う．安友

8.  〃薬理学的作用の解析手法を中心とした実習を行う．玉置・森田
9.  〃生理学的手法を中心とした医学研究について実習を行う．吉崎
10.  〃細胞周期制御に関わる分子機構の解析を中心とした実習を行う．原
11. ゲノム科学分野プロテオーム解析法を中心に，蛋白質の機能解析法について実習を行う．谷口
12.  〃ゲノム構造の解析法，遺伝学的解析法について実習を行う．板倉
13.  〃モデル動物の解析を中心としたゲノム操作法について実習を行う．松本満・松本耕
14.  〃個体発生および組織分化のメカニズム解析のための実習を行う．高濱
15.  〃病原微生物のゲノム機能解析のための実習を行う．足立

Evaluation Criteria⟩ 受講状況と授業内容の理解の程度に応じて評価する．再試験は行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217577 Contact⟩

⇒ Matsumoto (Room 701, +81-88-633-7432, mitsuru@ier.tokushima-u.ac.jp) (Office Hour: 火曜日の16: 00∼ 18:00 (e-mail により時間調節を適宜，行ないます))

Proteomics in Medicine 2 units (selection) 1st-year(1st semester)

Haruhiko Siomi ·, Suehiro Sakaguchi · Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ タンパク質の多様性が生み出される分子機構と，疾患においてどのようにその機構が変化/破綻するか

を解説し，実際の研究法を幾つかの例をあげて説明する． Outline⟩ 病態に伴い動態が変化するタンパク質群を探索同定するための研究法と，このようなプロテオーム

解析を通して得られた知見の診断法や新薬開発への応用について講義する． Manner⟩ オムニバス方式 Notice⟩ 特になし． Schedule⟩

大項目

1. {ゲノム一次配列の多様性とタンパク質が構築する機能ネットワークの多様性とを比較 k する基礎知識と研究法について講義 \atop する．}

Evaluation Criteria⟩ 受講状況と授業内容の理解の程度に応じて評価する． Re-evaluation⟩ 行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217565 Contact⟩

⇒ Siomi . ⇒ Sakaguchi (sakaguch@ier.tokushima-u.ac.jp)

Applied Molecular Enzymology and Pathogenesis 2 units (selection) 1st-year(1st semester)

Hiroshi Kido⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 蛋白質・酵素の機能解析の基礎理論の習得と，実験操作などの基礎的技術力の習得を目的とする． Outline⟩ 最初に蛋白質と酵素の機能解析の基礎理論について習得し，これらの知識を基に，病気の原因，治療

法，予防法の解明の実際例について学ぶ．講義では最新の情報を可能な限り取り入れた講義を行う． Notice⟩ 講義と演習を行う． Schedule⟩

大項目中項目内容担当
1. 酵素学の基礎知識の講義の後， 蛋白質分解酵素を例にした病 気の原因の解明と治療に関す る講義を行う．酵素学の基礎知識 1  酵素と蛋白質の基礎知識木戸博教官

2.  〃酵素学の基礎知識 2  蛋白質精製の基礎知識〃
3.  〃酵素学の基礎知識 3  〃〃
4.  〃酵素学の基礎知識 4  酵素，タンパク質研究の手技〃
5.  〃酵素学の基礎知識 5  〃〃
6.  〃{酵素学からアプローチする病気の原因 \atop 解明 1} インフルエンザ感染と酵素〃
7.  〃{酵素学からアプローチする病気の原因 \atop 解明 2} インフルエンザ脳症と酵素〃
8.  〃{酵素学からアプローチする病気の原因 \atop 解明 3} インフルエンザの予防，治療と酵素〃

9. 蛋白質の立体構造を監視して 構造のゆがみを修復する分子 シャペロンの機能を中心に，シ ャペロンが関与する病態につ いても講義する．酵素学の基礎知識 1  分子シャペロンの基礎知識矢野仁康教官

10.  〃酵素学の基礎知識 2  〃〃
11.  〃{酵素学からアプローチする病気の原因 \atop 解明 1} コンフォメーション病 1  〃
12.  〃{酵素学からアプローチする病気の原因 \atop 解明 2} コンフォメーション病 2  〃
13.  〃{酵素学からアプローチする病気の原因 \atop 解明 3} コンフォメーション病 3  〃
14.  〃{酵素学からアプローチする病気の原因 \atop 解明 4} 分子シャペロンと病態 1  〃
15.  〃{酵素学からアプローチする病気の原因 \atop 解明 5} 分子シャペロンと病態 2  〃

Evaluation Criteria⟩ 出席とレポート提出 (出席 2/3 以上をもって合格とする)． Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217588

Note⟩ e-Learning の取扱い無し．

Medical Electronics 2 units (selection) each-year(1st semester)

Kazuo Yoshizaki · Professor / Course of Preventive Medicine, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

, Yohsuke Kinouchi · 名誉教授

Target⟩ バイオメディカル研究者，医療従事者として必要な電気電子機器の基礎知識を身につけると共に，最

新応用知識を修学する． Outline⟩ バイオメディカル研究者，医療従事者として必要な電気電子機器の活用法，情報処理解析法，機器分

析法についての基本的な知識と医学応用について講義・演習を行う． Schedule⟩

大項目担当
1. メディカルエレクトロニクス授業ガイダンス吉崎教授
2. 心音図〃
3. 膜電位電流計測〃
4. 心電図〃

5. 機器分析法 (NMR 法など)  〃
6. 神経・筋活動勢井教授
7. 脳波・脳磁図・筋電図・姿勢制御〃
8. 循環・呼吸モニタ〃
9. 睡眠記録・解析〃
10. 磁気刺激〃

11.

Medical Information System 森口教授

12.

Personal information protection and P-Mark 〃

13.

Hospital strategic management 1 森川准教授

14.

Hospital strategic management 2 〃

15. 電子機器の医学への応用とその実際木内教授

Evaluation Criteria⟩ 出席とレポート提出 (出席は 2/3 以上をもって合格とする) Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217576 Contact⟩

⇒ Yoshizaki (+81-88-633-7055, kyoshi@basic.med.tokushima-u.ac.jp) (Office Hour: e-mailで時間調整の

上，面談して下さい． ) ⇒ 他の教員についても， e-mailにて時間調整の上，面談してください．

Medical Application of Nanotechnology 2 units (selection) 1st-year(1st semester)

Toshio Doi⟨Manager⟩ · Professor / Course of Biological Medicine, Medical Science

, Hirokazu Miyoshi · Associate Professor / Radioisotope Research Center

Tatsuhiro Ishida · Associate Professor / Institute of Health Biosciences

, Mitsuru Ishikawa · / National Institute of Advanced Industrial Science and Technology

Target⟩ ナノテクノロジーの現状を理解させる． Outline⟩ 近年のナノテクノロジーの進歩を背景として，この技術を医療応用するための基礎知識，臨床検査

上の現状と問題点を講義すると共に，実際の医療応用例を示して基礎研究と実用化のプロセスを学ぶ． Notice⟩ なし Schedule⟩

大項目担当
1. 疾病とナノテクノロジーの応用土井俊夫
2. 超微細加工技術と分子ナノテクノロジーの基礎と応用石川満
3. ナノテクノロジーの医療への応用石田竜弘
4. ナノテクノロジーと標識分子及びその応用三好弘一

Evaluation Criteria⟩ 出席及びレポートで評価する． Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217562 Note⟩ なし

Clinical Genetics 2 units (selection) 1st-year(1st semester)

Yutaka Nakahori ·, Youichi Sato · Associate Professor / Institute of Health Biosciences

Target⟩ 人類遺伝学の基礎的な知識を習得し，臨床的な応用，遺伝カウンセリングについて学習する． Outline⟩ 臨床遺伝学関連の本を選定し輪読会形式で読み進む．各学生が，予習と説明を担当する．最初の授

業で分担を決める予定． Notice⟩ 教科書に指定された本はコピーせずに買うこと． Schedule⟩

大項目

1. ゲノム，ヒトの遺伝子構造の基本を述べることができる．
2. 遺伝子の変化による疾病の発生機序を知っている．
3. 患者 (模擬患者) ・家族等に対し遺伝性疾患について話す際に注意すべき点を知っている．

Evaluation Criteria⟩ 出席，担当，報告に対する評価をもって成績を判定する． 1/3 以上の欠席は単位を与え

ない． Re-evaluation⟩ 出席が足らない分は，指定する授業，演習で満たすこと． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217583 Contact⟩

⇒ Nakahori . (Office Hour: 月曜日，金曜日の12時から13時(e-mail により時間調節を適宜おこないま

す)) ⇒ 他の教員についても,e-mail にて時間調節の上,面談して下さい.

Note⟩ 分担または当番がある．自分があたったところは，ただ訳したものを読むのではなく，人に説明できる程度には理解すること．

Development, Differentiation, and Regeneration 2 units (selection) each-year(1st semester)

Yousuke Takahama⟨Manager⟩ · Professor / Course of Preventive Medicine, Medical Science

Mitsuru Matsumoto · Professor / Course of Medicine for Biological Responses, Proteomics

, Sumihare Noji · Professor / Faculty of Engineering

· Part-time Lecturer

, Masahiko Hibi · Part-time Lecturer

Target⟩ 生体の発生・分化についての基礎知識を習得し，それにもとづく再生医療の可能性について理解する． Outline⟩ 発生現象の基礎とその分子機構および細胞分化の基礎とその分子機構について講義する．また，胚

性培養細胞等を用いた再生医療について解説する． Schedule⟩

大項目担当
1. オリエンテーション高濱洋介 (第 1 回目の講義開始時刻に開催)
2. 発生学 1  野地澄晴
3. 発生学 2  〃
4. 細胞分化 1  高濱洋介
5. 細胞分化 2  〃
6. 発生工学松本満
7. 再生医学〃
8. 発生学特別講義上野直人
9. 発生再生医学特別講義日比正彦

Evaluation Criteria⟩ 習得状況および出席状況により評価する．詳細はオリエンテーションにて説明する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217601 Contact⟩

⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

(Office Hour: 水曜日16:00∼ 18:00(電子メイルにより適宜面談時刻を調整すること)． ) ⇒ 他の教員についても電子メイルにて面談時刻を調節すること．

Project Search Study Course 4 units (compulsory) 1st-year(whole year), 2nd-year(whole year)

Kiyoshi Fukui⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Kazunori Ishimura · Professor / Course of Integrated Brain Sciences, Medical Science

, Yousuke Ebina · Professor / Course of Medicine for Biological Responses, Proteomics

Kazuo Yoshizaki · Professor / Course of Preventive Medicine, Medical Science

, Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics

Toshiaki Tamaki · Professor / Course of Biological Medicine, Medical Science

, Haruhiko Siomi ·

Yoshihiro Fukui · Professor / Course of Human Development, Medical Science

, Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics Mitsuru Matsumoto · Professor / Course of Medicine for Biological Responses, Proteomics

Hisaaki Taniguchi · Professor / Course of Medicine for Biological Responses, Proteomics

, Takuya Sasaki · Professor / Molecular Medicine, Proteomics

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

, Toshiaki Sano · Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Yutaka Nakahori ·, Shin-ichi Kubo ·, Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Toshio Matsumoto · Professor / Molecular Medicine, Proteomics

, Toshio Doi · Professor / Course of Biological Medicine, Medical Science

Shinji Nagahiro · Professor / Course of Integrated Brain Sciences, Medical Science

, Kyoji Morita ·

Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics

, Kozo Matsumoto ·

Itsuo Tokunaga · Associate Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Shoji Kagami · Professor / Course of Human Development, Medical Science

, Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science , Noriaki Takeda · Professor / Course of Sensory Neuroscience, Medical Science Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Target⟩ ヘルスバイオサイエンス領域の世界をリードする最先端研究の成果を学修して，課題探求能力の開発

と，独創的な発想力並びに研究プロジェクト推進能力を養成する． Outline⟩ 最新の内外の代表的学術論文を学生に抄読させ，作業仮説の構築とその実証過程について学修する．

研究計画の立案方法および研究実施方法とその結果に関して，学生による発表討論を行い，課題探求能力開発と研究推進能力育成を目的とした総括的な指導を行う．さらに，病態の解明と治療薬開発をめざす医科学研究の戦略を考察する．

Notice⟩ e-learning の採用とその取り扱いは各担当教員の実施計画に従う． Schedule⟩

大項目

1. 各担当指導教員の研究担当分野の学術論文等を教材として，各教員の研究指導計画に基づいて，通年にわたって実施する．

Evaluation Criteria⟩ 受講状況と演習内容の理解の程度に応じて評価する． Re-evaluation⟩ 行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217610 Contact⟩

⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 月曜日の11:00∼ 12:00(e-ma

ilにより時間調節を適宜おこないます)) ⇒ 他の教員についても， e-mailにて時間調節の上，面談をしてください．

メディカルサイエンス特別研究 12 units (compulsory) 2nd-year

修士課程全教員

Target⟩ 講義・実習で習得した知識，実験手法などを実際に研究の場でどう生かせるかを試行することを目的

とする． Outline⟩ 研究課題を設定させ，研究計画の作成，実験の進め方，実験手法の習熟，データの整理・解釈を指導

して，研究成果を修士論文としてまとめさせる． Notice⟩ 特になし． Schedule⟩

大項目内容
1. 形態・機能解析領域，分子細胞機能解析領域ホルモンの作用機構，脳・神経機能，細胞増殖・分化機構，体液循環調節，細胞情報伝達機 構，細胞内分子輸送，プロテオミクス医科学などの研究課題から研究テーマを選択させ，研究 指導を行う．

2. 病理・病態解析領域{内分泌病理，発癌機構，細菌ゲノミクス，感染・免疫機構などの研究課題から研究テーマを選 \atop 択させ，研究指導を行う．}

3. 医療社会学領域{性分化，薬毒物による中枢神経障害機構などの研究課題から研究テーマを選択させ，研究指導 \atop を行う．}

4. 臨床医学領域{内分泌・代謝疾患，消化器癌，臓器移植，脳血管障害，腎障害，重傷脳障害の救急・集中治療 \atop などの研究課題から研究テーマを選択させ，研究指導を行う．}

Evaluation Criteria⟩ 修士論文の内容を発表させ評価する．

Re-evaluation⟩ 行わない．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217578

Chapter 2

Doctor’s Degree

Graduate School of Medical Sciences (2011) ⟩ Doctor’s Degree ⟩ Course of Medicine

Course of Medicine Syllabus of Subjects

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ...17 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) .................17

Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka

・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) .......................................18

Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・

Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester)..............................................18 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao

・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ...................................................... 19 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・

Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga .....................................19

• 各専攻系間共通カリキュラム科目

Introduction to Clinical Medicine … Kaji ・ Ito ・ Soeki ・ Doi ・ Bando ・ Nakajo ・ Kawano ・ Takeda ・ Saito / each-year(1st

semester) .......................................................................................... 20 A Training of Microbiology and Immunology … Miyake ・ Adachi ・ Yasutomo ・ Ono ・ Kuwahara ・ Uchiyama ・

Maekawa ・ Hirota / 1st-year(2nd semester), 2nd-year(2nd semester) .......................................... 20

• 指定科目

Basic practice for physiological and pharmacological research … Tamaki ・ Yoshizaki ・ Sei ・ Sano ・

Minakuchi / 1st-year(1st semester) ......................................................................20

Introductory lab course for biochemical and molecular biology experiments … Takahama ・

Oyadomari / each-year(1st semester) .................................................................... 21

• 専門科目

Lecture for Birth Defects … Fukui ・ Sakata-Haga / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)21 Seminar for Birth Defects … Fukui ・ Sakata-Haga / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

21

Developmental Pathophysiology … Kagami・Mori・Hayabuchi・Watanabe・Kondo / 1st-year(whole year), 2nd-year(whole

year), 3rd-year(whole year) ...........................................................................22 Practical developmental pathophysiology … Kagami ・ Mori ・ Hayabuchi ・ Watanabe ・ Kondo / 1st-year(whole year),

2nd-year(whole year), 3rd-year(whole year) .............................................................22 Obstetrics and Gynecology … Irahara ・ Furumoto ・ Yasui ・ Matsuzaki ・ Maeda ・ Kuwahara ・ Yamamoto / 1st-year(whole

year), 2nd-year(whole year), 3rd-year(whole year) ....................................................... 22 Practice of Obstetrics and Gynecology … Irahara ・ Furumoto ・ Yasui ・ Matsuzaki ・ Maeda ・ Kuwahara ・

Yamamoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................ 22 発生発達医学実験実習 … Fukui ・ Kagami ・ Irahara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)23 Physiology … Yoshizaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................23 Practice of Physiology … Yoshizaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........23 Gastroenterology and Oncology … Takayama ・ Okahisa ・ Muguruma / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 23 Practice of Gastroenterology and Oncology … Takayama・Okahisa・Muguruma / 1st-year(whole year), 2nd-year(whole

year), 3rd-year(whole year) ...........................................................................23 Immune system development … Takahama ・ Takada / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................24 Immune system development … Takahama ・ Takada / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................24 Immune Regulation … Okazaki ・ Okazaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..... 24 Practice of Immune Regulation … Okazaki ・ Okazaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................24 Experimental practice for preventive medicine … Yasutomo ・ Yoshizaki ・ Takahama ・ Kishihara ・ Shimizu ・ Tomita

・ Honda / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................. 24 Immunology … Yasutomo ・ Kishihara ・ Maekawa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .24 Practice for Immunology … Yasutomo・Kishihara・Maekawa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................24 Microbiology … Adachi ・ Kido ・ Uchiyama / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)......25 Practice for Microbiology … Adachi ・ Uchiyama / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)25 Practice for Microbiology and Immunology … Yasutomo ・ Adachi / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 25 Preventive Medicine … Arisawa ・ Uemura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..... 25 Practice of Preventive Medicine … Arisawa ・ Uemura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................26 Community Medicine … Tani / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............... 26 Practice of Community Medicine … Tani / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...26 Practice of Social and Environmental Medicine … Arisawa / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 26 Anatomy and Cell Biology … Ishimura ・ Nakamura / 1st-year(1st semester, 2nd semester), 2nd-year(1st semester, 2nd

semester), 3rd-year(1st semester, 2nd semester)..........................................................26 Practice of Anatomy and Cell Biology … Ishimura ・ Nakamura / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 27 Integrative Physiology … Sei / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............... 27 Practice of Integrative Physiology … Sei ・ Shiuchi ・ Fujihara ・ Chikahisa / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 27 Psychiatry … Ohmori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................27 Practical Psychiatry … Ohmori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............27 Brain Pathophysiology and Medicine … Nagahiro / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................27 Practice of Brain Pathophysiology and Medicine … Nagahiro / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 27 Strategic medical informatics … Moriguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .. 27 Practice of strategic medical informatics … Moriguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................28 Molecular Biology … Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............28 Practice of Molecular Biology … Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .28

Practice for Integrated Brain Sciences and Medical Information … Ishimura ・ Sei ・ Ohmori ・ Nagahiro ・

Moriguchi ・ Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................28 Pharmacology … Tamaki ・ Minakuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........28 Practice for Pharmacology … Tamaki ・ Minakuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

28

Anesthesiology … Oshita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).....................29 Practice of Anesthesiology … Oshita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........ 29 Nephrology … Doi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........................29 Seminar for Nephrology … Doi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............29 救急集中治療医学 … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............. 29 救急集中治療医学演習 … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........29 Experimental Practice for Biological Medicine … Tamaki・Oshita・Nishimura / 1st-year(whole year), 2nd-year(whole

year), 3rd-year(whole year) ...........................................................................29 Ophthalmology … Mitamura ・ Eguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........29 Practice of Ophthalmology … Mitamura ・ Eguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)29 Otorhinolaryngology … Takeda ・ Tamura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...... 30 Experimental Practice for Otorhinolaryngology … Takeda ・ Tamura / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 30 Clinical Neuroscience … Kaji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............... 30 Clinical Neuroscience Workshop … Kaji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .... 30 Sensory Neuroscience Laboratory Experiment … Takeda ・ Shiota ・ Kaji ・ Tamura ・ Naito ・ Izumi / 1st-year(whole

year), 2nd-year(whole year), 3rd-year(whole year) ....................................................... 30 Human Pathology … Sakashita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...............31 Practice of Human Pathology … Sakashita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .. 31 Gastroenterological Surgery and Transplantation … Shimada ・ Utsunomiya ・ Kurita ・ Iwata ・ Nishioka ・ Morine ・

Imura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................... 31

Practice of Gastroenterological Surgery and Transplantation … Shimada ・ Utsunomiya ・ Kurita ・ Iwata ・ Nishioka

・ Morine ・ Imura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................31 Cardiovascular Surgery … Kitagawa ・ Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).31 Practice of Cardiovascular Surgery … Kitagawa・Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................32 Urology … Kanayama ・ Fukumori ・ Takahashi ・ Nakatsuji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................32 Experimental Practice for Urology … Kanayama ・ Fukumori ・ Takahashi ・ Nakatsuji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................32 Cardiovascular Medicine … Sata / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............32 Cardiovascular Medicine Practice … Sata / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...32

Laboratory Experiment Practice of Diseased Organ and Repair Medicine … Kitagawa ・ Sakashita ・ Shimada

・ Kanayama ・ Ishibashi ・ Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)..............33 Experimental Oncology … Izumi ・ Uehara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .... 33 Practice for Experimental Oncology … Izumi ・ Uehara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................33 Radiology … Harada ・ Otsuka ・ Ikushima ・ Ueno ・ Matsuzaki ・ Takeuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................ 33 Practice of Radiology … Harada ・ Ikushima ・ Otsuka ・ Matsuzaki ・ Takeuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................ 33 Respiratory Medicine and Rheumatology … Sone・Nishioka・Hanibuchi・Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................................34 Practice of Respiratory Medicine and Rheumatology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................34 Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................34 Practice of Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................ 34 Thoracic, endocrine surgery and Oncology … Tangoku ・ Sakiyama ・ Kondo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................................34 Surgical Oncology … Tangoku / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...............35

Experimental Practice for Bioregulation and Medical Oncology … Izumi ・ Nishitani ・ Sone ・ Tangoku / 1st-

year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................. 35 Reconstructional Pathology … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).....35 Practice of Reconstructional Pathology … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................35 Dermatology … Kubo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................... 35 Practice for Dermatology … Kubo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......... 35 Orthopedic Surgery … Yasui ・ Katoh ・ Sairyo ・ Takahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year), 4th-year......................................................................................35

Orthopedic Surgery Practice … Yasui / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......35

Plastic Surgery and Reconstructive Surgery … Nakanishi ・ Hashimoto / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 36

Practice for Plastic Surgery and Reconstructive Surgery … Nakanishi / 1st-year(whole year), 2nd-year(whole

year), 3rd-year(whole year) ...........................................................................36

Laboratory Experiment of Pathologic Medicine for Cutaneo-motoric System … Nishimura ・ Yasui ・

Nakanishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................36

Introduction to Biological Ethics 2 units (compulsory) each-year(1st semester)

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

, Seiichiro Kitamura · Professor / Institute of Health Biosciences

Junji Terao · Professor / Institute of Health Biosciences , Nobuyoshi Nakajo · Professor

, Kazuo Minakuchi · Professor / Institute of Health Biosciences

Target⟩ バイオサイエンスおよび医療に従事する者は，人権，生命倫理に十分な配慮を行い，個人情報保護，実験動物愛護にも同じく目を向けなければならない．本授業は生命倫理に関わる基本的知識を修得することを目的としている．

Outline⟩ 生命倫理学，臨床倫理学，社会倫理，個人情報保護，実験動物愛護などの問題に日頃接することの多

い講師が，経験に基づいた講義をオムニバス方式で行う． Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 「ヒトゲノム · 遺伝子解析に関する倫理指針」について板倉光夫
2. 「実験動物管理と情報処理にまつわる倫理学」太田房雄
3. 「キャンパスハラスメントを通して「医療従事者の倫理」を考える」 北村清一郎
4. 「法規や制度に関する基本的考え方—法医学の立場から」久保真一
5. 「大学における人権問題」佐野壽昭
6. 「組織検体取り扱い上の倫理的問題」〃
7. 「食品の機能性 · 安全性の評価と社会倫理」寺尾純二
8. 「尊厳死や安楽死など生命倫理について」中條信義
9. 「遺伝医学の視点」中堀豊
10. 「先天異常， (or 神経疾患，家族性腫瘍) の遺伝カウンセリング」〃
11. 「動物実験倫理」松本耕三
12. 「臨床治験に関する倫理」水口和生
13. 「ヒト胚の倫理学的な地位」山野修司
14. 「臓器移植の倫理的問題 (仮題)」井藤久雄 · 予定
15. (未定)

Evaluation Criteria⟩ 出席状況 (厳格にする予定) を踏まえ，さらに小テストないし面接試験で評価することを

検討している． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218234 Contact⟩

⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp)

⇒ Kitamura (口腔顎顔面形態学分野教授室, +81-88-633-7319, kitamura@dent.tokushima-u.ac.jp) (Office

Hour: 前後期:金曜日 12 時 ∼ 13 時) ⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp) ⇒ Nakajo (+81-88-633-7366, nakajo@dentclin.tokushima-u.ac.jp) ⇒ Minakuchi (+81-88-633-7212, minakuti@clin.med.tokushima-u.ac.jp)

Clinical Psychology 2 units (selection) 1st-year(1st semester)

Kenji Sato · Professor / Institute of Socio-Arts and Sciences

, Mayumi YAMAMOTO · Professor / Institute of Socio-Arts and Sciences

Motohiro Sakai · Associate Professor / Institute of Socio-Arts and Sciences , Takaki Fukumori · Associate Professor / Institute of Socio-Arts and Sciences

Masato Mitome · Professor / Institute of Health Biosciences

Target⟩ 臨床心理学の基礎的理論・技法および今日的課題の概説

Outline⟩ 心と身体は密接につながっている．したがって「心の問題」の理解と制御を扱う臨床心理学の基礎の習得は，精神医学・心身医学のみならず，ヘルスバイオサイエンスを基盤とする医学，歯学，薬学，栄養学，保健学領域において重要である．そこで，本講義では，臨床心理学の定義，対象，方法 (代表的な心理検査，心理療法) について，初学者を考慮して，その基礎と今日的課題を概説する．心理療法に関しては，医学領域などでエビデンスを示す認知行動療法について詳述する．

Notice⟩ e-learning 対応．

Goal⟩ 臨床心理学の基礎的理論・技法および今日的課題を説明できる

Schedule⟩

大項目中項目
1. 臨床心理学の定義と対象 (佐藤)
2. 臨床心理学の測定方法 (1)  パーソナリティの査定法 (福森)
3. 心理療法・カウンセリングとは (福森)
4. 心理療法の基礎 (1)  精神分析 (福森)
5. 臨床心理学の測定方法 (2)  知能，発達の査定法 (山本)
6. 心理療法の基礎 (2)  分析心理学，クライエント中心療法 (山本)
7. 発達障害と特別支援教育の現状と課題 (山本)
8. 行動理論と行動療法 (境)
9. 認知療法うつ病 (境)
10. 認知行動療法不安障害 (境)
11. 臨床心理的地域援助の基礎と実際ひきこもりに焦点を当てて (境)
12. 外傷後ストレス障害の認知行動療法 (佐藤)
13. 摂食障害の認知行動療法 (佐藤)
14. 体重減量・糖尿病の認知行動療法 (佐藤)

Evaluation Criteria⟩ 対面講義， e-learning とも受講とレポート提出を以て出席扱い．対面講義の場合，授業内の 15 分間で書く「出席・質問票・ミニレポート」の提出を以て， e-learning の場合，視聴記録 (受講) と MLS オンラインのレポート提出を以て出席とみなす．どちらのレポート課題とも「講義内容を A4 一枚でまとめる」．出席が三分の二以上の受講生が評価対象者

Re-evaluation⟩ 再試験無し

Textbook⟩ 教科書は使用しない．参考書などは，適宜，授業時に紹介する

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218235

Contact⟩ ⇒ Mitome (mitome@dent.tokushima-u.ac.jp) (Office Hour: 三留:水曜日の12:00∼ 13:00(e-mail により調

整可能))

Introduction to social medicine, epidemiology and biostatistics 2 units (selection)

each-year(2nd semester)

Kokichi Arisawa · Professor / Course of Social and Environmental Medicine, Medical Science

, Hiro-O Ito · Professor / Institute of Health Biosciences

Hirokazu Uemura · Associate Professor / Course of Social and Environmental Medicine, Medical Science

, Aiko Yamauchi · Professor / Institute of Health Biosciences

Tetsuya Tanioka · Professor / Institute of Health Biosciences

, Akira Tokumura · Professor / Institute of Health Biosciences

Daisuke Hinode · Professor / Institute of Health Biosciences , Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Hiroaki Yanagawa · Associate Professor / University Hospital

Target⟩ 1, Clinical trial as a tool to establish clinical evidence 2. The role of the IT in the medical treatment and the hospital management are outlined. 3. Outline analytical skills for clarifying the phenomenon about Mental Health and Welfare. 4. Outline epidemiology and analytical methods for oral health promotion and/or oral disease prevention. 6. Outline damages from medicines and its etiology and prevention. 7. Outline designing of surveys and statistical analysis of data for writing a thesis.

Outline⟩ 1) Principles and practice of clinical trial from scientific and ethical viewpoint will be reviewed. 2) The usage of IT and the method of the management in the university hospital are outlined. 3) We will explain the statistical method for clarifying the feature of phenomenon based on previous work in Mental Health and Welfare. 4) We will explain the etiology of oral disease and its prevention based on the evidence of oral sciences. 6) Instruct background and etiology of several examples of damages from medicines and discuss possible prevention of the damages. 7) Basic aspects of statistical analysis, such as type of data and appropriate treatment of data, and meaning of hypothesis testing, will be summarized. In addition, practice of data analysis will be done using SPSS.

Notice⟩ It is possible to take a lesson by e-learning. Goal⟩ The goal is to deepen one’s understanding of contents presented during lectures, regarding social

medicine, pharmaceutical science and dentistry.

Schedule⟩

大項目担当

1.

Basic handling and statistical analysis of data for article makingHirokazu Uemura

2.  〃〃

3.

Practice of epidemiology and statistical data analysis for writing a thesis

Kokichi Arisawa

4.  〃〃

5.

Principles and practice of clinical trial Hiroaki Yanagawa

6.  〃〃

7.

The role of IT in a hospital Hiroki Moriguchi

8.  〃〃

9.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (1)

Shu-ichi Ueno

10.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (2)

Tetsuya Tanioka

11.

Epidemiology of dental caries and periodontal disease, and the relationship between oral disease and systemic disease

Hiro-o Ito

12.

〃

Daisuke Hinode

13. 医薬品の情報管理，医薬品の安全対策木原勝
14.  〃〃

15.

What is damages from medicines? Why did the damages occurred? How to prevent the damages?

Akira Tokumura

16.

What is the damages from medicines? Why did the damages occurred? How to prevent the damages?

〃 Evaluation Criteria⟩ Reports should be submitted on questions presented during lectures. Grades are

determined by summing points from all instructors. Re-evaluation⟩ In principle, there is no reexaminations. Textbook⟩ Teaching materials designated by each instructor wil be used.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218231

Contact⟩

⇒ Arisawa (+81-88-633-7071, arisawa@basic.med.tokushima-u.ac.jp) (Office Hour: Please contact eac

h instructor by an e-mail.)

Note⟩ Different classrooms may be used according instructors.

Introduction to How to Write Up Scientific Manuscript in English 2 units (selection)

each-year(2nd semester)

Shuhei Tomita · Professor / Course of Biological Medicine, Medical Science

, Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Yutaka Nakaya · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Hirofumi Nagashino · Professor / Institute of Health Biosciences

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Bukasa Kalubi · Assistant Professor / Medical Science

Target⟩ 21 世紀に医学，歯学，薬学，栄養学，保健学の各分野で活躍する人材は英語が堪能であることが要求される．本授業ではこれらの領域で用いられる独特の英語表現法に関わる基本的知識を修得することを目的とする．

Outline⟩ 研究成果を国際的な学術雑誌に発表したり，海外の学会やシンポジウムで発表や講演をしたりすることは，研究者養成を目的とする大学院教育において必須の履修目標である．そこで，本授業では医学英語論文，用紙のの作成方法について系統的な講義をビデオ並びにマルチメディア教材等を積極的に活用しながら行う．更に，医科学用英語の聴き取り及び英語による討論の訓練を行い，発表技術の向上を目指す． Notice⟩ 1) 授業は後期のみに開講される．時間帯は原則的に火曜日の午後であるが，講師の事情等により変更されることがある． 2) 講義の一部は e-learning 化されているので， e-learning 学習も出席として取り扱う． 3)20 分以上の遅刻は出席と見なさない．

Schedule⟩

大項目担当

1.

生物医学雑誌投稿に要求される条件 (I)

Fukui Kiyoshi

2. 生物医学雑誌投稿に要求される条件について (II)  〃
3. 研究，論文，学会発表の進め方 (I)  市原明

4.

研究，論文，学会発表の進め方 (II)

Ichihara Akira

5. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (I) 長篠博文
6. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (II)  〃
7. 学会口演要旨作成上の留意点について実例をまじえて冨田修平
8. 論文作成に役立つコンピュータの利用法ー実例を示した結果，考察の書き方ー 中屋豊

9.

英語による PowerPoint presentation についての要点

梶龍兒

10. 英語論文の特性と日本語論文との比較林良夫
11. 口頭による英語発表と短報の書き方に関する基本を講義する I  岡崎拓
12.  〃〃
13. 研究の区切りとしての論文作成 I  福井裕行
14. II

Evaluation Criteria⟩ 講義への出席，随時の試験，受講態度を総合して行う． 2/3 以上の出席がなければ不合

格とする．再試験はない． Textbook⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218236

Psychosomatic health and environmental stress 2 units (selection)

Eiji Takeda · Professor / Institute of Health Biosciences

, Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Junji Terao · Professor / Institute of Health Biosciences

, Ken-ichi Miyamoto · Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Rintaro Yamanishi · Associate Professor / Institute of Health Biosciences

Hiroshi Sakaue · Associate Professor / Institute of Health Biosciences

, Toshio Hosaka · Assistant Professor / Institute of Health Biosciences

Target⟩ The aim of this program is to study on psychosomatic stress in daily life. Contents of this program

include gene expression profile under stress, stress response in health and diseases. Outline⟩ By the systems biology, clinical neurology, nutritional science techniques, the effects of stress on

development and growth in child will be assessed and discussed. �������⟩ psychosomatic health, environment, stress Goal⟩ To understand stress assessment technology

Schedule⟩

大項目担当

1.

Anxiety and depression (Tetsurou Ohmori) Noriaki Okazaki

2.

The development of the immune system (Yousuke Takahama)

〃

3.

Mechanism of sleep regulation (Hiroyoshi Sei)

Ryuji Kaji

4.

Endoplasmic reticulum stress in health and diseases (Seiichi Oyadomari)

Kazuhito Rokutan

5.

Stress genomics (Kazuhito Rokutan)

〃

6.

New concept and innovation of functional foods in Japan (Junji Terao)

Hiroyoshi Sei

7.

Why do we need ”stress control” through foods and food style? (Junji Terao)

〃

8.

Aging and mineral nutrition (ken-ichi Miyamoto)

Megumi Goto

9.

Caloric restriction and bone metabolism (Ken-ichi Miyamoto)

〃

10.

Breakfast and chewing for stress (Eiji Takeda) Hiroshi Chuman

11.

Food and stress related hormone (Eiji Takeda) Suehiro Sakaguchi

12.

Physical impact caused by stress (Yutaka Nakaya)

Eiji Takeda

13.

Gut immunity and stress (Yutaka Nakaya) Tetsurou Ohmori

14.

Oxidative stress and immunity (rintaro Yamanishi)

Gen Tamiya

15.

The role of adipokines on metabolic stress in obesity and obesity-related diseases (hiroshi Sakaue)

〃

16.

Diabetes mellitus and stress (Toshio Hosaka)

Eiji Takeda

Evaluation Criteria⟩ Reports on more than two third of the studies will be required for evaluation. Re-evaluation⟩ 再試験無し Textbook⟩ none Reference⟩ 授業時に適宜，紹介する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218232 Contact⟩

⇒ Takeda (+81-88-633-7093, takeda@nutr.med.tokushima-u.ac.jp) ⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour:

8:30-17:30) ⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp) ⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp) ⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp)

⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp)

⇒ Miyamoto (Molecular Nutrition, +81-88-633-7081, miyamoto@nutr.med.tokushima-u.ac.jp)

⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Yamanishi (+81-88-633-9366, rintaro@nutr.med.tokushima-u.ac.jp) ⇒ Sakaue (303, +81-88-633-7091, hsakaue@nutr.med.tokushima-u.ac.jp) ⇒ Hosaka .

Basic methods in life sciences 2 units (selection)

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Hiroyuki Mizuguchi · Associate Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Tatsuji Haneji · Professor / Institute of Health Biosciences

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Kazuko Seta · Part-time Lecturer

Kouji Itou · Professor / Institute of Health Biosciences

, Takahiro Matsumoto · Associate Professor / Institute of Health Biosciences

Keiko Kataoka · Associate Professor / Course of Microbiology and Immunology, Medical Science

, Tetsuo Yamazaki · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Takeshi Nikawa · Professor / Institute of Health Biosciences

Hironori Yamamoto · Assistant Professor / Institute of Health Biosciences

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics , Toshihiko Nagata · Professor / Institute of Health Biosciences Tetsuya Yoshinaga · Professor / Institute of Health Biosciences

Target⟩ 大学院に進学した直後の院生に生命科学に関する基礎的な実験方法を理解させる．

Outline⟩ 実験動物の取り扱い方，細胞と器官の培養方法とその応用，蛋白質の取り扱い方とその解析方法，遺伝子解析の方法とその応用，抗体を用いた研究方法とその蛋白質，免疫組織細胞化学に対する応用，数理モデルによる生体機能発現機構の解明等生命科学の基礎的な研究技法を講義する．

�������⟩ 生命科学, 実験手法

Notice⟩ 受講生は必ず出席すること．

Goal⟩ 現在行われている生命科学研究方法の基礎を大学院進学直後に理解し，研究生活にスムーズに入れるよ

うにする． Schedule⟩

大項目

1. {大学院に入学直後に現在用いられている研究手法を集中的に講義する．第 1 回目は授業ガイダンスと生命科学の研究手法総 \atop 論 (羽地，研究部長)}
2. 実験動物を利用した研究 (松本)
3. 細胞と器官の培養 (羽地)
4. 細胞培養の応用 (宮本)
5. 蛋白研究 1(福井清)
6. 蛋白研究 2(藤原)
7. 抗体を用いた生化学的研究方法 (二川，山本)
8. 抗体を用いた研究の臨床応用 (中屋)
9. 免疫学研究の基礎 (安友)
10. 免疫学研究の最近の進歩 (安友)
11. 遺伝子解析 1(福井裕)
12. 遺伝子解析 2(伊藤)
13. 遺伝子解析 3(高濱)
14. 遺伝子解析 4(高濱)
15. 数理モデルによる生命機能の解析 (吉永)

Evaluation Criteria⟩ Web によるレポート，出席状況，受講態度等により総合的に判定する． Re-evaluation⟩ しない Textbook⟩ 指定しないが，講義の都度プリント等資料を配布する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218233 Contact⟩

⇒ Fukui (hfukui@ph.tokushima-u.ac.jp) ⇒ Mizuguchi (+81-88-633-7264, guchi003@ph.tokushima-u.ac.jp) ⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Haneji (+81-88-633-7321, tat-hane@dent.tokushima-u.ac.jp) ⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 8:30 to 17:30) ⇒ Seta . ⇒ Itou (Institute for Medicinal Resources, +81-88-633-7290, kitoh@ph.tokushima-u.ac.jp) (Office Hour:

Monday, 17:0019:00) ⇒ Matsumoto . ⇒ Kataoka (kataoka@basic.med.tokushima-u.ac.jp) ⇒ Yamazaki (医学臨床 B 棟 5 階医薬品病態生化学研究室, +81-88-633-7886, yamazakt@ph.tokushima-u

.ac.jp) (Office Hour: 特に定めない) ⇒ Okazaki (414, +81-88-633-9158, tokazaki@genome.tokushima-u.ac.jp) ⇒ Nikawa (nikawa@nutr.med.tokushima-u.ac.jp) ⇒ Yamamoto (+81-88-633-7235, yamamoto@nutr.med.tokushima-u.ac.jp) ⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp) ⇒ Nagata (+81-88-633-7343, nagata@dent.tokushima-u.ac.jp) ⇒ Yoshinaga (保健学 B 棟教員研究室 (吉永), +81-88-633-9050, yosinaga@medsci.tokushima-u.ac.jp)

(Office Hour: 18:00-19:00 every Friday)

Introduction to Clinical Medicine 2 units (selection) each-year(1st semester)

Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

, Susumu Ito · Professor

Takeshi Soeki · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science

, Toshio Doi · Professor / Course of Biological Medicine, Medical Science

Eiichi Bando · Professor , Nobuyoshi Nakajo · Professor

, Fumiaki Kawano · Professor / Institute of Health Biosciences

Eiji Takeda · Professor / Institute of Health Biosciences

, Ken Saito · Professor / Institute of Health Biosciences

Target⟩ 臨床医学の基礎知識について概説する． Outline⟩ 講義 Goal⟩ 循環器，呼吸器，消化器，腎臓，神経 · 筋，内分泌 · 代謝，血液の各臨床領域における代表的な疾病に

つき，発生機序および原因となる遺伝子などの異常，そして各々の疾患の病態生理を理解させ，最新の診断および治療法の理論と実践を学ばせる．

Schedule⟩

大項目中項目
1. 消化器疾患の診断と治療 (担当伊東進)
2. 循環器疾患の診断と治療 (担当添木武)
3. 臨床医科学概論・神経系総論 1 ・ 2 ・ 3(担当 梶龍兒)
4. 腎臓病 1 ・ 2(担当土井俊夫)
5. 顎口腔機能入門 1 ・ 2(担当坂東永一)
6. 麻酔薬と精神機能 1 ・ 2(担当中條信義)
7. 顎顔面領域の手術後のリハビリテーション 1 ・ 2(担当 河野文昭)

8.

Medical nutrition theory

9.

Nutrition and psychotic health

10. 循環器疾患の病態解析と時間医学 (担当齋藤憲)

Evaluation Criteria⟩ 出席状況と小テストまたはレポート等により評価する．

Textbook⟩ なし

Reference⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217052 Contact⟩

⇒ Kaji (rkaji@clin.med.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00-18:00 (e-mail により時間調節を適宜おこないます: rkaji@clin.med.tokushima-u.ac.jp)．他の教員についても， e-mail にて時間調節の上，面談して下さい． )

A Training of Microbiology and Immunology 2 units (selection)

1st-year(2nd semester), 2nd-year(2nd semester)

Yoichiro Miyake · Professor / Institute of Health Biosciences

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science

, Tsuneko Ono · Professor / Institute of Health Biosciences

Tomomi Kuwahara · Associate Professor / Course of Microbiology and Immunology, Medical Science Tsuneo Uchiyama · Associate Professor / Course of Microbiology and Immunology, Medical Science Yoichi Maekawa · Associate Professor / Course of Microbiology and Immunology, Medical Science Katsuhiko Hirota · Associate Professor / Institute of Health Biosciences

Target⟩ 微生物の取り扱いの基本，及び免疫学的実験法の基本を学ぶ

Outline⟩ · ·

Notice⟩ 1．病原微生物実習室で行うので，それ専用の自分の白衣とスリッパ (上履き) を持参のこと (桑原知

巳)． 2．白衣と筆記用具を持参のこと (三宅洋一郎 · 弘田克彦)． Schedule⟩

大項目中項目
1. (1˜ 4) 細菌学実験の基本操作と常在菌や病原菌の分離 (桑原知巳) {培地の作製と培養，滅菌 · 消毒法，無菌操作などの基本的手技を学び，常在菌 \atop や病原菌の分離 · 同定を行い，身近な細菌の性状を理解する．}

2. (5, 6) 分子ウイルス学的解析法 1(足立昭夫) バイオハザードや分 子遺伝学等，現代ウイルス学の基礎知識を習得するとともに基 本的技能について学ぶ．

3. (7, 8) 分子ウイルス学的解析法 2(内山恒夫) ウイルスの同定法や ウイルス感染の解析法に関する知識を習得し，その基本的手技 について学習する．

4. {(9,10) リンパ球活性化機構 (安友康二) リンパ球の分離方法の基 \atop 本的手技と，その活性化測定方法を学ぶ．}
5. {(11) マラリア原虫の診断 (前川洋一) マラリア原虫の分類とその \atop 鑑別方法を学ぶ．}

6. (12˜ 14) 口腔常在菌とバイオフィルム (三宅洋一郎 · 弘田克彦) 口 ; 腔常在菌の分離培養 · 同定などに関する基本的手技と，細菌バ ; イオフィルムの作成法及びそれを用いた各種の実験法を学ぶ．

7. {(15) 細菌感染症の分子遺伝学的検出法 (小野恒子) 菌種特異的遺 \atop 伝子マーカーを用いた細菌の検出と定量法について学ぶ．}

Evaluation Criteria⟩ 出席率と実習レポートによる．

Re-evaluation⟩ (再試験) 実習なので再試験はしない．出席率が 60%に満たない学生は，受講しなかった実習

を，次学年後期に追加して受けること． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217022 Contact⟩

⇒ Miyake (+81-88-633-7329, miyake@dent.tokushima-u.ac.jp) (Office Hour: 8:00∼ 8:45(e-mailでも対応します))

Basic practice for physiological and pharmacological research 2 units (selection)

1st-year(1st semester)

Toshiaki Tamaki⟨Manager⟩ · Professor / Course of Biological Medicine, Medical Science

Kazuo Yoshizaki · Professor / Course of Preventive Medicine, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Atsuko Sano ·, Kazuo Minakuchi⟨Manager⟩ · Professor / Institute of Health Biosciences

Target⟩ 生命科学研究の基礎となる生理学・薬理学研究法を習得するように指導する． Outline⟩ 生命科学を研究する上で必要となる生理学・薬理学的研究法について，適応方法および結果の解析

方法などの基本を指導する． Non-MD の大学院生については，生体内の調節機構，薬力学・薬物動態学などについても指導を行う． (分光分析法，磁気共鳴法，電気生理学的研究法，培養細胞法，実験動物の扱い方，行動生理学・薬理学的研究法，無線による生体信号記録法，腎微小循環研究法，生体内活性酸素・フリーラジカル測定法，細胞内情報伝達研究法，臨床薬理学的研究方法など)

Manner⟩ オムニバス方式 Notice⟩ 全ての実習に参加すること．実習のため， e-learning は利用できない． Schedule⟩

大項目中項目担当
1. 授業ガイダンス+実験動物の扱い方勢井宏義
2. 行動生理薬理学的研究法の概説勢井宏義・佐野敦子
3∼4. {ラット脳定位固定装置の操作，慢性電極等の植込み \atop 手術の見学・実習体験}勢井・佐野
5∼6. 上皮膜輸送の測定吉崎和男
7∼8. 磁気共鳴法〃
9∼10. {睡眠覚醒状態，中枢作用薬の睡眠覚醒に対する効果 \atop などの記録及び解析}勢井・佐野

11∼12.タンパク定量法玉置俊晃・森田恭二

13∼15.細胞損傷の研究生存率測定法玉置俊晃・水口和生・森田恭二

Evaluation Criteria⟩ 実習のレポートにて評価する Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217599 Contact⟩

⇒ e-mail にて時間調節の上，面談して下さい．

Introductory lab course for biochemical and molecular biology experiments

2 units (selection) each-year(1st semester)

Yousuke Takahama⟨Manager⟩ · Professor / Course of Preventive Medicine, Medical Science Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Target⟩ 医科学研究の基礎技術としての生化学・分子生物学的研究手法の基礎を修得し，それらの適応と限界

を理解する． Outline⟩ 遺伝子組換え技術の基礎を解説するとともに実習により指導する．また，ゲノム解析や遺伝子発現解

析の原理と応用について解説するとともに， PCR に基づくゲノム・遺伝子解析技術を指導する．更に，タンパク質の分離精製法，同定法，定量法について基礎理論を解説するとともに実習により基礎技術を指導する．いずれの実習についても，実際にそれぞれの技術を活用している研究室にて実施する．

Schedule⟩

大項目内容担当

1.

{組換え DNA とゲノム解析 (講義 45 分 ×2， \atop 実習 45 分 ×4)} 講義項目: ゲノム診断・法医学的検査・ DNA 抽出・サザン・ PCR ・データベース利用 / 実習項目: PCR

高浜

2.

{遺伝子発現解析 (講義 45 分 ×2，実習 45 \atop 分 ×4)} 講義項目: ノザン・ RT-PCR ・ライブラリスクリーニング・アレイ解析 / 実習項目: RT-PCR

親泊

3. {タンパク質の精製 (講義 45 分 ×2，実習 45 \atop 分 ×4)} ; {講義項目: カラムの種類と特性・タグの種類と特性・ゲル染色法 / 実習項目: \atop アフィニティクロマトグラフィ・ SDS-PAGE} ; 〃

4.

{タンパク質の同定 (講義 45 分 ×2，実習 45 \atop 分 ×4)} 講義項目: 電気泳動法の種類と特性・ウェスタン・マススペック・ ELISA / 実習項目: ELISA

〃

5. {DNA クローニング (講義 45 分 ×2，実習 \atop 45 分 ×4)} ; 講義項目: DNA クローニング / 実習項目: 制限酵素処理と電気泳動 ; 高浜

Evaluation Criteria⟩ 技術習得状況および出席状況により評価する．単位認定には，講義と実習から構成され

る上記 5 セットのうち，最低 3 セットへの出席が必要である． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217596 Contact⟩

⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

(Office Hour: 水曜日16:00∼ 18:00(電子メイルにより適宜面談時刻を調整すること)． ) ⇒ 他の教員についても電子メイルにて面談時刻を調節すること．

Lecture for Birth Defects 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshihiro Fukui⟨Manager⟩ · Professor / Course of Human Development, Medical Science Hiromi Sakata-Haga · Associate Professor / Course of Human Development, Medical Science

Target⟩ 環境因子，遺伝因子によって引き起こされる中枢神経系の発生・発達障害の成立機構を神経解剖学的，

分子生物学的に把握し，先端的研究にも対応できるようになる． Outline⟩ 子宮内環境の影響を強く受ける胎児中枢神経系の発生異常と生後の学習障害や精神発達遅滞 (胎児性

アルコール症候群等) との関連性を神経解剖学的，分子生物学的視点から解説，講義する．また，遺伝性小脳性運動失調の発祥機序を学ぶ．情報伝達系 (細胞内 Ca2+調節機構) の異常，神経修飾因子及びこれらの受容体・トランスポーターの分布を量的変化に注目して免疫組織化学的，分子生物学的視点から体系的に学ぶ．

Manner⟩ オムニバス方式 Notice⟩ 受講前に書く授業の担当教官と連絡を取り，受講上の注意事項について確認すること．

授業の e-learning 化は準備中である．完成後，速やかに導入する． Schedule⟩

大項目内容
1. 講義毎週火曜日 17:30∼ 19:00
2. 研究指導・技術指導随時

Evaluation Criteria⟩ 出席状況等，授業への取り組み姿勢にもとづく評価と，各教員による評価を併用して

行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217604 Contact⟩

⇒ Fukui (fukui@basic.med.tokushima-u.ac.jp) ⇒ Sakata-Haga (+81-88-633-7052, haga@basic.med.tokushima-u.ac.jp)

Note⟩ ・中枢神経系の発生異常に関して，分子レベルから個体レベルまで広い範囲で学ぶことができる．・ヒトとモデル動物の先天性神経疾患の発症機構，病態を理解することができる．・同一大口座の他分野等との人的交流を活発に行なう．

Seminar for Birth Defects 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshihiro Fukui⟨Manager⟩ · Professor / Course of Human Development, Medical Science Hiromi Sakata-Haga · Associate Professor / Course of Human Development, Medical Science

Target⟩ 中枢神経系の発生・発達障害の成立機構解明に関する実験技術を体得し，さらに先天異常への興味を

深くする． Outline⟩ 子宮内環境の影響を強く受ける胎児中枢神経系の発生異常と生後の学習障害や精神発達遅滞 (胎児性

アルコール症候群等) との関連性に関する神経解剖学的，分子生物学的解析法を学習する．遺伝性小脳性運動失調の発症機序を免疫組織化学的，分子生物学的に解析し，さらに発生，発達と関連した学術論文を抄読し，仮説や研究計画の立案のしかたおよびその実証方法について理解する．

Manner⟩ オムニバス方式 Notice⟩ 受講前に書く授業の担当教官と連絡を取り，受講上の注意事項について確認すること．

授業の e-learning 化は準備中である．完成後，速やかに導入する． Schedule⟩

大項目内容
1. 抄読会・リサーチカンファレンス ; 毎週火曜日 19:00∼ 20:30
2. 研究指導・技術指導随時

Evaluation Criteria⟩ 出席状況等，授業への取り組み姿勢にもとづく評価と，各教員による評価を併用して

行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217605 Contact⟩

⇒ Fukui (fukui@basic.med.tokushima-u.ac.jp) ⇒ Sakata-Haga (+81-88-633-7052, haga@basic.med.tokushima-u.ac.jp)

Note⟩ ・中枢神経系研究の基礎的手法の方法論を理解できる・発生学研究の方法論の理解も併せて行う．同一大口座の他分野等との人的交流を活発に行なう．

Developmental Pathophysiology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shoji Kagami⟨Manager⟩ · Professor / Course of Human Development, Medical Science

, Kenji Mori · Associate Professor / Course of Human Development, Medical Science Yasunobu Hayabuchi · Associate Professor / Course of Human Development, Medical Science

Hiroyoshi Watanabe · Associate Professor / Course of Human Development, Medical Science

, Shuji Kondo · Assistant Professor / Institute of Health Biosciences

Target⟩ To gain knowledge about pathophysiology, and advances in the diagnosis and treatment about

pediatric and developmental diseases. To learn how to perform basic and clinical research. Outline⟩ We would like to provide opportunities that our staff and attendances try to systematically understand

the diagnostic procedure, pathophysiology, treatment, and preventives of pediatric diseases including renal disease, nervous disease, heart disease, malignant disease, rheumatic diseases, etc. Furthermore, we would like to discuss experimental design, methods, technique, and results on research of molecular and cell biology for pediatric and developmental diseases.

Notice⟩ n/a

Schedule⟩

大項目担当

1∼4.

Guidance for the course and pediatric nephrology

Shoji Kagami

5∼8.

Pediatric neurology

Kenji Mori

9∼11.

Pediatric cardiology Yasunobu Hayabuchi

12∼14.

Pediatric hematology Hiroyoshi Watanabe

15∼16.

Pediatric rheumatology

Shuji Kondo

Evaluation Criteria⟩ n/a

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217602 Contact⟩

⇒ Kagami (kagami@clin.med.tokushima-u.ac.jp) (Office Hour: Tuesday 17:30-18:30) Note⟩ 臨床医学研究者，研究マインドをもった臨床医および臨床的知識をもった基礎医学研究者の育成を目

指す．

Practical developmental pathophysiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shoji Kagami⟨Manager⟩ · Professor / Course of Human Development, Medical Science

, Kenji Mori · Associate Professor / Course of Human Development, Medical Science Yasunobu Hayabuchi · Associate Professor / Course of Human Development, Medical Science

Hiroyoshi Watanabe · Associate Professor / Course of Human Development, Medical Science

, Shuji Kondo · Assistant Professor / Institute of Health Biosciences

Target⟩ To gain knowledge about pathophysiology, and advances in the diagnosis and treatment about

pediatric and developmental diseases. To learn how to perform basic and clinical research. Outline⟩ We would like to provide opportunities that our staff and attendances try to systematically understand

the diagnostic procedure, pathophysiology, treatment, and preventives of pediatric diseases including renal disease, nervous disease, heart disease, malignant disease, rheumatic diseases, etc. Furthermore, we would like to discuss experimental design, methods, technique, and results on research of molecular and cell biology for pediatric and developmental diseases.

Notice⟩ n/a

Schedule⟩

大項目担当

1∼4.

Guidance for the course and pediatric nephrology

Shoji Kagami

5∼8.

Pediatric neurology

Kenji Mori

9∼11.

Pediatric cardiology Yasunobu Hayabuchi

12∼14.

Pediatric hematology Hiroyoshi Watanabe

15∼16.

Pediatric rheumatology

Shuji Kondo

Evaluation Criteria⟩ n/a

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217603

Contact⟩

⇒ Kagami (kagami@clin.med.tokushima-u.ac.jp) (Office Hour: Thursday 19:00-20:00) Note⟩ 臨床医学研究者，研究マインドをもった臨床医および臨床的知識をもった基礎医学研究者の育成を目

指す．

Obstetrics and Gynecology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Minoru Irahara · Professor / Course of Human Development, Medical Science

, Hiroyuki Furumoto · Designated Professor / University Hospital

Toshiyuki Yasui · Professor / Institute of Health Biosciences , Toshiya Matsuzaki · Associate Professor / Course of Human Development, Medical Science Kazuhisa Maeda · Associate Professor / Course of Human Development, Medical Science Akira Kuwahara · Associate Professor / Course of Human Development, Medical Science Satoshi Yamamoto · Associate Professor / Course of Human Development, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217498

Contact⟩ ⇒ Irahara (+81-88-633-7177, irahara@clin.med.tokushima-u.ac.jp) ⇒ Furumoto (furumoto@clin.med.tokushima-u.ac.jp) ⇒ Yasui (yasui@clin.med.tokushima-u.ac.jp) ⇒ Matsuzaki (mat@clin.med.tokushima-u.ac.jp) ⇒ Maeda (maeda@clin.med.tokushima-u.ac.jp) ⇒ Kuwahara (akira@clin.med.tokushima-u.ac.jp) ⇒ Yamamoto .

Practice of Obstetrics and Gynecology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Minoru Irahara · Professor / Course of Human Development, Medical Science

, Hiroyuki Furumoto · Designated Professor / University Hospital

Toshiyuki Yasui · Professor / Institute of Health Biosciences , Toshiya Matsuzaki · Associate Professor / Course of Human Development, Medical Science Kazuhisa Maeda · Associate Professor / Course of Human Development, Medical Science Akira Kuwahara · Associate Professor / Course of Human Development, Medical Science Satoshi Yamamoto · Associate Professor / Course of Human Development, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217499

Contact⟩ ⇒ Irahara (+81-88-633-7177, irahara@clin.med.tokushima-u.ac.jp) ⇒ Furumoto (furumoto@clin.med.tokushima-u.ac.jp) ⇒ Yasui (yasui@clin.med.tokushima-u.ac.jp) ⇒ Matsuzaki (mat@clin.med.tokushima-u.ac.jp) ⇒ Maeda (maeda@clin.med.tokushima-u.ac.jp) ⇒ Kuwahara (akira@clin.med.tokushima-u.ac.jp) ⇒ Yamamoto .

発生発達医学実験実習 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshihiro Fukui · Professor / Course of Human Development, Medical Science , Shoji Kagami · Professor / Course of Human Development, Medical Science Minoru Irahara · Professor / Course of Human Development, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217534 Contact⟩

⇒ Fukui (fukui@basic.med.tokushima-u.ac.jp) ⇒ Kagami (kagami@clin.med.tokushima-u.ac.jp) ⇒ Irahara (+81-88-633-7177, irahara@clin.med.tokushima-u.ac.jp)

Physiology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazuo Yoshizaki⟨Manager⟩ · Professor / Course of Preventive Medicine, Medical Science

Target⟩ 生体の生理機能とエネルギー代謝について分子レベルからの解析方法を習得する． Outline⟩ 生体機能の非侵襲的計測法，細胞生理学的研究および体力医学・環境生理学的研究について，最新の

知見を含めセミナー形式で習学する． Notice⟩ セミナー形式毎週金曜日 9:30-11:00 e-Learning には対応していない． Schedule⟩

大項目担当
1. 授業ガイダンス吉崎
2∼15. 生理学セミナー〃

Evaluation Criteria⟩ 本授業はセミナー形式で行うが，出席状況や授業への取り組み姿勢などにもとづく評価

を期末に行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217524 Contact⟩

⇒ Yoshizaki (+81-88-633-7055, kyoshi@basic.med.tokushima-u.ac.jp) (Office Hour: 火・水曜日 17時 ∼ 18時 (e-mailにより時間調節を適宜おこないます))

Practice of Physiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazuo Yoshizaki⟨Manager⟩ · Professor / Course of Preventive Medicine, Medical Science

Target⟩ 生体の生理機能とエネルギー代謝について分子レベルからの解析方法を習得し，その応用研究が推進

できる． Outline⟩ 生体機能の非侵襲的計測法，細胞生理学的研究および体力医学・環境生理学的研究について代表的

学術論文・成書を輪読あるいは精読し，仮説や研究計画の立案のしかたおよびその実証方法について考案できる．

Notice⟩ 輪読会形式，後期後半 (火，水曜日午前) ほか随時， e-Learning には対応していない Schedule⟩

大項目担当
1. 授業ガイダンス吉崎
2∼15. 生理学カンファレンス〃

Evaluation Criteria⟩ 出席状況や授業への取り組み姿勢などにもとづく評価を期末に行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217525 Contact⟩

⇒ Yoshizaki (+81-88-633-7055, kyoshi@basic.med.tokushima-u.ac.jp) (Office Hour: 火・水曜日17時∼ 1 8時(e-mailにより時間調節を適宜おこないます))

Gastroenterology and Oncology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuji Takayama · Professor / Course of Preventive Medicine, Medical Science

Toshiya Okahisa · Associate Professor / Course of Preventive Medicine, Medical Science

, Naoki Muguruma · Associate Professor / University Hospital

Target⟩ 消化器および循環器に関する研究について，その方法論を修得させ，それに基づき各研究分野の先端

的研究を推進するよう指導する． Outline⟩ ポストゲノム時代の内視鏡学，肝発癌と肝線維化に関する分子生物学，肝炎・肝癌の診断と治療，新

しい内視鏡診断と治療，心臓病の新しい診断と治療法について講義する．オムニバス方式で行う． Manner⟩ オムニバス方式 Notice⟩ 授業はプリントやパワーポイントを使用するが，各講義は独立した内容なので，講義に集中し，ノー

トの取り方を工夫すること．遅刻の場合，「出席」扱いはしない． Schedule⟩

大項目担当
1. 赤外線蛍光内視鏡学清水一郎
2. 肝線維化と肝発癌清水一郎
3. 肝炎・肝癌の診断と治療本田浩仁
4. 新しい内視鏡診断と治療六車直樹
5. 心臓病の新しい診断と治療法添木武

Evaluation Criteria⟩ 授業は講義形式で行うが，出席状況，質疑応答といった授業への取組み姿勢などにもとづく平常点での評価を中心に行う．また，理解度の確認のために質問用紙を配布し，授業終了後に回収することも行う．

Re-evaluation⟩ 行わない Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217507 Contact⟩

⇒ Takayama (+81-88-633-7122, takayama@clin.med.tokushima-u.ac.jp)

⇒ Okahisa (okahisa@clin.med.tokushima-u.ac.jp)

⇒ Muguruma (+81-88-633-7124, muguruma@clin.med.tokushima-u.ac.jp) Note⟩ non medical の学生に消化器・循環器疾患に関する臨床医学を修得させる．

Practice of Gastroenterology and Oncology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuji Takayama · Professor / Course of Preventive Medicine, Medical Science

Toshiya Okahisa · Associate Professor / Course of Preventive Medicine, Medical Science

, Naoki Muguruma · Associate Professor / University Hospital

Target⟩ 臓器病態治療医学で修得した研究手法を実践し，各研究分野の先端的研究を推進するよう指導する． Outline⟩ 肝発癌と肝線維化に関する分子生物学，非観血的検査法による心機能評価と冠循環動態，難治性不

整脈の病態と治療および赤外線蛍光内視鏡を用いた新しい不可視病変の内視鏡診断と治療法について，代表的学術論文を学生に抄読させ，仮説や研究計画の立案の仕方，及びその実証方法について指導する．オムニバス方式で行う．

Manner⟩ オムニバス方式 Notice⟩ 各演習項目の目的，理論等を解説する．解説書やプリント類を前もって熟読する．十分に理解した上

で演習に取り組む．なお，遅刻の場合，「出席」扱いはしない． Schedule⟩

大項目担当
1. 赤外線蛍光内視鏡学臓器病態治療医学
2. 肝線維化と肝発癌〃
3. 心機能検査法〃
4. 不整脈〃

Evaluation Criteria⟩ 演習では，画像，内視鏡機器，心電図，心エコー図等を前に，質疑応答形式で行い，出席状況や授業への取り組む姿勢などにもとづく平常点での評価を中心に行う．また，理解度の確認のために質問用紙を配布し，演習終了後に回収することも行う．

Re-evaluation⟩ 行わない Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217508 Contact⟩

⇒ Takayama (+81-88-633-7122, takayama@clin.med.tokushima-u.ac.jp)

⇒ Okahisa (okahisa@clin.med.tokushima-u.ac.jp)

⇒ Muguruma (+81-88-633-7124, muguruma@clin.med.tokushima-u.ac.jp) Note⟩ non medical の学生に消化器・循環器疾患に関する臨床医学を修得させる．

Immune system development 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science Kensuke Takada · Associate Professor / Course of Preventive Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217547 Contact⟩

⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

⇒ Takada (+81-88-633-9472, takada@genome.tokushima-u.ac.jp)

Immune system development 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science Kensuke Takada · Associate Professor / Course of Preventive Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217548 Contact⟩

⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

⇒ Takada (+81-88-633-9472, takada@genome.tokushima-u.ac.jp)

Immune Regulation 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Il-mi Okazaki · Assistant Professor / Institute for Genome Research

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217549 Contact⟩

⇒ Okazaki (414, +81-88-633-9158, tokazaki@genome.tokushima-u.ac.jp)

⇒ Okazaki (+81-88-633-9158, ilmi@genome.tokushima-u.ac.jp)

Practice of Immune Regulation 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Il-mi Okazaki · Assistant Professor / Institute for Genome Research

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217550 Contact⟩

⇒ Okazaki (414, +81-88-633-9158, tokazaki@genome.tokushima-u.ac.jp)

⇒ Okazaki (+81-88-633-9158, ilmi@genome.tokushima-u.ac.jp)

Experimental practice for preventive medicine 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Koji Yasutomo⟨Manager⟩ · Professor / Course of Microbiology and Immunology, Medical Science

Kazuo Yoshizaki · Professor / Course of Preventive Medicine, Medical Science

, Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

Kenji Kishihara ·, Ichiro Shimizu ·, Shuhei Tomita · Associate Professor / Course of Biological Medicine, Medical Science

Hirohito Honda · Target⟩ 病態予防医学研究を遂行する上で必須の基本理論・方法を習得する． Outline⟩ 感染に対する生体防御機構，感染症の制圧に向けた宿主免疫系の操作， T リンパ球と胸腺器官の発生と再生をはじめとする免疫システムの発生構築，遺伝子改変モデル動物を用いた先天性免疫不全病の遺伝子治療，肝発癌と肝線維感化に関する分子生物学，超音波血流計測法・核磁気共鳴法の応用，非観血的検査法による心機能評価，難治性不整脈の診断，赤外線蛍光内視鏡を用いた新しい不可視病変の内視鏡診断について，実際に実験を行いながら技術を習得する．実験成績に基づいた作業仮説の立案と実証，修正のプロセスを学び，最終的に研究成果を学術論文にまとめる．

Notice⟩ e-learning には対応していない．

Schedule⟩

大項目内容
1. 抄読会毎週水・金曜日 9:00-10:30
2. 研究・技術指導随時

Evaluation Criteria⟩ 出席状況とレポート等で判定する．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217600

Contact⟩

⇒ Yasutomo (+81-88-633-7048, yasutomo@basic.med.tokushima-u.ac.jp) (Office Hour: 水曜日11時から

12時) ⇒ Yoshizaki (+81-88-633-7055, kyoshi@basic.med.tokushima-u.ac.jp) (Office Hour: 水曜日17時から18

時 (e-mailにより時間調節を適宜おこないます)) ⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

(Office Hour: 水曜日16時∼ 18時)

Immunology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Koji Yasutomo⟨Manager⟩ · Professor / Course of Microbiology and Immunology, Medical Science

, Kenji Kishihara ·

Yoichi Maekawa · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ 免疫システムを中心とした生体防御医学全般に関する基礎的理論を習得させる．また生体防御医学に

関する最新知見を解説する． Outline⟩ Basic concepts of immune system and recent advances of immunology. Manner⟩ オムニバス方式 Notice⟩ e-learning には対応していない．

Schedule⟩

大項目内容
1. リサーチカンファレンス毎週月曜日 9:30-11:00
2. 抄読会毎週水曜日 9:00-10:30
3. 研究・技術指導随時

Evaluation Criteria⟩ 出席状況とレポート等によって判定する．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217594

Contact⟩

⇒ Yasutomo (+81-88-633-7048, yasutomo@basic.med.tokushima-u.ac.jp) (Office Hour: 水曜日11時から 12時)

Practice for Immunology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Koji Yasutomo⟨Manager⟩ · Professor / Course of Microbiology and Immunology, Medical Science

, Kenji Kishihara ·

Yoichi Maekawa · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ 免疫システムを中心とした生体防御医学に関する基礎的方法論を習得させる．

Outline⟩ 生体防御医学に関する基礎的な概念・方法論についての代表的学術論文を学生に抄読させ，仮説や

研究計画の立案の方法およびその実証方法について指導する． (オムニバス方式) Manner⟩ オムニバス方式 Notice⟩ e-learning には対応していない

Schedule⟩

大項目内容
1. リサーチカンファレンス毎週月曜日 9:30-11:00
2. 抄読会毎週水曜日 9:00-10:30
3. 研究・技術指導随時

Evaluation Criteria⟩ 出席状況とレポート等によって判定する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217595 Contact⟩

⇒ Yasutomo (+81-88-633-7048, yasutomo@basic.med.tokushima-u.ac.jp) (Office Hour: 水曜日11時から 12時)

Microbiology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akio Adachi⟨Manager⟩ · Professor / Course of Microbiology and Immunology, Medical Science

Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics

Tsuneo Uchiyama · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ 主としてヒト免疫不全ウイルス (HIV) を対象に，ウイルス複製機序に関与する分子遺伝学的要因と感染に対する細胞応答について解析し，ウイルス学，免疫学，蛋白質化学，分子遺伝学などの主要な技法を習得させる．

Outline⟩ ウイルス複製過程の分子遺伝学的解析を中心に，細胞死を含めたウイルス感染に対する細胞応答，さらにウイルス病原性発現機構について，学生を指導する．これらにより，作業仮説や研究計画の立案の仕方およびその実証法について習得させる． (オムニバス方式)

Manner⟩ オムニバス方式 Notice⟩ e-learning 化はしていない．授業は英語で行われる． Schedule⟩

大項目内容担当
1∼15. 研究成果検討会毎週火曜日 16:15∼ 18:15  {足立昭夫，木戸博，内山 \atop 恒夫}
16. 研究指導・技術指導随時〃

Evaluation Criteria⟩ 出席点と学習態度などによる． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217536 Contact⟩

⇒ e-mailにて時間調節の上,面談する. Note⟩ 学習態度とは，発表準備状況，発表内容，質疑応答などをさす．

Practice for Microbiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akio Adachi⟨Manager⟩ · Professor / Course of Microbiology and Immunology, Medical Science Tsuneo Uchiyama · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ ウイルスの複製ならびに病原性の機序に関与するウイルス及び細胞要因について総合的に理解を与え

る．また，ウイルスを基盤とした技術の医学応用についても学習する． Outline⟩ ウイルス宿主細胞相互作用におけるウイルス複製とそれに対する細胞応答の機序について，それらを

規定する要因に関し最新の知見を含め体系的に学習する．さらに，ウイルス病の新しい診断，治療，予防法に関しても，また，ウイルスベクターの利用等ウイルスを用いた応用面などについて，包括的に学習する．

Manner⟩ オムニバス方式 Notice⟩ e-learning 化はしていない．授業は英語で行われる． Schedule⟩

大項目内容担当
1. 論文抄読会毎週木曜日 9:15∼ 10:15(通年受講)  足立昭夫，内山恒夫

Evaluation Criteria⟩ 出席点と学習態度などによる． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217537 Contact⟩

⇒ e-mailにて時間調節の上,面談する. Note⟩ 学習態度とは，発表準備状況，発表内容，質疑応答などをさす．

Practice for Microbiology and Immunology 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217535 Contact⟩

⇒ Yasutomo (+81-88-633-7048, yasutomo@basic.med.tokushima-u.ac.jp) ⇒ Adachi (+81-88-633-7078, adachi@basic.med.tokushima-u.ac.jp)

Preventive Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kokichi Arisawa⟨Manager⟩ · Professor / Course of Social and Environmental Medicine, Medical Science Hirokazu Uemura · Associate Professor / Course of Social and Environmental Medicine, Medical Science

Target⟩ 疫学の基本的事項を理解し，ヒト集団における健康問題を正しく評価して疾病予防に貢献できるため

の基礎知識を身につける． Outline⟩ ヒト集団における疾病の頻度を測定し，その規定因子を明らかにするための方法について解説する．

また，疫学研究の結果の妥当性を脅かす要因とその対策について解説する． Notice⟩ 現在のところ， e-learning に対応していません． Schedule⟩

大項目

1. 疾病頻度の測定
2. 関連の指標
3. 地域相関研究
4. コホート研究，介入研究
5. 患者対照研究，コホート内患者対照研究
6. 偏りおよび交絡とその制御
7. 層別解析 (マンテルヘンツエル推定量)
8. 標準化
9. マッチング
10. 偶然誤差と系統誤差
11. ROC 分析
12. 疫学研究の倫理指針
13. 疫学研究のデザイン (1)
14. 疫学研究のデザイン (2)
15. 疫学研究のデザイン (3)

Evaluation Criteria⟩ 筆記試験および課題についてのレポートを総合して行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217579 Contact⟩

⇒ 授業担当者とメールで相談の上決めて下さい．

Practice of Preventive Medicine 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kokichi Arisawa⟨Manager⟩ · Professor / Course of Social and Environmental Medicine, Medical Science Hirokazu Uemura · Associate Professor / Course of Social and Environmental Medicine, Medical Science

Target⟩ 疫学論文を情報データベースから検索し，その質を批判的に吟味する (critical appraisal) 能力を身に

つける． Outline⟩ 疫学論文を一定の形式 (別紙) に従って批判的に吟味する．題材の半分は教員が指定し，残りの半分

は学生が選択する．レポート提出とそれについての討論により，疫学研究論文を正しく評価できる能力を養成する．

Notice⟩ 現在のところ， e-learning に対応していません． Schedule⟩

大項目

1∼8. 論文の批判的吟味 (題材は教官が選択)
9∼15. 論文の批判的吟味 (題材は受講者が選択)

Evaluation Criteria⟩ 成績評価の方法レポートをもとに行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217580 Contact⟩

⇒ 授業担当者とメールで相談の上決めて下さい．

Community Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kenji Tani · Designated Professor / Course of Social and Environmental Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217528 Contact⟩

⇒ Tani (+81-88-633-7127, kenjikt@clin.med.tokushima-u.ac.jp) (Office Hour: 9am-5:30pm)

Practice of Community Medicine 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kenji Tani · Designated Professor / Course of Social and Environmental Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217529 Contact⟩

⇒ Tani (+81-88-633-7127, kenjikt@clin.med.tokushima-u.ac.jp) (Office Hour: 9am-5:30pm)

Practice of Social and Environmental Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kokichi Arisawa⟨Manager⟩ · Professor / Course of Social and Environmental Medicine, Medical Science

Target⟩ 疫学データを解析し，論文にまとめ，医学雑誌に受理されるための技術を身につける． Outline⟩ 実際の疫学データ (臨床データを含む) を統計パッケージ (SAS， SPSS) で解析し，記述統計，単変量

解析，多変量解析の順に結果をまとめる．題材は，前半は教員が選択し，後半は受講者が選択する．最終的な結果を研究論文としてまとめる方法について指導する．

Notice⟩ 現在のところ， e-learning に対応していません． Schedule⟩

大項目

1. 記述統計

2.

平均値の差の検定， Wilcoxon 順位和検定

3. カイ 2 乗検定， Fisher の直接確率法
4. 相関，回帰，重回帰分析
5. 生存分析 (カプランマイヤー法，ログランク検定)
6. 生存分析 (Cox 比例ハザードモデル)
7. 多重ロジステイック回帰解析
8. ポアソン回帰分析
9∼11. SAS 入門

12∼15.疫学データ解析 (題材は受講者が選択)

Evaluation Criteria⟩ 成績評価の方法レポートをもとに行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217606 Contact⟩

⇒ 授業担当者とメールで相談の上決めて下さい．

Anatomy and Cell Biology 2 units (selection)

1st-year(1st, 2nd semester), 2nd-year(1st, 2nd semester), 3rd-year(1st, 2nd semester)

Kazunori Ishimura⟨Manager⟩ · Professor / Course of Integrated Brain Sciences, Medical Science Michihiro Nakamura · Associate Professor / Course of Integrated Brain Sciences, Medical Science

Target⟩ 形態学的方法論を理解し，これに基づいて主として神経伝達機構を対象とした基本的および先端的形

態学的研究手技を身につける．ナノバイオテクノロジーの形態学的方法への応用についても検討する． Outline⟩ 生体の統合性の維持における脂質性情報伝達物質の意義を，中枢神経系を中心として，それらの合

成酵素とレセプターの局在の組織化学的，微細構造的視点から解説，講義する．また新規なナノバイオテクノロジーの開発とその形態学的への応用についても解説する．

Notice⟩ 特になし．一部 e-learning 化を検討している． Schedule⟩

大項目担当
1. 授業ガイダンス，顕微解剖学概論石村和敬
2. 細胞構造学 (1)  〃
3. 細胞構造学 (2)  〃
4. 光学顕微鏡法 (1)  〃
5. 光学顕微鏡法 (2)  〃
6. 神経組織の形態学 (1)  樋田一徳
7. 神経組織の形態学 (2)  〃
8. 共焦点レーザー顕微鏡法 (1)  〃
9. 共焦点レーザー顕微鏡法 (2)  〃
10. 共焦点レーザー顕微鏡お法 (3)  〃
11. 免疫組織化学法 (1)  中村教泰
12. 免疫組織化学法 (2)  〃
13. 免疫組織化学法 (3)  〃
14. 電子顕微鏡法 (1)  〃
15. 電子顕微鏡法 (2)  〃

Evaluation Criteria⟩ 出席状況とレポート，口頭試問等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217612 Contact⟩

⇒ Ishimura (+81-88-633-7049, ishimura@basic.med.tokushima-u.ac.jp) (Office Hour: 金曜日の16:30∼ 1

8:00(e-mailにより適宜時間調整を行います)) ⇒ 他の教員については， e-mailで時間調整の上面談してください．

Practice of Anatomy and Cell Biology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazunori Ishimura⟨Manager⟩ · Professor / Course of Integrated Brain Sciences, Medical Science Michihiro Nakamura · Associate Professor / Course of Integrated Brain Sciences, Medical Science

Target⟩ 神経伝達機構の形態学的研究方法を習得する． Outline⟩ ステロイドホルモンやプロスタグランジンなどの脂質性情報伝達物質の合成酵素やレセプターの中

枢神経系における局在，ナノバイオテクノロジーを用いた研究に関する代表的学術論文を学生に抄読させ，仮説や研究計画の立案の仕方及びその実証方法について指導する．

Notice⟩ 論文のまとめ方，発表態度，考察のレベルなどを総合して判定する． Schedule⟩

大項目担当
1. 論文検討会 (1)  石村和敬
2. 論文検討会 (2)  〃
3. 論文検討会 (3)  〃
4. 論文検討会 (4)  〃
5. 論文検討会 (5)  〃
6. 論文検討会 (6)  樋田一徳
7. 論文検討会 (7)  〃
8. 論文検討会 (8)  〃
9. 論文検討会 (9)  〃
10. 論文検討会 (10)  〃
11. 論文検討会 (11)  中村教泰
12. 論文検討会 (12)  〃
13. 論文検討会 (13)  〃
14. 論文検討会 (14)  〃
15. 論文検討会 (15)  〃

Evaluation Criteria⟩ 特になし． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217613 Contact⟩

⇒ Ishimura (+81-88-633-7049, ishimura@basic.med.tokushima-u.ac.jp) (Office Hour: 金曜日の16:30∼ 1

8:00(e-mailにより適宜時間調整を行います)) ⇒ 他の教員については， e-mailで時間調整の上面談してください．

Note⟩ なし

Integrative Physiology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217530 Contact⟩

⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp)

Practice of Integrative Physiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science , Tetsuya Shiuchi · Associate Professor / Institute of Health Biosciences Hiroaki Fujihara · Assistant Professor / Institute of Health Biosciences Sachiko Chikahisa · Assistant Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217531 Contact⟩

⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp)

⇒ Shiuchi (+81-88-633-7057, shiuchi@basic.med.tokushima-u.ac.jp)

⇒ Fujihara .

⇒ Chikahisa (+81-88-633-7057, chika@basic.med.tokushima-u.ac.jp)

Psychiatry 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217518 Contact⟩

⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour: 8:30-17:30)

Practical Psychiatry 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217519 Contact⟩

⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour: 8:30-17:30)

Brain Pathophysiology and Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shinji Nagahiro · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217532 Contact⟩

⇒ Nagahiro (+81-88-633-7148, snagahi@clin.med.tokushima-u.ac.jp)

Practice of Brain Pathophysiology and Medicine 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shinji Nagahiro · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217533 Contact⟩

⇒ Nagahiro (+81-88-633-7148, snagahi@clin.med.tokushima-u.ac.jp)

Strategic medical informatics 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217527 Contact⟩

⇒ Moriguchi (+81-88-633-9410, h moriguchi@ap6.mopera.ne.jp)

Practice of strategic medical informatics 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217526 Contact⟩

⇒ Moriguchi (+81-88-633-9410, h moriguchi@ap6.mopera.ne.jp)

Molecular Biology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217520 Contact⟩

⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp)

Practice of Molecular Biology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217521 Contact⟩

⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp)

Practice for Integrated Brain Sciences and Medical Information 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazunori Ishimura · Professor / Course of Integrated Brain Sciences, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science , Shinji Nagahiro · Professor / Course of Integrated Brain Sciences, Medical Science Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217509 Contact⟩

⇒ Ishimura (+81-88-633-7049, ishimura@basic.med.tokushima-u.ac.jp) ⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp) ⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour:

8:30-17:30) ⇒ Nagahiro (+81-88-633-7148, snagahi@clin.med.tokushima-u.ac.jp) ⇒ Moriguchi (+81-88-633-9410, h moriguchi@ap6.mopera.ne.jp) ⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp)

Pharmacology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshiaki Tamaki⟨Manager⟩ · Professor / Course of Biological Medicine, Medical Science Kazuo Minakuchi⟨Manager⟩ · Professor / Institute of Health Biosciences

Target⟩ 薬理学的研究に関する方法論を習得させ，循環器疾患を対象とした研究が推進できるように指導する． Outline⟩ 循環調節機序に関する細胞間および細胞内情報伝達機構とその異常により発症する病態について解

説するとともに，循環器疾患に対する新しい予防法や薬物療法さらに創薬について体系的に講義する． Manner⟩ オムニバス方式 Notice⟩ 全ての講義に参加のこと．未だ， e-learning 化した講義は利用できない． Schedule⟩

大項目担当
1. 授業ガイダンス玉置俊晃
2∼3. 生体内情報伝達と細胞内情報伝達機構〃

4∼5.

Pharmacodynamics & Pharmacokinetics 〃

6∼7. 循環器疾患とその治療薬〃
8∼9. 腎疾患とその治療薬〃

10∼11.循環器疾患に関する薬物療法水口和生

12∼13.分子生物学的実験方法森田恭二

14∼15.細胞培養法と培養細胞を用いた実験方法〃

Evaluation Criteria⟩ レポートの提出と研究計画立案書の提出 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217551 Contact⟩

⇒ e-mail にて時間調節の上,面談して下さい. Note⟩ 情報伝達薬理学分野以外の共同研究者とも交流を活発に行い指導を受ける．

Practice for Pharmacology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshiaki Tamaki⟨Manager⟩ · Professor / Course of Biological Medicine, Medical Science Kazuo Minakuchi⟨Manager⟩ · Professor / Institute of Health Biosciences

Target⟩ 情報伝達機構とその異常により発症する病態を理解し，薬理学的研究を計画出来るように指導する． Outline⟩ 循環調節機序に関する細胞間および細胞内情報伝達機構とその異常により発症する病態，循環器疾

患に対する新しい予防法や薬物療法について代表的な学術論文を学生に抄読させ，仮説や研究計画の立案の仕方及びその実証方法について指導する．

Manner⟩ オムニバス方式 Notice⟩ 全ての演習に参加のこと．演習のため， e-learning は利用できない． Schedule⟩

大項目担当
1. 授業ガイダンス玉置俊晃
2∼3. 細胞内情報伝達機構に関する抄読会〃
4∼5. 循環器疾患に関する抄読会〃
6∼7. 研究計画の立案の仕方及びその実証方法〃
8∼9. 実験結果のまとめ方・統計処理方法〃

10∼11.薬物療法に関する抄読会水口和生

12∼13.英語論文の投稿方法および検索方法森田恭二

14∼15.英語論文の書き方〃

Evaluation Criteria⟩ レポートの提出と研究計画立案書の提出 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217552 Contact⟩

⇒ e-mail にて時間調節の上,面談して下さい. Note⟩ 情報伝達薬理学分野以外の共同研究者とも交流を活発に行い指導を受ける．

Anesthesiology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shuzo Oshita · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217545 Contact⟩

⇒ Oshita (+81-88-633-7181, shuzo@clin.med.tokushima-u.ac.jp)

Practice of Anesthesiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Shuzo Oshita · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217546 Contact⟩

⇒ Oshita (+81-88-633-7181, shuzo@clin.med.tokushima-u.ac.jp)

Nephrology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshio Doi · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217512 Contact⟩

⇒ Doi (+81-88-633-7183, doi@clin.med.tokushima-u.ac.jp)

Seminar for Nephrology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshio Doi · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217513 Contact⟩

⇒ Doi (+81-88-633-7183, doi@clin.med.tokushima-u.ac.jp)

救急集中治療医学 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masaji Nishimura · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217487 Contact⟩

⇒ Nishimura (+81-88-633-9425, )

救急集中治療医学演習 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masaji Nishimura · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217488 Contact⟩

⇒ Nishimura (+81-88-633-9425, )

Experimental Practice for Biological Medicine 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshiaki Tamaki · Professor / Course of Biological Medicine, Medical Science

, Shuzo Oshita · Professor / Course of Biological Medicine, Medical Science Masaji Nishimura · Professor / Course of Biological Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217540 Contact⟩ ⇒ Tamaki (+81-88-633-7061, tamaki@basic.med.tokushima-u.ac.jp) ⇒ Oshita (+81-88-633-7181, shuzo@clin.med.tokushima-u.ac.jp) ⇒ Nishimura (+81-88-633-9425, )

Ophthalmology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshinori Mitamura · Professor / Course of Sensory Neuroscience, Medical Science

, Hiroshi Eguchi · Associate Professor / University Hospital

Target⟩ 視覚疾患に対する病態を理解し，最新の方法論を習得すると共に，先端的治療研究ができるように学

習する． Outline⟩ 感覚器感染症としての眼感染症の病態，遺伝子診断法，最新の治療法について体系的に講義する．感

覚器障害としての糖尿病網膜症の病態と治療法についても解説する． Manner⟩ オムニバス方式 Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目内容
1. 臨床カンファレンス毎週水曜日 16:00∼ 17:30
2. 抄読会毎週水曜日 17:30∼ 19:00
3. 研究指導・技術指導随時
4. {視覚病態学分野に関する高いレベルでの知識・技 \atop 術が習得できる．}

5. {同一大講座の他分野，他大学などとの人的交流を \atop 活発におこなう．}
6. 国際的見地からの思考力を養う．

Evaluation Criteria⟩ 出席状況とテストまたはレポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217480 Contact⟩

⇒ Mitamura (+81-88-633-7162, ymita@clin.med.tokushima-u.ac.jp) ⇒ Eguchi (8th Lab., +81-88-633-7163, hiroegu@clin.med.tokushima-u.ac.jp)

Practice of Ophthalmology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshinori Mitamura · Professor / Course of Sensory Neuroscience, Medical Science

, Hiroshi Eguchi · Associate Professor / University Hospital

Target⟩ 視覚疾患に対する病態を理解し，最新の方法論を習得すると共に，先端的治療研究ができるように学

習する． Outline⟩ 感覚器疾患として，眼感染症，糖尿病網膜症などに関する代表的学術論文を抄読し，作業仮説や研究

計画の立案とその実証方法について指導する． Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目内容
1. リサーチカンファレンス毎週月曜日 16:00∼ 17:30
2. 抄読会毎週木曜日 17:00∼ 18:30
3. 研究指導・技術指導随時
4. {視覚病態学分野に関する高いレベルでの知識・技 \atop 術が習得できる．}

5. {同一大講座の他分野，他大学などとの人的交流を \atop 活発におこなう．}

6. {国際的見地からの思考力を養い，将来留学できる \atop 実力を養う．}

Evaluation Criteria⟩ 出席状況とテストまたはレポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217481 Contact⟩

⇒ Mitamura (+81-88-633-7162, ymita@clin.med.tokushima-u.ac.jp) ⇒ Eguchi (8th Lab., +81-88-633-7163, hiroegu@clin.med.tokushima-u.ac.jp)

Otorhinolaryngology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Noriaki Takeda⟨Manager⟩ · Professor / Course of Sensory Neuroscience, Medical Science Koichi Tamura · Associate Professor / Course of Sensory Neuroscience, Medical Science

Target⟩ 感覚情報の統合と感覚・運動連関を統合的に理解，習得する目的で，神経科学に関する最新の知見，

トピックス，研究方法について講義を行う． Outline⟩ 感覚情報の中枢神経系におけるプロセッシングについて，特に聴覚や平衡覚に関して体系的に講義す

る．感覚・運動連関の出力としての顔面神経や喉頭神経の機能について，音声言語医学を含めて講義する． Manner⟩ オムニバス方式 Notice⟩ 授業の一部は e-learining 化しているので， e-learining 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 授業ガイダンス，聴覚言語医学序論武田憲昭
2∼3. 聴覚・平衡覚の受容メカニズム〃
4. 遺伝子難聴〃
5∼6. 聴覚と言語の獲得〃
7∼8. 聴覚の中枢プロセッシング〃
9∼10. 平衡覚の中枢プロセッシング〃
11. 空間識〃

12∼13.発声のメカニズム田村公一

14. 音声言語コミュニケーション医学〃
15. 顔面神経の障害と再生〃

Evaluation Criteria⟩ 口頭試問により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217607 Contact⟩

⇒ e-mailにてアポイントを取ってから,面談して下さい. Note⟩ 大講座の他分野の教官も講義をする他，内外の著名な講師を招いて講義を行う．

Experimental Practice for Otorhinolaryngology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Noriaki Takeda⟨Manager⟩ · Professor / Course of Sensory Neuroscience, Medical Science Koichi Tamura · Associate Professor / Course of Sensory Neuroscience, Medical Science

Target⟩ 感覚情報の統合と感覚・運動連関を総合的に理解，習得する目的で，神経科学に関する研究デザイン

の演習を行う． Outline⟩ 感覚情報の中枢神経系におけるプロセッシング，感覚・運動連関の出力としての顔面神経や喉頭神

経の機能，音声言語コミュニケーション医学に関する代表的学術論文を学生に抄読させ，作業仮説や研究計画の立案とその実証方法について指導する．

Manner⟩ オムニバス方式 Notice⟩ 授業の一部は e-learining 化しているので， e-learining 学習も出席として取り扱う． Schedule⟩

大項目中項目
1. 人工内耳
2. 人工中耳
3. 好酸球性中耳炎
4. 耳硬化症
5. 遺伝子難聴
6. メニエール病
7. 良性発作性頭位めまい症
8. アレルギー性鼻炎
9. 好酸球性副鼻腔炎
10. 顎顔面骨折
11. 上咽頭癌
12. 頭頸部領域のウイルス感染 (担当田村公一)
13. 睡眠時無呼吸症候群
14. 喉頭癌
15. 甲状腺癌

Evaluation Criteria⟩ 口頭試問により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217608 Contact⟩

⇒ e-mailにてアポイントを取ってから,面談して下さい. Note⟩ 大講座の他分野の教官や他施設の講師による研究指導も行う．

Clinical Neuroscience 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217556 Contact⟩

⇒ Kaji (rkaji@clin.med.tokushima-u.ac.jp)

Clinical Neuroscience Workshop 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217557 Contact⟩

⇒ Kaji (rkaji@clin.med.tokushima-u.ac.jp)

Sensory Neuroscience Laboratory Experiment 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Noriaki Takeda⟨Manager⟩ · Professor / Course of Sensory Neuroscience, Medical Science

, Hiroshi Shiota ·

Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

, Koichi Tamura · Associate Professor / Course of Sensory Neuroscience, Medical Science

Takeshi Naito · Associate Professor / Course of Sensory Neuroscience, Medical Science , Yuishin Izumi · 診療支援医師 Target⟩ 感覚情報の統合と感覚・運動関連を研究する上で，必要な基本理論と実験方法を習得することを目的

とする． Outline⟩ 感覚器疾患としての眼感染症と糖尿病網膜症の分子生物学的診断法と新しい治療法の開発，聴覚，平

衡覚の中枢神経系内でのプロセッシングの解析，感覚・運動連関の出力としての顔面神経障害，音声言語障害の新しい診断法，治療の開発，不随意運動を含めた運動障害とその診断法，治療法，予防法および脳神経機能解析法の開発について，実際に実験を行いながら技能を修得させる．また，実験成績に基づいた作業仮説の立案と実証，修正のプロセスを学ばせ，最終的に研究成果を学術論文にまとめる方法を指導する．

Notice⟩ 授業の一部は e-learining 化しているので， e-learining 学習も出席として取り扱う． Schedule⟩

大項目

1∼3. 角膜感染症に対する新しい診断法と治療法の開発
4∼5. 糖尿病網膜症に対する増悪因子とその遺伝子解明
6∼8. 平衡覚の中枢プロセッシングに関する分子生物学的実験
9∼10. 顔面神経の再生に関する電気生理学的実験

11∼13.神経変性疾患における遺伝子発現の異常の解析法に関する実験

14∼15.基底核疾患における感覚運動連関の評価法に関する実験

Evaluation Criteria⟩ 口頭試問により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217589 Contact⟩

⇒ e-mailにてアポイントを取ってから,面談して下さい. Note⟩ 大講座の他分野の教官も講義をする他，内外の著名な講師を招いて講義を行う

Human Pathology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Naomi Sakashita · Professor / Institute of Health Biosciences

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217485 Contact⟩

⇒ Sakashita .

Practice of Human Pathology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Naomi Sakashita · Professor / Institute of Health Biosciences

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217486 Contact⟩

⇒ Sakashita .

Gastroenterological Surgery and Transplantation 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuo Shimada · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Toru Utsunomiya · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Nobuhiro Kurita · Designated Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Takashi Iwata · Associate Professor / Institute of Health Biosciences

Masanori Nishioka · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science Yuji Morine · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science Satoru Imura · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217505 Contact⟩ ⇒ Shimada (+81-88-633-7137, mshimada@clin.med.tokushima-u.ac.jp) ⇒ Utsunomiya . ⇒ Kurita (kurita@clin.med.tokushima-u.ac.jp) ⇒ Iwata . ⇒ Nishioka (mnishioka@clin.med.tokushima-u.ac.jp) ⇒ Morine (ymorine@clin.med.tokushima-u.ac.jp) ⇒ Imura (s-imura@clin.med.tokushima-u.ac.jp)

Practice of Gastroenterological Surgery and Transplantation 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuo Shimada · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Toru Utsunomiya · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Nobuhiro Kurita · Designated Professor / Course of Diseased Organ and Repair Medicine, Medical Science Takashi Iwata · Associate Professor / Institute of Health Biosciences Masanori Nishioka · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Yuji Morine · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science Satoru Imura · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217506 Contact⟩

⇒ Shimada (+81-88-633-7137, mshimada@clin.med.tokushima-u.ac.jp) ⇒ Utsunomiya . ⇒ Kurita (kurita@clin.med.tokushima-u.ac.jp) ⇒ Iwata . ⇒ Nishioka (mnishioka@clin.med.tokushima-u.ac.jp) ⇒ Morine (ymorine@clin.med.tokushima-u.ac.jp) ⇒ Imura (s-imura@clin.med.tokushima-u.ac.jp)

Cardiovascular Surgery 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuya Kitagawa⟨Manager⟩ · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Takashi Kitaichi · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Target⟩ 循環器領域における臓器保護，再生・修復医学，組織移植医療の意義と展望についての理解を通して，

先端的研究の推進が可能となるよう指導する． Outline⟩ 循環器領域，特に先天性心疾患における再生医療，特に肺動脈を中心とした血管新生療法について，

最新の知見を解説，講義する． Notice⟩ 遅刻しないこと． e-learning 化は検討中．

毎週月曜日 7:45 リサーチカンファレンスを行っています． Schedule⟩

大項目中項目担当

1.

{授業ガイダンス+臓器保護, 組織移植医療, 再生・修 \atop 復医学序論} Tetsuya Kitagawa

2. 臓器保護概論〃

3.

臓器保護各論 1 心筋保護の基礎

Takashi Kitaichi

4.

臓器保護各論 2 心筋保護の実際

Tetsuya Kitagawa

5. 再生医療概論〃
6. 組織移植各論 1  先天性心疾患における再生医療開発〃
7. 再生・修復医学〃

8.

〃

Takashi Kitaichi

9.  〃〃
10. 組織修復学・医療の将来〃

Evaluation Criteria⟩ 知識・理解が十分得られ，新たな方法論を考察できる． Re-evaluation⟩ 行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217510 Contact⟩

⇒ Kitagawa (心臓血管外科教授室, +81-88-633-7151, kitagawa@clin.med.tokushima-u.ac.jp) (Office

Hour: 木曜日の14:00-17:00． ) ⇒ Kitaichi (+81-88-633-7152, kitaichi@clin.med.tokushima-u.ac.jp) (Office Hour: 木曜日の15:00-17:00．

)

Practice of Cardiovascular Surgery 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuya Kitagawa⟨Manager⟩ · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Takashi Kitaichi · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Target⟩ 課題研究に関する代表的な学術論文，最新の学術論文を抄読させることにより，研究計画の立案の仕

方，実証的方法を学ばせる． Outline⟩ 循環器領域，特に先天性心疾患・弁膜症外科における組織移植医療，組織工学，血管新生を中心とし

た再生医療等について，代表的な学術論文を学生に抄読させ，研究計画の立案の仕方およびその，実証方法について指導する．

Schedule⟩

大項目中項目内容
1. 公開抄読会リサーチカンファレンス毎週月曜 7:45 医局

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217511 Contact⟩

⇒ Kitagawa (心臓血管外科教授室, +81-88-633-7151, kitagawa@clin.med.tokushima-u.ac.jp) (Office

Hour: 16:00-17:00 every Tuesday) ⇒ Kitaichi (+81-88-633-7152, kitaichi@clin.med.tokushima-u.ac.jp)

Note⟩ 同一講座の他分野等の知的交流を活発にする．

Urology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiro-omi Kanayama · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Tomoharu Fukumori · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Masayuki Takahashi · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Hiroyoshi Nakatsuji · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Target⟩ 泌尿器科学的研究に関する方法論を習得させ，それに基づいて尿路性器癌の基礎研究と臨床応用に関

する先端的研究，排尿障害，泌尿器科の低侵襲手術について解説，指導する． Outline⟩ 尿路性器癌における浸潤，転移の分子生物学機構及びそれを標的分子とした治療への応用について

と泌尿器科腫瘍に対する低侵襲手術について解説，講義する．また，男性・女性での排尿障害の特徴と実際の治療法について解説する．

Notice⟩ 特になし Schedule⟩

大項目内容担当
1. 抄読会毎週月曜日，第 2， 4 水曜日 7:30∼ 8:30
2. 臨床カンファレンス毎週火曜日 15:00∼ 18:00
3. 泌尿器科腫瘍に対する低侵襲手術金山
4. 排尿障害西谷
5. {尿路性器癌に対する分子生物学的アプロー \atop チ}福森

Evaluation Criteria⟩ 実習の取り組みをみて総合的に判断する

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217585

Contact⟩ ⇒ Kanayama (+81-88-633-7158, kanahiro@clin.med.tokushima-u.ac.jp) ⇒ Fukumori (+81-88-633-7159, fukumori@clin.med.tokushima-u.ac.jp) ⇒ Takahashi (+81-88-633-7159, matakaha@clin.med.tokushima-u.ac.jp) ⇒ Nakatsuji (hnakatsu@clin.med.tokushima-u.ac.jp)

Note⟩ 特になし

Experimental Practice for Urology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiro-omi Kanayama⟨Manager⟩ · Professor / Course of Diseased Organ and Repair Medicine, Medical Science Tomoharu Fukumori · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Masayuki Takahashi · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science Hiroyoshi Nakatsuji · Assistant Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Target⟩ 泌尿器科学的研究に関する方法論を習得させ，それに基づいて尿路性器癌の基礎研究と臨床応用に関

する先端的研究，排尿障害，泌尿器科の低侵襲手術について解説，指導する． Outline⟩ 尿路性器癌における浸潤，転移の分子生物学機構及びそれを標的分子とした治療への応用について

と泌尿器科腫瘍に対する低侵襲手術について解説，講義する．また，男性・女性での排尿障害の特徴と実際の治療法について解説する．

Notice⟩ 特になし

Schedule⟩

大項目内容担当
1. 抄読会毎週月曜日，第 2， 4 水曜日 7:30∼ 8:30
2. 臨床カンファレンス毎週火曜日 15:00∼ 18:00
3. 泌尿器科腫瘍に対する低侵襲手術金山
4. 排尿障害西谷
5. {尿路性器癌に対する分子生物学的アプロー \atop チ}福森

Evaluation Criteria⟩ 実習の取り組みをみて総合的に判断する

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217586

Contact⟩ ⇒ Kanayama (+81-88-633-7158, kanahiro@clin.med.tokushima-u.ac.jp) ⇒ Fukumori (+81-88-633-7159, fukumori@clin.med.tokushima-u.ac.jp) ⇒ Takahashi (+81-88-633-7159, matakaha@clin.med.tokushima-u.ac.jp) ⇒ Nakatsuji (hnakatsu@clin.med.tokushima-u.ac.jp)

Note⟩ 特になし

Cardiovascular Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masataka Sata · Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217503

Contact⟩ ⇒ Sata (3F, +81-88-633-7851, sata@clin.med.tokushima-u.ac.jp) (Office Hour: 8:30-17:30)

Cardiovascular Medicine Practice 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masataka Sata · Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217504 Contact⟩

⇒ Sata (3F, +81-88-633-7851, sata@clin.med.tokushima-u.ac.jp) (Office Hour: 8:30-17:30)

Laboratory Experiment Practice of Diseased Organ and Repair Medicine

12 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Tetsuya Kitagawa⟨Manager⟩ · Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Naomi Sakashita · Professor / Institute of Health Biosciences

, Mitsuo Shimada · Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Hiro-omi Kanayama · Professor / Course of Diseased Organ and Repair Medicine, Medical Science

, Hiroki Ishibashi ·

Takashi Kitaichi · Associate Professor / Course of Diseased Organ and Repair Medicine, Medical Science

Target⟩ 課題研究に関して，作成した仮説を立証するための実験を行いながら，技能を習得させる．そして研

究成果を学術論文にまとめる． Outline⟩ 内分泌腫瘍細胞における機能と形態の相関性に関する分子病理学的解析，肝移植における免疫寛容

と熱ショック蛋白の非侵襲性誘導剤による肝保護，心・血管領域における血管新生を中心とした再生医療，また，実験成績に基づいた作業仮説の立案と実証，修正のプロセスを学ばせ，最終的に研究成果を学術論文にまとめる方法を指導する．

Schedule⟩

大項目

1. 実験実習

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217584 Contact⟩

⇒ Kitagawa (心臓血管外科教授室, +81-88-633-7151, kitagawa@clin.med.tokushima-u.ac.jp) (Office

Hour: 16:00-17:00 every Tuesday) ⇒ Sakashita . ⇒ Shimada (+81-88-633-7137, mshimada@clin.med.tokushima-u.ac.jp) ⇒ Kanayama (+81-88-633-7158, kanahiro@clin.med.tokushima-u.ac.jp) ⇒ Ishibashi . ⇒ Kitaichi (+81-88-633-7152, kitaichi@clin.med.tokushima-u.ac.jp)

Note⟩ 同一大講座の他分野等の器官・組織の再生・修復医学領域に関する研究技能を併せて修得する．

Experimental Oncology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science Hisanori Uehara · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217500 Contact⟩

⇒ Izumi (+81-88-633-7065, izumi@basic.med.tokushima-u.ac.jp) ⇒ Uehara (+81-88-633-7066, uehara@basic.med.tokushima-u.ac.jp)

Practice for Experimental Oncology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science Hisanori Uehara · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217501 Contact⟩

⇒ Izumi (+81-88-633-7065, izumi@basic.med.tokushima-u.ac.jp) ⇒ Uehara (+81-88-633-7066, uehara@basic.med.tokushima-u.ac.jp)

Radiology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masafumi Harada · Professor / Institute of Health Biosciences

, Hideki Otsuka · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Hitoshi Ikushima · Professor / Institute of Health Biosciences

, Junji Ueno · Professor / Institute of Health Biosciences

Kenji Matsuzaki · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Mayumi Takeuchi · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Target⟩ 放射線診断学と放射線腫瘍学の研究に関する方法論を習得させ，それに基づいて病態放射線医学の先

端的研究を推進するよう指導する Outline⟩ 放射線画像診断学と放射線腫瘍学について体系的に講義を行う．特に，画像診断学では 3 次元画像

による立体的な形態診断と核医学および磁気共鳴スペクトロスコピーによる機能診断について解説するとともに，放射線腫瘍学では放射線治療法および腫瘍の集学的治療法に関して解説する．

Manner⟩ オムニバス方式 Notice⟩ あらかじめ授業内容に関連した文献を読んでおき，疑問点について積極的に発言し議論に参加する

こと． Schedule⟩

大項目担当
1. 画像診断学 (3 次元画像の基礎理論)  西谷
2. 画像診断学 (3 次元画像の臨床応用)  上野
3. 画像診断学 (核医学による機能診断の基礎理論)  大塚
4. 画像診断学 (核医学による機能診断の臨床応用)  〃
5. 画像診断学 (磁気共鳴スペクトロスコピーによる機能診断の基礎理論) 原田
6. 画像診断学 (磁気共鳴スペクトロスコピーによる機能診断の臨床応用) 〃
7. 画像診断学 (形態と機能の総合的画像診断 1. 歴史と現況)  西谷
8. 画像診断学 (形態と機能の総合的画像診断 2. 将来展望)  〃
9. 放射線生物学生島
10. 高エネルギー放射線治療システム〃
11. 外部放射線治療の基礎〃
12. 外部放射線治療の臨床〃
13. 密封小線源治療の基礎〃
14. 密封小線源治療の臨床〃
15. 放射線治療における品質保障・精度管理〃

Evaluation Criteria⟩ 出席状況，質疑応答といった授業への取り組み姿勢などに基づく平常点での評価を行う． Re-evaluation⟩ 行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217543 Contact⟩

⇒ Harada (masafumi@medsci.tokushima-u.ac.jp) (Office Hour: Friday PM5:00∼ 7:00) ⇒ Otsuka (hotsuka@clin.med.tokushima-u.ac.jp) ⇒ Ikushima (+81-88-633-9051, ikushima@medsci.tokushima-u.ac.jp) ⇒ Ueno (+81-88-633-9020, ueno@medsci.tokushima-u.ac.jp) (Office Hour: 毎週金曜日 16:30-17:30) ⇒ Matsuzaki (kenji@clin.med.tokushima-u.ac.jp) ⇒ Takeuchi (mayumi@clin.med.tokushima-u.ac.jp)

Practice of Radiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Masafumi Harada · Professor / Institute of Health Biosciences

, Hitoshi Ikushima · Professor / Institute of Health Biosciences Hideki Otsuka · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Kenji Matsuzaki · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Mayumi Takeuchi · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Target⟩ 放射線診断学と放射線腫瘍学の研究を進めるため臨床の現場でデータ収集，研究計画立案，成果の解

析などを目的とする． Outline⟩ 3 次元画像診断，機能診断およびに放射線治療法に関する論文を読ませ，研究計画の立て方を学ばせ

る． (オムニバス方式) Manner⟩ オムニバス方式 Notice⟩ 臨床の現場での授業であるため，言動や服装には十分注意すること． Schedule⟩

大項目担当
1. 臨床現場での放射線診断の演習西谷弘
2. 臨床現場での放射線治療の演習生島仁史

Evaluation Criteria⟩ 出席状況，質疑応答といった授業への取り組み姿勢などに基づく平常点での評価を行う． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217544 Contact⟩

⇒ Harada (masafumi@medsci.tokushima-u.ac.jp) (Office Hour: Friday PM5:00∼ 7:00)

⇒ Ikushima (+81-88-633-9051, ikushima@medsci.tokushima-u.ac.jp)

⇒ Otsuka (hotsuka@clin.med.tokushima-u.ac.jp)

⇒ Matsuzaki (kenji@clin.med.tokushima-u.ac.jp)

⇒ Takeuchi (mayumi@clin.med.tokushima-u.ac.jp)

Respiratory Medicine and Rheumatology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Masaki Hanibuchi · Associate Professor / University Hospital Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217494 Contact⟩

⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp)

⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp)

⇒ Hanibuchi .

⇒ Goto .

Practice of Respiratory Medicine and Rheumatology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Masaki Hanibuchi · Associate Professor / University Hospital Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217495 Contact⟩

⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp)

⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp)

⇒ Hanibuchi .

⇒ Goto .

Clinical Oncology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Masaki Hanibuchi · Associate Professor / University Hospital Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217554

Contact⟩ ⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp) ⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp) ⇒ Hanibuchi . ⇒ Goto .

Practice of Clinical Oncology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Masaki Hanibuchi · Associate Professor / University Hospital Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217555

Contact⟩ ⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp) ⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp) ⇒ Hanibuchi . ⇒ Goto .

Thoracic, endocrine surgery and Oncology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akira Tangoku⟨Manager⟩ · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Shoji Sakiyama · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

, Kazuya Kondo · Professor / Institute of Health Biosciences

Target⟩ ヒト癌に関する基礎的，臨床的知識を踏まえて発癌過程，担癌生体の病態生理を理解させ，これに基づいて発癌ならびに治療に関する有用な先進的研究が行えるように指導する．また外科治療の本質的事項である外科手術侵襲と生体反応に関する理解を深め，外科治療学の発展に寄与し得るよう指導する．

Outline⟩ 以下の項目を主題として最新の知見をもとに講義を行う．担癌生体における様々な病態生理，外科的侵襲による生体反応．食道癌の発生，転移のメカニズム．抗癌剤耐性獲得のメカニズム．たばこやクロム酸塩によるヒト肺扁平上皮癌の段階的発癌に関する癌関連遺伝子の変化大腸発癌における各種胆汁酸の核レセプターを介した発癌関連遺伝子への調節機構．乳癌治療薬である選択的エストロゲン作用作動薬タモキシフェンの DNA 結合能と子宮内膜発癌の解析臓器移植，ことに肺移植におけるレシピントの生体反応，拒絶反応および移植臓器の再灌流障害．

Notice⟩ 講義の一部は e-learning 化の準備中

Schedule⟩

大項目内容
1. 抄読会毎週月・火曜日， 8:00∼ 8:30
2. 手術カンファレンス毎週火・木曜日 7:30∼ 8:30
3. 臨床カンファレンス毎週火曜日 17:00∼ 19:00
4. 研究指導随時

Evaluation Criteria⟩ 出席状況，随時口頭試問を行う．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217489

Contact⟩

⇒ Tangoku (+81-88-633-7141, tangoku@clin.med.tokushima-u.ac.jp) (Office Hour: 金曜日の16:00∼ 18:

00(e-mailにて時間調節を行います． ) ⇒ 他の教官についても,e-mailにて時間調節の上,面談してください.

Note⟩ 外科治療を本質的に理解し得る能力と外科系医師として求められる診断能力の習得を併せて目的とする．他分野との積極的な交流を図る．

Surgical Oncology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akira Tangoku · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217502 Contact⟩

⇒ Tangoku (+81-88-633-7141, tangoku@clin.med.tokushima-u.ac.jp)

Experimental Practice for Bioregulation and Medical Oncology 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Keisuke Izumi · Professor / Course of Bioregulation and Medical Oncology, Medical Science

, Hiromu Nishitani ·

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science Akira Tangoku · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217523 Contact⟩

⇒ Izumi (+81-88-633-7065, izumi@basic.med.tokushima-u.ac.jp) ⇒ Nishitani . ⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp) ⇒ Tangoku (+81-88-633-7141, tangoku@clin.med.tokushima-u.ac.jp)

Reconstructional Pathology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akiyoshi Nishimura · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217496 Contact⟩

⇒ Nishimura (Basic Bwing 3rd floor, +81-88-633-7084, ncc1701abcde@basic.med.tokushima-u.ac.jp)

Practice of Reconstructional Pathology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akiyoshi Nishimura · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217497 Contact⟩

⇒ Nishimura (Basic Bwing 3rd floor, +81-88-633-7084, ncc1701abcde@basic.med.tokushima-u.ac.jp)

Dermatology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshiaki Kubo · Associate Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217541 Contact⟩

⇒ Kubo (+81-88-633-7154, ykubo@clin.med.tokushima-u.ac.jp)

Practice for Dermatology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yoshiaki Kubo · Associate Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217542 Contact⟩

⇒ Kubo (+81-88-633-7154, ykubo@clin.med.tokushima-u.ac.jp)

Orthopedic Surgery 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year), 4th-year

Natsuo Yasui⟨Manager⟩ · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Shinsuke Katoh · Associate Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

, Koichi Sairyo ·

Mitsuhiko Takahashi · Assistant Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Target⟩ To study the fundamental mechanism of musculo-skeletal development, growth, repair, and

regeneration. To learn the current techniques and knowledges for advanced research. Outline⟩ We give systemic lectures elucidationg fundamental mechanism of musculo-skeletal development,

growth, repair and regeneration. We focus our attention to introduce the current tecniques and knowlegdes to study the fracture healing process, genes causing skeletal dysplasias, and molecular mechanism of the growth plate. We also explain the growth factors and cytokines that may affect repair and regeneration after spinal cord injury.

Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当

1. guidance for musculo-skeletal tissue research

Natsuo Yasui

2.

Matrix contents of bone and cartilage 〃

3.

Skeletal growth and development 〃

4.

Fracture repair mechanism 〃

5.

Bone formation and resorption 〃

6.

Aging and regeneration of bone 〃

7.

Development, growth and repair of the muscle Koichi Sairyo

8.

Structure and function of the tendon Mitsuhiko Takahashi

9.

Structure and function of the nerve 〃

10.

Reair and regeneration of the peripheral nerve 〃

11.

Development and growth of the spinal vertebra Koichi Sairyo

12.

Deveropment and grwoth oa the spinal cord Shinsuke katoh

13.

Intervertebral disc degeneration 〃

14.

Ossification of spinal ligaments Koichi Sairyo

15.

Genes causing skeletal dysplasias

Natsuo Yasui

Evaluation Criteria⟩ 出席状況と小テストまたはレポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217611 Contact⟩

⇒ Yasui (+81-88-633-7238, nyasui@clin.med.tokushima-u.ac.jp) (Office Hour: 毎週木曜午前9時―12時(

e-mail により時間調節を適宜おこないます)) ⇒ 他の教員についても， e-mail にて時間調節の上，面談して下さい．

Orthopedic Surgery Practice 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Natsuo Yasui · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217479 Contact⟩

⇒ Yasui (+81-88-633-7238, nyasui@clin.med.tokushima-u.ac.jp) (Office Hour: 8:00-18:00)

Plastic Surgery and Reconstructive Surgery 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hideki Nakanishi · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science Ichiro Hashimoto · Associate Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217483 Contact⟩

⇒ Nakanishi (+81-88-633-7296, nhideki@clin.med.tokushima-u.ac.jp)

⇒ Hashimoto (+81-88-633-7296, ichiro@clin.med.tokushima-u.ac.jp)

Practice for Plastic Surgery and Reconstructive Surgery 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hideki Nakanishi · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217484 Contact⟩

⇒ Nakanishi (+81-88-633-7296, nhideki@clin.med.tokushima-u.ac.jp)

Laboratory Experiment of Pathologic Medicine for Cutaneo-motoric System

12 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Akiyoshi Nishimura · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Natsuo Yasui · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science Hideki Nakanishi · Professor / Course of Pathologic Medicine for Senso-motoric System, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217482 Contact⟩

⇒ Nishimura (Basic Bwing 3rd floor, +81-88-633-7084, ncc1701abcde@basic.med.tokushima-u.ac.jp) ⇒ Yasui (+81-88-633-7238, nyasui@clin.med.tokushima-u.ac.jp) (Office Hour: 8:00-18:00) ⇒ Nakanishi (+81-88-633-7296, nhideki@clin.med.tokushima-u.ac.jp)

Graduate School of Medical Sciences (2011) ⟩ Doctor’s Degree ⟩ Course of Proteomics Medical Sciences

Course of Proteomics Medical Sciences Syllabus of Subjects

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ...38 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) .................38

Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka

・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) .......................................38

Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・

Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester)..............................................39 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao

・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ...................................................... 39 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・

Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga .....................................40

• 各専攻系間共通カリキュラム科目

A Training of Microbiology and Immunology … Miyake ・ Adachi ・ Yasutomo ・ Ono ・ Kuwahara ・ Uchiyama ・ Maekawa ・ Hirota / 1st-year(2nd semester), 2nd-year(2nd semester) .......................................... 41

• 指定科目

Introductory lab course for protein experiments … Takahama ・ Oyadomari ・ Okazaki ・ Katagiri / each-year(1st

semester) .......................................................................................... 41 Introductory practice of proteomic analysis … Sakaguchi ・ Taniguchi ・ Ebina ・ Uchiyama ・ Yuasa / 1st-year(2nd

semester) .......................................................................................... 41 Introductory practice for functional genomics … Matsumoto ・ Itakura ・ Kuwahara / 1st-year(1st semester).....42 Application of proteomics to medicine … Matsumoto ・ Adachi ・ Nakahori ・ Rokutan / 1st-year(2nd semester) ... 42

• 専門科目

Disease Proteomics … Sasaki ・ Sakane / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........ 43 Disease Proteomics Seminar … Sasaki ・ Sakane / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)43 Human Genetics … Imoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...................43 Human Genetics seminor … Imoto ・ Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)44 Bioregulatory Sciences and Medicine … Matsumoto ・ Azuma ・ Abe ・ Inoue ・ Akaike / 1st-year(whole year),

2nd-year(whole year), 3rd-year(whole year) .............................................................44 Bioregulatory Sciences and Medicine … Matsumoto ・ Azuma ・ Abe ・ Inoue ・ Akaike / 1st-year(whole year),

2nd-year(whole year), 3rd-year(whole year) .............................................................44 Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................44 Practice of Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 45 Stress Science … Rokutan ・ Tanahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)..........45 Stress Science Seminar … Rokutan ・ Tanahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) 45 Molecular Medicine Seminar … Sasaki・Adachi・Nakahori・Matsumoto・Rokutan・Kuwahara・Uchiyama / 1st-year(whole

year), 2nd-year(whole year), 3rd-year(whole year) ....................................................... 45 Proteomics in Enzymology … Kido ・ Yano / 1st-year(1st semester), 2nd-year(1st semester), 3rd-year(1st semester) 45 Practice in Proteomics Enzymology … Kido ・ Sakaguchi ・ Yano / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 46 Proteomics … Taniguchi ・ Fujiwara ・ Konishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .... 46 Proteomics seminar … Taniguchi ・ Fujiwara ・ Konishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................46 Molecular immunology … Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........46 Practice for molecular immunology … Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................47 Systems Biology for Molecular Signailng … Sakaguchi ・ Uchiyama ・ Mori / 1st-year(whole year), 2nd-year(whole

year), 3rd-year(whole year) ...........................................................................47

Practice on Systems Biology for Molecular Signailng … Sakaguchi ・ Uchiyama ・ Mori / 1st-year(whole year),

2nd-year(whole year), 3rd-year(whole year) .............................................................47 Proteome Genetics … Ebina ・ Yuasa ・ Uchiyama ・ Nagaya / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................47 Laboratory Practice of Proteome Genetics … Ebina ・ Uchiyama ・ Yuasa ・ Nagaya / 1st-year(whole year),

2nd-year(whole year), 3rd-year(whole year) .............................................................48 Research Progress in Enzyme Pathophysiology … Fukui ・ Sakai / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 48 Selected Topics on Enzyme Pathophysiology … Fukui ・ Sakai / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year) ................................................................................ 48 Course on Genomic Information … Itakura ・ Inoue / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................48 Practice on Genomic Information … Itakura ・ Inoue / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole

year) ..............................................................................................48 Genome Medicine … Katagiri / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)................49 Practice on Genome Medicine … Katagiri / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...49

Experimental practice of medicine for biological responses … Ebina ・ Kido ・ Taniguchi ・ Matsumoto ・ Sakaguchi

・ Fukui ・ Itakura ・ Hara ・ Tanigami ・ Takahashi ・ Yano ・ Fujiwara ・ Inoue ・ Sakai / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................................49 ゲノム機能解析学 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................49

ゲノム機能解析学演習 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................. 49

生体対応医学実験実習 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................. 49

Introduction to Biological Ethics 2 units (compulsory) each-year(1st semester)

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

, Seiichiro Kitamura · Professor / Institute of Health Biosciences

Junji Terao · Professor / Institute of Health Biosciences , Nobuyoshi Nakajo · Professor

, Kazuo Minakuchi · Professor / Institute of Health Biosciences

Target⟩ バイオサイエンスおよび医療に従事する者は，人権，生命倫理に十分な配慮を行い，個人情報保護，実験動物愛護にも同じく目を向けなければならない．本授業は生命倫理に関わる基本的知識を修得することを目的としている．

Outline⟩ 生命倫理学，臨床倫理学，社会倫理，個人情報保護，実験動物愛護などの問題に日頃接することの多

い講師が，経験に基づいた講義をオムニバス方式で行う． Notice⟩ 講義の一部は e-learning 化しているので， e-learning 学習も出席として取り扱う． Schedule⟩

大項目担当
1. 「ヒトゲノム · 遺伝子解析に関する倫理指針」について板倉光夫
2. 「実験動物管理と情報処理にまつわる倫理学」太田房雄
3. 「キャンパスハラスメントを通して「医療従事者の倫理」を考える」 北村清一郎
4. 「法規や制度に関する基本的考え方—法医学の立場から」久保真一
5. 「大学における人権問題」佐野壽昭
6. 「組織検体取り扱い上の倫理的問題」〃
7. 「食品の機能性 · 安全性の評価と社会倫理」寺尾純二
8. 「尊厳死や安楽死など生命倫理について」中條信義
9. 「遺伝医学の視点」中堀豊
10. 「先天異常， (or 神経疾患，家族性腫瘍) の遺伝カウンセリング」〃
11. 「動物実験倫理」松本耕三
12. 「臨床治験に関する倫理」水口和生
13. 「ヒト胚の倫理学的な地位」山野修司
14. 「臓器移植の倫理的問題 (仮題)」井藤久雄 · 予定
15. (未定)

Evaluation Criteria⟩ 出席状況 (厳格にする予定) を踏まえ，さらに小テストないし面接試験で評価することを

検討している． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218234 Contact⟩

⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp) ⇒ Kitamura (口腔顎顔面形態学分野教授室, +81-88-633-7319, kitamura@dent.tokushima-u.ac.jp) (Office

Hour: 前後期:金曜日 12 時 ∼ 13 時) ⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp) ⇒ Nakajo (+81-88-633-7366, nakajo@dentclin.tokushima-u.ac.jp) ⇒ Minakuchi (+81-88-633-7212, minakuti@clin.med.tokushima-u.ac.jp)

Clinical Psychology 2 units (selection) 1st-year(1st semester)

Kenji Sato · Professor / Institute of Socio-Arts and Sciences

, Mayumi YAMAMOTO · Professor / Institute of Socio-Arts and Sciences

Motohiro Sakai · Associate Professor / Institute of Socio-Arts and Sciences , Takaki Fukumori · Associate Professor / Institute of Socio-Arts and Sciences

Masato Mitome · Professor / Institute of Health Biosciences

Target⟩ 臨床心理学の基礎的理論・技法および今日的課題の概説

Outline⟩ 心と身体は密接につながっている．したがって「心の問題」の理解と制御を扱う臨床心理学の基礎の習得は，精神医学・心身医学のみならず，ヘルスバイオサイエンスを基盤とする医学，歯学，薬学，栄養学，保健学領域において重要である．そこで，本講義では，臨床心理学の定義，対象，方法 (代表的な心理検査，心理療法) について，初学者を考慮して，その基礎と今日的課題を概説する．心理療法に関しては，医学領域などでエビデンスを示す認知行動療法について詳述する．

Notice⟩ e-learning 対応．

Goal⟩ 臨床心理学の基礎的理論・技法および今日的課題を説明できる

Schedule⟩

大項目中項目
1. 臨床心理学の定義と対象 (佐藤)
2. 臨床心理学の測定方法 (1)  パーソナリティの査定法 (福森)
3. 心理療法・カウンセリングとは (福森)
4. 心理療法の基礎 (1)  精神分析 (福森)
5. 臨床心理学の測定方法 (2)  知能，発達の査定法 (山本)
6. 心理療法の基礎 (2)  分析心理学，クライエント中心療法 (山本)
7. 発達障害と特別支援教育の現状と課題 (山本)
8. 行動理論と行動療法 (境)
9. 認知療法うつ病 (境)
10. 認知行動療法不安障害 (境)
11. 臨床心理的地域援助の基礎と実際ひきこもりに焦点を当てて (境)
12. 外傷後ストレス障害の認知行動療法 (佐藤)
13. 摂食障害の認知行動療法 (佐藤)
14. 体重減量・糖尿病の認知行動療法 (佐藤)

Evaluation Criteria⟩ 対面講義， e-learning とも受講とレポート提出を以て出席扱い．対面講義の場合，授業内の 15 分間で書く「出席・質問票・ミニレポート」の提出を以て， e-learning の場合，視聴記録 (受講) と MLS オンラインのレポート提出を以て出席とみなす．どちらのレポート課題とも「講義内容を A4 一枚でまとめる」．出席が三分の二以上の受講生が評価対象者

Re-evaluation⟩ 再試験無し

Textbook⟩ 教科書は使用しない．参考書などは，適宜，授業時に紹介する

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218235

Contact⟩

⇒ Mitome (mitome@dent.tokushima-u.ac.jp) (Office Hour: 三留:水曜日の12:00∼ 13:00(e-mail により調整可能))

Introduction to social medicine, epidemiology and biostatistics 2 units (selection)

each-year(2nd semester)

Kokichi Arisawa · Professor / Course of Social and Environmental Medicine, Medical Science

, Hiro-O Ito · Professor / Institute of Health Biosciences

Hirokazu Uemura · Associate Professor / Course of Social and Environmental Medicine, Medical Science

, Aiko Yamauchi · Professor / Institute of Health Biosciences

Tetsuya Tanioka · Professor / Institute of Health Biosciences

, Akira Tokumura · Professor / Institute of Health Biosciences

Daisuke Hinode · Professor / Institute of Health Biosciences , Hiroki Moriguchi · Professor / Course of Integrated Brain Sciences, Medical Science

Hiroaki Yanagawa · Associate Professor / University Hospital Target⟩ 1, Clinical trial as a tool to establish clinical evidence 2. The role of the IT in the medical treatment

and the hospital management are outlined. 3. Outline analytical skills for clarifying the phenomenon about Mental Health and Welfare. 4. Outline epidemiology and analytical methods for oral health promotion and/or oral disease prevention. 6. Outline damages from medicines and its etiology and prevention. 7. Outline designing of surveys and statistical analysis of data for writing a thesis.

Outline⟩ 1) Principles and practice of clinical trial from scientific and ethical viewpoint will be reviewed. 2) The usage of IT and the method of the management in the university hospital are outlined. 3) We will explain the statistical method for clarifying the feature of phenomenon based on previous work in Mental Health and Welfare. 4) We will explain the etiology of oral disease and its prevention based on the evidence of oral sciences. 6) Instruct background and etiology of several examples of damages from medicines and discuss possible prevention of the damages. 7) Basic aspects of statistical analysis, such as type of data and appropriate treatment of data, and meaning of hypothesis testing, will be summarized. In addition, practice of data analysis will be done using SPSS.

Notice⟩ It is possible to take a lesson by e-learning. Goal⟩ The goal is to deepen one’s understanding of contents presented during lectures, regarding social

medicine, pharmaceutical science and dentistry.

Schedule⟩

大項目担当

1.

Basic handling and statistical analysis of data for article makingHirokazu Uemura

2.  〃〃

3.

Practice of epidemiology and statistical data analysis for writing a thesis

Kokichi Arisawa

4.  〃〃

5.

Principles and practice of clinical trial Hiroaki Yanagawa

6.  〃〃

7.

The role of IT in a hospital Hiroki Moriguchi

8.  〃〃

9.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (1)

Shu-ichi Ueno

10.

Statistical procedure for analyzing issues facing the Mental Health and Welfare (2)

Tetsuya Tanioka

11.

Epidemiology of dental caries and periodontal disease, and the relationship between oral disease and systemic disease

Hiro-o Ito

12.

〃

Daisuke Hinode

13. 医薬品の情報管理，医薬品の安全対策木原勝
14.  〃〃

15.

What is damages from medicines? Why did the damages occurred? How to prevent the damages?

Akira Tokumura

16.

What is the damages from medicines? Why did the damages occurred? How to prevent the damages?

〃

Evaluation Criteria⟩ Reports should be submitted on questions presented during lectures. Grades are

determined by summing points from all instructors. Re-evaluation⟩ In principle, there is no reexaminations. Textbook⟩ Teaching materials designated by each instructor wil be used.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218231 Contact⟩

⇒ Arisawa (+81-88-633-7071, arisawa@basic.med.tokushima-u.ac.jp) (Office Hour: Please contact eac

h instructor by an e-mail.)

Note⟩ Different classrooms may be used according instructors.

Introduction to How to Write Up Scientific Manuscript in English 2 units (selection)

each-year(2nd semester)

Shuhei Tomita · Professor / Course of Biological Medicine, Medical Science

, Ryuji Kaji · Professor / Course of Sensory Neuroscience, Medical Science

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Yutaka Nakaya · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Hirofumi Nagashino · Professor / Institute of Health Biosciences

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Bukasa Kalubi · Assistant Professor / Medical Science

Target⟩ 21 世紀に医学，歯学，薬学，栄養学，保健学の各分野で活躍する人材は英語が堪能であることが要求される．本授業ではこれらの領域で用いられる独特の英語表現法に関わる基本的知識を修得することを目的とする．

Outline⟩ 研究成果を国際的な学術雑誌に発表したり，海外の学会やシンポジウムで発表や講演をしたりすることは，研究者養成を目的とする大学院教育において必須の履修目標である．そこで，本授業では医学英語論文，用紙のの作成方法について系統的な講義をビデオ並びにマルチメディア教材等を積極的に活用しながら行う．更に，医科学用英語の聴き取り及び英語による討論の訓練を行い，発表技術の向上を目指す． Notice⟩ 1) 授業は後期のみに開講される．時間帯は原則的に火曜日の午後であるが，講師の事情等により変更されることがある． 2) 講義の一部は e-learning 化されているので， e-learning 学習も出席として取り扱う． 3)20 分以上の遅刻は出席と見なさない．

Schedule⟩

大項目担当

1.

生物医学雑誌投稿に要求される条件 (I)

Fukui Kiyoshi

2. 生物医学雑誌投稿に要求される条件について (II)  〃
3. 研究，論文，学会発表の進め方 (I)  市原明

4.

研究，論文，学会発表の進め方 (II)

Ichihara Akira

5. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (I) 長篠博文
6. 論文投稿申込書，経歴書，研究計画書，履歴書などの書き方 (II)  〃
7. 学会口演要旨作成上の留意点について実例をまじえて冨田修平
8. 論文作成に役立つコンピュータの利用法ー実例を示した結果，考察の書き方ー 中屋豊

9.

英語による PowerPoint presentation についての要点

梶龍兒

10. 英語論文の特性と日本語論文との比較林良夫
11. 口頭による英語発表と短報の書き方に関する基本を講義する I  岡崎拓
12.  〃〃
13. 研究の区切りとしての論文作成 I  福井裕行
14. II

Evaluation Criteria⟩ 講義への出席，随時の試験，受講態度を総合して行う． 2/3 以上の出席がなければ不合

格とする．再試験はない． Textbook⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218236

Psychosomatic health and environmental stress 2 units

Eiji Takeda · Professor / Institute of Health Biosciences

, Tetsuro Ohmori · Professor / Course of Integrated Brain Sciences, Medical Science

Yousuke Takahama · Professor / Course of Preventive Medicine, Medical Science

, Hiroyoshi Sei · Professor / Course of Integrated Brain Sciences, Medical Science

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

Junji Terao · Professor / Institute of Health Biosciences

, Ken-ichi Miyamoto · Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Rintaro Yamanishi · Associate Professor / Institute of Health Biosciences

Hiroshi Sakaue · Associate Professor / Institute of Health Biosciences , Toshio Hosaka · Assistant Professor / Institute of Health Biosciences Target⟩ The aim of this program is to study on psychosomatic stress in daily life. Contents of this program

include gene expression profile under stress, stress response in health and diseases. Outline⟩ By the systems biology, clinical neurology, nutritional science techniques, the effects of stress on

development and growth in child will be assessed and discussed. �������⟩ psychosomatic health, environment, stress Goal⟩ To understand stress assessment technology

Schedule⟩

大項目担当

1.

Anxiety and depression (Tetsurou Ohmori) Noriaki Okazaki

2.

The development of the immune system (Yousuke Takahama)

〃

3.

Mechanism of sleep regulation (Hiroyoshi Sei)

Ryuji Kaji

4.

Endoplasmic reticulum stress in health and diseases (Seiichi Oyadomari)

Kazuhito Rokutan

5.

Stress genomics (Kazuhito Rokutan)

〃

6.

New concept and innovation of functional foods in Japan (Junji Terao)

Hiroyoshi Sei

7.

Why do we need ”stress control” through foods and food style? (Junji Terao)

〃

8.

Aging and mineral nutrition (ken-ichi Miyamoto)

Megumi Goto

9.

Caloric restriction and bone metabolism (Ken-ichi Miyamoto)

〃

10.

Breakfast and chewing for stress (Eiji Takeda) Hiroshi Chuman

11.

Food and stress related hormone (Eiji Takeda) Suehiro Sakaguchi

12.

Physical impact caused by stress (Yutaka Nakaya)

Eiji Takeda

13.

Gut immunity and stress (Yutaka Nakaya) Tetsurou Ohmori

14.

Oxidative stress and immunity (rintaro Yamanishi)

Gen Tamiya

15.

The role of adipokines on metabolic stress in obesity and obesity-related diseases (hiroshi Sakaue)

〃

16.

Diabetes mellitus and stress (Toshio Hosaka)

Eiji Takeda

Evaluation Criteria⟩ Reports on more than two third of the studies will be required for evaluation. Re-evaluation⟩ 再試験無し Textbook⟩ none Reference⟩ 授業時に適宜，紹介する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218232 Contact⟩

⇒ Takeda (+81-88-633-7093, takeda@nutr.med.tokushima-u.ac.jp) ⇒ Ohmori (臨床研究棟 8F 教授室, +81-88-633-7130, tohmori@clin.med.tokushima-u.ac.jp) (Office Hour:

8:30-17:30) ⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp) ⇒ Sei (+81-88-633-7056, sei@basic.med.tokushima-u.ac.jp) ⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Oyadomari (408, +81-88-633-9450, oyadomar@genome.tokushima-u.ac.jp) ⇒ Terao (+81-88-633-7087, terao@nutr.med.tokushima-u.ac.jp) ⇒ Miyamoto (Molecular Nutrition, +81-88-633-7081, miyamoto@nutr.med.tokushima-u.ac.jp) ⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Yamanishi (+81-88-633-9366, rintaro@nutr.med.tokushima-u.ac.jp) ⇒ Sakaue (303, +81-88-633-7091, hsakaue@nutr.med.tokushima-u.ac.jp) ⇒ Hosaka .

Basic methods in life sciences 2 units

Hiroyuki Fukui · Professor / Institute of Health Biosciences

, Hiroyuki Mizuguchi · Associate Professor / Institute of Health Biosciences

Yutaka Nakaya · Professor / Institute of Health Biosciences

, Tatsuji Haneji · Professor / Institute of Health Biosciences

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Kazuko Seta · Part-time Lecturer

Kouji Itou · Professor / Institute of Health Biosciences

, Takahiro Matsumoto · Associate Professor / Institute of Health Biosciences

Keiko Kataoka · Associate Professor / Course of Microbiology and Immunology, Medical Science

, Tetsuo Yamazaki · Professor / Institute of Health Biosciences

Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science

, Takeshi Nikawa · Professor / Institute of Health Biosciences

Hironori Yamamoto · Assistant Professor / Institute of Health Biosciences

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics , Toshihiko Nagata · Professor / Institute of Health Biosciences Tetsuya Yoshinaga · Professor / Institute of Health Biosciences

Target⟩ 大学院に進学した直後の院生に生命科学に関する基礎的な実験方法を理解させる．

Outline⟩ 実験動物の取り扱い方，細胞と器官の培養方法とその応用，蛋白質の取り扱い方とその解析方法，遺伝子解析の方法とその応用，抗体を用いた研究方法とその蛋白質，免疫組織細胞化学に対する応用，数理モデルによる生体機能発現機構の解明等生命科学の基礎的な研究技法を講義する．

�������⟩ 生命科学, 実験手法

Notice⟩ 受講生は必ず出席すること．

Goal⟩ 現在行われている生命科学研究方法の基礎を大学院進学直後に理解し，研究生活にスムーズに入れるよ

うにする． Schedule⟩

大項目

1. {大学院に入学直後に現在用いられている研究手法を集中的に講義する．第 1 回目は授業ガイダンスと生命科学の研究手法総 \atop 論 (羽地，研究部長)}
2. 実験動物を利用した研究 (松本)
3. 細胞と器官の培養 (羽地)
4. 細胞培養の応用 (宮本)
5. 蛋白研究 1(福井清)
6. 蛋白研究 2(藤原)
7. 抗体を用いた生化学的研究方法 (二川，山本)
8. 抗体を用いた研究の臨床応用 (中屋)
9. 免疫学研究の基礎 (安友)
10. 免疫学研究の最近の進歩 (安友)
11. 遺伝子解析 1(福井裕)
12. 遺伝子解析 2(伊藤)
13. 遺伝子解析 3(高濱)
14. 遺伝子解析 4(高濱)
15. 数理モデルによる生命機能の解析 (吉永)

Evaluation Criteria⟩ Web によるレポート，出席状況，受講態度等により総合的に判定する． Re-evaluation⟩ しない Textbook⟩ 指定しないが，講義の都度プリント等資料を配布する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218233 Contact⟩ ⇒ Fukui (hfukui@ph.tokushima-u.ac.jp) ⇒ Mizuguchi (+81-88-633-7264, guchi003@ph.tokushima-u.ac.jp) ⇒ Nakaya (栄養学科棟 302, +81-88-633-7090, yutaka-nakaya@nutr.med.tokushima-u.ac.jp) (Office

Hour: Tuesday from 11:00 to 14:00) ⇒ Haneji (+81-88-633-7321, tat-hane@dent.tokushima-u.ac.jp) ⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 8:30 to 17:30) ⇒ Seta . ⇒ Itou (Institute for Medicinal Resources, +81-88-633-7290, kitoh@ph.tokushima-u.ac.jp) (Office Hour:

Monday, 17:0019:00) ⇒ Matsumoto . ⇒ Kataoka (kataoka@basic.med.tokushima-u.ac.jp) ⇒ Yamazaki (医学臨床 B 棟 5 階医薬品病態生化学研究室, +81-88-633-7886, yamazakt@ph.tokushima-u

.ac.jp) (Office Hour: 特に定めない) ⇒ Okazaki (414, +81-88-633-9158, tokazaki@genome.tokushima-u.ac.jp) ⇒ Nikawa (nikawa@nutr.med.tokushima-u.ac.jp) ⇒ Yamamoto (+81-88-633-7235, yamamoto@nutr.med.tokushima-u.ac.jp) ⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp) ⇒ Nagata (+81-88-633-7343, nagata@dent.tokushima-u.ac.jp) ⇒ Yoshinaga (保健学 B 棟教員研究室 (吉永), +81-88-633-9050, yosinaga@medsci.tokushima-u.ac.jp)

(Office Hour: 18:00-19:00 every Friday)

A Training of Microbiology and Immunology 2 units (selection)

1st-year(2nd semester), 2nd-year(2nd semester)

Yoichiro Miyake · Professor / Institute of Health Biosciences

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Koji Yasutomo · Professor / Course of Microbiology and Immunology, Medical Science

, Tsuneko Ono · Professor / Institute of Health Biosciences

Tomomi Kuwahara · Associate Professor / Course of Microbiology and Immunology, Medical Science Tsuneo Uchiyama · Associate Professor / Course of Microbiology and Immunology, Medical Science Yoichi Maekawa · Associate Professor / Course of Microbiology and Immunology, Medical Science Katsuhiko Hirota · Associate Professor / Institute of Health Biosciences

Target⟩ 微生物の取り扱いの基本，及び免疫学的実験法の基本を学ぶ Outline⟩ · · Notice⟩ 1．病原微生物実習室で行うので，それ専用の自分の白衣とスリッパ (上履き) を持参のこと (桑原知

巳)． 2．白衣と筆記用具を持参のこと (三宅洋一郎 · 弘田克彦)． Schedule⟩

大項目中項目
1. (1˜ 4) 細菌学実験の基本操作と常在菌や病原菌の分離 (桑原知巳) {培地の作製と培養，滅菌 · 消毒法，無菌操作などの基本的手技を学び，常在菌 \atop や病原菌の分離 · 同定を行い，身近な細菌の性状を理解する．}

2. (5, 6) 分子ウイルス学的解析法 1(足立昭夫) バイオハザードや分 子遺伝学等，現代ウイルス学の基礎知識を習得するとともに基 本的技能について学ぶ．

3. (7, 8) 分子ウイルス学的解析法 2(内山恒夫) ウイルスの同定法や ウイルス感染の解析法に関する知識を習得し，その基本的手技 について学習する．

4. {(9,10) リンパ球活性化機構 (安友康二) リンパ球の分離方法の基 \atop 本的手技と，その活性化測定方法を学ぶ．}

5. {(11) マラリア原虫の診断 (前川洋一) マラリア原虫の分類とその \atop 鑑別方法を学ぶ．}

6. (12˜ 14) 口腔常在菌とバイオフィルム (三宅洋一郎 · 弘田克彦) 口 ; 腔常在菌の分離培養 · 同定などに関する基本的手技と，細菌バ ; イオフィルムの作成法及びそれを用いた各種の実験法を学ぶ．

7. {(15) 細菌感染症の分子遺伝学的検出法 (小野恒子) 菌種特異的遺 \atop 伝子マーカーを用いた細菌の検出と定量法について学ぶ．}

Evaluation Criteria⟩ 出席率と実習レポートによる．

Re-evaluation⟩ (再試験) 実習なので再試験はしない．出席率が 60%に満たない学生は，受講しなかった実習

を，次学年後期に追加して受けること． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217022 Contact⟩

⇒ Miyake (+81-88-633-7329, miyake@dent.tokushima-u.ac.jp) (Office Hour: 8:00∼ 8:45(e-mailでも対応します))

Introductory lab course for protein experiments 2 units (selection) each-year(1st semester)

Yousuke Takahama⟨Manager⟩ · Professor / Course of Preventive Medicine, Medical Science

Seiichi Oyadomari · Professor / Course of Integrated Brain Sciences, Medical Science

, Taku Okazaki · Professor / Course of Preventive Medicine, Medical Science Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 医科学研究の基礎技術としての蛋白・酵素機能解析学的研究手法の基礎を修得し，それらの適応と限

界を理解する． Outline⟩ 遺伝子組換え技術の基礎を解説するとともに実習により指導する．また，ゲノム解析や遺伝子発現解析の原理と応用について解説するとともに， PCR に基づくゲノム・遺伝子解析技術を指導する．更に，タンパク質の分離精製法，同定法，定量法について基礎理論を解説するとともに実習により基礎技術を指導する．いずれの実習についても，実際にそれぞれの技術を活用している研究室にて実施する．

Schedule⟩

大項目内容担当
1. {DNA 組換えとクローニング (講義 45 分 \atop ×2，実習 45 分 ×4)} ; {講義項目: プラスミドの特性・クローニング・シーケンシング / 実習項目: 制 \atop 限酵素処理・電気泳動・断片精製・ライゲーション・形質転換} ; 冨田

2.

ゲノム解析 (講義 45 分 ×2，実習 45 分 ×4)

講義項目: ゲノム診断・法医学的検査・ DNA 抽出・サザン・ PCR ・データベース利用 / 実習項目: PCR

高浜

3.

{遺伝子発現解析 (講義 45 分 ×2，実習 45 \atop 分 ×4)} 講義項目: ノザン・ RT-PCR ・ライブラリスクリーニング・アレイ解析 / 実習項目: RT-PCR

塩見 (美)

4. {タンパク質の精製 (講義 45 分 ×2，実習 45 \atop 分 ×4)} ; {講義項目: カラムの種類と特性・タグの種類と特性・ゲル染色法 / 実習項目: \atop アフィニティクロマトグラフィ・ SDS-PAGE} ; 桑島

5.

{タンパク質の同定 (講義 45 分 ×2，実習 45 \atop 分 ×4)} 講義項目: 電気泳動法の種類と特性・ウェスタン・マススペック・ ELISA / 実習項目: ELISA

塩見 (春)

Evaluation Criteria⟩ 技術習得状況および出席状況により評価する．単位認定には，講義と実習から構成され

る上記 5 セットのうち，最低 3 セットへの出席が必要である． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217609 Contact⟩

⇒ Takahama (ゲノム機能研究センター 208, +81-88-633-9452, takahama@genome.tokushima-u.ac.jp)

(Office Hour: 水曜日16:00∼ 18:00(電子メイルにより適宜面談時刻を調整すること)． ) ⇒ 他の教員についても電子メイルにて面談時刻を調節すること.

Introductory practice of proteomic analysis 2 units (selection) 1st-year(2nd semester)

Suehiro Sakaguchi · Professor / Course of Medicine for Biological Responses, Proteomics Hisaaki Taniguchi · Professor / Course of Medicine for Biological Responses, Proteomics

Yousuke Ebina⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Keiji Uchiyama · Associate Professor / Course of Medicine for Biological Responses, Proteomics Tomoyuki Yuasa · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 疾患原因の解明と新しい治療法の確立を目指したプロテオミクス研究の実践的研究能力を高める． Outline⟩ 二次元電気泳動・高速液体クロマトグラフィ・高感度質量分析計による自動的な構造解析法・分子間

相互作用解析法などプロテオーム解析に必要な基礎実験の実習を行う． Notice⟩ e-learning 学習は扱っていない

Schedule⟩

大項目担当
1. 授業ガイダンス+プロテオーム解析入門実習序論蛯名
2. 蛋白質について (1)  内山
3. 蛋白質について (2)  〃

4.

蛋白質の電気泳動原理と Western blot(1)

湯浅

5.

蛋白質の電気泳動原理と Western blot(2)

〃

6. 質量分析法の原理谷口
7. ゲル内消化〃
8. 質量分析法の実際〃
9. データベース検索〃
10. 蛋白結合と情報伝達について (1)  坂口
11. 蛋白結合と情報伝達について (2)  〃
12. 結合分子の同定法について (1)  〃
13. 結合分子の同定法について (2)  〃
14. 免疫共沈法の原理と手技 (1)  〃
15. 免疫共沈法の原理と手技 (2)  〃

Evaluation Criteria⟩ 出席状況・実習態度・レポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217575 Contact⟩

⇒ 他の教員についても， e-mailにて時間調節の上，面談してください．

Introductory practice for functional genomics 2 units (selection) 1st-year(1st semester)

Mitsuru Matsumoto⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics

Tomomi Kuwahara · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ ゲノムのもつ遺伝情報と，実際に生体内で発揮されている遺伝子機能との関連を研究するための方法

論を学ぶ． Outline⟩ 生物ゲノムは多数の遺伝子を含む．これらの遺伝子がコードする遺伝情報の意義づけを蛋白質の機

能とその発現調節機構から明らかにするための実習を行う．例えば，具体的な遺伝子をとりあげたバイオインフォマティックス等について実習を行う．

Notice⟩ 第一回目に開講するオリエンテーションには，必ず出席すること．また，予定の変更もありうるので，受講に当たっては事前に担当教官に電話等で確認すること．講義の一部については e-learning 化も検討しているが，利用可能になった場合には e-learning 学習も出席として取り扱う．

Schedule⟩

大項目担当
1. 各種解析方法の指導を担当する教官と事前に連絡を取り，該当する研究が行われる時期を設定し て実習を行う．遺伝子から細胞へ，細胞から個体へという一連のゲノム機能研究方法について， 総括的に習得する．

2. 授業ガイダンス
3∼6. ゲノム機能解析序論松本・板倉・桑原
7∼11. ゲノム機能解析理論〃

12∼16.ゲノム機能解析応用〃 Evaluation Criteria⟩ 受講状況と授業内容の理解の程度に応じて評価する．再試験は行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217559 Contact⟩

⇒ Matsumoto (Room 701, +81-88-633-7432, mitsuru@ier.tokushima-u.ac.jp) (Office Hour: 火曜日の16:

00∼ 18:00 (e-mail により時間調節を適宜，行います)) ⇒ 他の教員についても, e-mail にて時間調節の上,面談して下さい.

Application of proteomics to medicine 2 units (selection) 1st-year(2nd semester)

Toshio Matsumoto⟨Manager⟩ · Professor / Molecular Medicine, Proteomics

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Yutaka Nakahori ·, Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

Target⟩ <GIO> プロテオミクス研究の医学への応用の理論と手法を理解し，基本的手技を習得することにより，疾患の発症機構の解明および治療法の開発に貢献し得る人材となる． <SBO>

1. HIV の発症機構，特に HIV と宿主との蛋白質相互作用を理解する．

2. ウイルス感染によるアポトーシスの制御機構とその意義を理解する．

3. フローサイトメトリーの原理を理解し，疾患解析への応用を実践できる．

4. 遺伝子組替えの原理と手法を理解し，その基本手技を実践できる．

5. 染色体マーカー，遺伝子多型の概念を理解し， PCR などによる解析を実践できる．

6. 疫学研究の手法を理解し，基本的な調査法や統計解析を実践できる．

7. 遺伝子の転写・発現および蛋白質の翻訳と修飾の基本を理解する．

8. DNA チップによる解析の原理を理解し，疾患関連遺伝子同定のためのバイオインフォマティックスを実践できる．

Outline⟩ プロテオミクスの研究法を用いた，病因の解析と治療法の開発を行う教育プログラムである．ゲノム (遺伝子) とプロテオーム (蛋白質) の情報を駆使して，病因の解析と治療法に結びつく標的遺伝子と蛋白質を選び出して研究する方法について実習する．

Notice⟩ 生体情報内科学，ウイルス病原学，分子予防医学，ストレス制御医学の各分野の研究室において，講義および演習を行う．プロテオミクス研究の手法，考え方とその疾患へのアプローチを理解してほしい．授業は英語で行われることがある．

Schedule⟩

大項目中項目担当

1.

HIV の分子遺伝学的解析エイズの病因ウイルス HIV を対象に，ウイルス病の発症機構を解 明しその制御法を確立するための病原と方法論を概説する． HIV 蛋白質と宿主細胞蛋白質の機能的相互作用の解析を通じてウイルス 学，免疫学蛋白質化学，分子遺伝学等の主要な技法を修得させる．足立

2. {細胞のアポトーシス制御機構に作用するウイルス蛋 \atop 白の機能解析}ウイルス感染に伴うアポトーシスについて，生物学的意義ならび にその制御に関与する蛋白質相互作用を明らかにすることを目的 とする．ウイルス感染細胞もしくはウイルス遺伝子を導入した形 質転換細胞が示すアポトーシス誘導処理やサイトカインに対する 挙動を細胞生物学的手法や生化学的手法を用いて解析する．〃

3. フローサイトメトリー法による細胞の分析・分画法{フローサイトメトリー法の原理，方法とその各種実験への応用を \atop 概説し，これを用いた細胞の分析・分画採取法の実習を行う．}松本

4. 遺伝子組み替え蛋白の発現による細胞機能解析組み替え遺伝子をベクターに導入し，発現させた変異蛋白による 細胞機能変化を捉える手法の原理と方法を概説し，その応用を含 めた実習を行う．〃

5. 実際にヒトの材料を用いて研究している現場で，プ ロテオミクスの基本的手技がどのように行われてい るかを経験する．中堀

6. {遺伝疫学的な考え方にもとづいて，集団データがど \atop のように解析されているかを経験する．}〃

7. ストレス応答を例に，シグナル伝達，遺伝子の活性化 と転写， RNA の加工，タンパク質翻訳の機構を解明 するために必要な考え方，アプローチの仕方，及び実 際の研究手法を習得させる．六反

8. DNA チップ解析による遺伝子発現のバイオインフォ マテイクスを実際に行い，各種疾患のバイオマーカー を同定する手法を習得させる．〃

Evaluation Criteria⟩ 講義，実習における質疑応答や取り組む姿勢，理解度，実技内容により判断する．再試

験は必要に応じて試行する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217564

Disease Proteomics 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Takuya Sasaki⟨Manager⟩ · Professor / Molecular Medicine, Proteomics

, Ayuko Sakane · Assistant Professor / Institute of Health Biosciences

Target⟩ 細胞機能の制御機構の研究に必須の基礎知識を習得させるとともに，最近のトピックスを紹介する．

特に，細胞内・細胞間シグナル伝達機構，細胞骨格，細胞内小胞輸送をキーワードとした研究を取り上げる． Outline⟩ ヒトの疾病を細胞機能の制御機構の破綻として捉え，その制御に関わる機能蛋白分子群を同定し，

個々の分子の機能および分子間の相互作用を解析していくことが疾病の病態解明につながりつつあることを，最近のプロテオミクス研究の成果を中心に総括的に解説する．

Notice⟩ 遅刻しないこと． Schedule⟩

大項目

1. 授業ガイダンス
2. プロテオミクス病態学の基礎
3.  〃
4.  〃
5.  〃
6.  〃
7. 最近のトピックス (細胞内シグナル伝達機構)
8.  〃
9. 最近のトピックス (細胞間シグナル伝達機構)
10.  〃
11. 最近のトピックス (細胞骨格)
12.  〃
13. 最近のトピックス (細胞内小胞輸送)
14.  〃
15. 授業総括

Evaluation Criteria⟩ 出席のみ (講義の出席を原則としているので，出席が満たない場合は再履修)． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217567 Contact⟩

⇒ いつでも.ただしあらかじめメールに用件を書くとともに,時間のアポイントをとること. Note⟩ 本授業科目を担当する分子病態学分野は，あらゆる細胞機能の制御において基本となるしくみに注目し

て研究を進めていることから，多くの研究領域との関わりがあり，どの研究領域の受講者であっても，本授業科目で習得した知識を受講者自身の研究テーマにつなげることが可能であるという点が特色である．

Disease Proteomics Seminar 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Takuya Sasaki⟨Manager⟩ · Professor / Molecular Medicine, Proteomics , Ayuko Sakane · Assistant Professor / Institute of Health Biosciences Target⟩ 細胞機能の制御機構の研究に必要な基礎知識と基本技術を習得させることを目的とする． Outline⟩ プロテオミクス研究を基盤とした，細胞機能の制御に関わる機能蛋白分子群の同定とそれらの作用

機構および相互作用の解析について演習を行う．さらに，研究成果を疾患の病態解明や診断，治療法の開発につなげる方法論についても指導する．

Notice⟩ 遅刻しないこと． Schedule⟩

大項目

1. 授業ガイダンス
2. プロテオミクス病態学演習の基礎
3.  〃
4.  〃
5.  〃
6.  〃
7. プロテオミクス病態学演習の方法論 (細胞内シグナル伝達機構)
8.  〃
9. プロテオミクス病態学演習の方法論 (細胞間シグナル伝達機構)
10.  〃
11. プロテオミクス病態学演習の方法論 (細胞骨格)
12.  〃
13. プロテオミクス病態学演習の方法論 (細胞内小胞輸送)
14.  〃
15. 授業総括

Evaluation Criteria⟩ 出席のみ (講義の出席を原則としているので，出席が満たない場合は再履修)． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217568 Contact⟩

⇒ いつでも.ただしあらかじめメールに用件を書くとともに,時間のアポイントをとること. Note⟩ 本授業科目を担当する分子病態学分野は，あらゆる細胞機能の制御において基本となるしくみに注目し

て研究を進めていることから，多くの研究領域との関わりがあり，どの研究領域の受講者であっても，本授業科目で習得した知識を受講者自身の研究テーマにつなげることが可能であるという点が特色である．

Human Genetics 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Issei Imoto · Professor / Molecular Medicine, Proteomics

Target⟩ 人類遺伝学の基礎的な知識を習得し，健康の保持，疾病予防について，人類遺伝学的知識の疫学的応

用について学ぶ． Outline⟩ 人類遺伝学，遺伝疫学の本を選定し輪読会形式で読み進み解説を加える．各学生が，予習と説明を

担当する． (平成 18 年度は日本語訳された Essential 細胞生物学を読み，人類遺伝的な視点で解説する) Notice⟩ 教科書に指定された本はコピーせずに買うこと． e-learning 対象ではない． Schedule⟩

大項目担当
1. 染色体，遺伝子，ゲノム，連鎖解析， SNPs 等の基本的事項を理解していることを目標とする．
2. 役割分担佐藤陽一
3∼16. 数回で 1 章を読むことを目標にする．各自が自習してくることが望ましい． 〃

Evaluation Criteria⟩ 出席，担当，報告に対する評価をもって成績を判定する． 1/3 以上の欠席は単位を与え

ない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217514 Contact⟩

⇒ Imoto (+81-88-633-7455, issehgen@basic.med.tokushima-u.ac.jp) Note⟩ 選択した学生は，分担または当番があるので，自分が当たったところは完全に理解できるまで読んでく

ること．

Human Genetics seminor 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Issei Imoto · Professor / Molecular Medicine, Proteomics

, Toshio Matsumoto · Professor / Molecular Medicine, Proteomics

Target⟩ 原則的に遺伝疫学を履修していることを前提としている．遺伝疫学で学んだことを実際の人間集団の

データを用いて演習を行う． Outline⟩ 分子予防分野で行っている研究のフィールド活動から得られたデータの入力，データ解析，解析の解

釈などを実際に経験する．また，自分自身でデータを持っている学生については，そのデータを利用する． Notice⟩ アンケートのパソコン入力をいい加減にすると，結果はまったく信頼できないものになります．この

ことを心して下さい． e-learning 対象ではない． Schedule⟩

大項目担当
1. 1∼ 5 を各 3 回ずつ程度に分けて行う
2. 有効なアンケートの作り方を学ぶ勢井雅子
3. データの入力を行う〃
4. 遺伝子解析の実際を見学する〃
5. 統計的処理を行う〃
6. 結果について小発表会を行う〃

Evaluation Criteria⟩ 演習態度，理解度の評価 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217515 Contact⟩

⇒ Imoto (+81-88-633-7455, issehgen@basic.med.tokushima-u.ac.jp) ⇒ Matsumoto (+81-88-633-7119, toshimat@clin.med.tokushima-u.ac.jp)

Note⟩ 他の人に迷惑がかかるので，何回も休む人は取らないで下さい．

Bioregulatory Sciences and Medicine 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshio Matsumoto⟨Manager⟩ · Professor / Molecular Medicine, Proteomics

, Hiroyuki Azuma ·

Masahiro Abe · Associate Professor / Molecular Medicine, Proteomics , Daisuke Inoue ·

, Masashi Akaike · Professor / Institute of Health Biosciences

Target⟩ <GIO> 加齢に伴い発症あるいは増悪する多因子病の病態・発症機序の解明および予防・治療法の開発に関わる研究能力を養う． <SBO> 1．骨粗鬆症の発症機構と研究手法を理解する． 2．癌転移や多発性骨髄腫の骨病変の発症・進展機構とその研究手法を理解する． 3．動脈硬化症の発症・進展機構とその研究手法を理解する． 4．糖尿病・高脂血症の発症機序とその研究手法を理解する．

Outline⟩ 加齢に伴い発症，進行あるいは頻度が増加する糖尿病，骨粗鬆症，動脈硬化症などの内分泌・代謝系，循環器系の多因子病や多発性骨髄腫などの血液・腫瘍疾患について，その背景となる遺伝子多型や変異，染色体転座，蛋白修飾などによりもたらされるプロテオーム系の変化による生体情報の異常に視点をおき，病態および発症機序の解析と予防・治療法の開発に関わる体系的講義を行う．

Notice⟩ 特になし Schedule⟩

大項目担当
1. 主に下記の疾患や病態に関する研究への理解を深め独創性ある思考を養う
2. {骨粗鬆症，癌に伴う骨病変や多発性骨髄腫における骨病変の成立・進展機序の解明と治療法の開 \atop 発に関わる研究} 松本・井上・安倍

3. {心血管障害や脳血管障害の原因となる動脈硬化症の発症・進展機序の解明およびその予防・治療 \atop 法の開発に関わる研究} 東・赤池

4. {糖尿病，高脂血症などの代謝疾患の発症機序，疾患関連遺伝子などの解明および治療法の開発に \atop 関わる研究} 松本・東・井上

Evaluation Criteria⟩ 講義，実習に取り組む姿勢，理解度，実技内容により判断する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217592

Bioregulatory Sciences and Medicine 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toshio Matsumoto⟨Manager⟩ · Professor / Molecular Medicine, Proteomics

, Hiroyuki Azuma ·

Masahiro Abe · Associate Professor / Molecular Medicine, Proteomics , Daisuke Inoue ·

, Masashi Akaike · Professor / Institute of Health Biosciences

Target⟩ <GIO> 加齢に伴い発症あるいは増悪する多因子病の病態・発症機序の解明および予防・治療法の開発に関わる研究手法を理解し，遂行能力を養う． <SBO> 1．骨粗鬆症の発症機構の解明および予防・治療法の開発に関わる研究手法を理解する． 2．癌転移や多発性骨髄腫の発症・進展機構およびその解明に向けた研究手法を理解する． 3．動脈硬化の発症・進展機構およびその予防・治療法の開発に関わる研究手法を理解する． 4．糖尿病・高脂血症の発症機序およびその治療法の開発に関わる研究手法を理解する．

Outline⟩ 内分泌・代謝系，循環器系の多因子病や多発性骨髄腫などの血液・腫瘍疾患について，その背景となる遺伝子多型や変異，染色体転座，蛋白修飾などによりもたらされるプロテオーム系の変化による生体制御の異常に視点をおき，病態および発症機序の解析と予防・治療法の開発に関する輪読会または抄読会形式での演習を行う．

Notice⟩ 特になし Schedule⟩

大項目担当
1. {骨粗鬆症，癌に伴う骨病変や多発性骨髄腫における骨病変の成立・進展機序の解明と治療法の開 \atop 発に関わる研究の遂行に必要な演習を行う} 松本，井上，安倍

2. {心血管障害や脳血管障害の原因となる動脈硬化症の発症・進展機序の解明およびその予防・治療 \atop 法の開発に関わる研究の遂行に必要な演習を行う} 東，赤池

3. {糖尿病・高脂血症などの代謝疾患の発症機序，疾患関連遺伝子などの解明および治療法の開発に \atop 関わる研究の遂行に必要な演習を行う} 松本，東，井上

Evaluation Criteria⟩ 講義，実習に取り組む姿勢，理解度，実技内容により判断する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217593

Clinical Oncology 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science Masaki Hanibuchi · Associate Professor / University Hospital Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217554 Contact⟩

⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp) ⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp) ⇒ Hanibuchi . ⇒ Goto .

Practice of Clinical Oncology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Saburo Sone · Professor / Course of Bioregulation and Medical Oncology, Medical Science

Yasuhiko Nishioka · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Masaki Hanibuchi · Associate Professor / University Hospital

Hisatsugu Goto · Associate Professor / Course of Bioregulation and Medical Oncology, Medical Science

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217555 Contact⟩

⇒ Sone (+81-88-633-7126, ssone@clin.med.tokushima-u.ac.jp) ⇒ Nishioka (+81-88-633-7127, yasuhiko@clin.med.tokushima-u.ac.jp) ⇒ Hanibuchi . ⇒ Goto .

Stress Science 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Toshihito Tanahashi · Associate Professor / Molecular Medicine, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217516 Contact⟩

⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Tanahashi (+81-88-633-9004, tana@basic.med.tokushima-u.ac.jp)

Stress Science Seminar 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

, Toshihito Tanahashi · Associate Professor / Molecular Medicine, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217517 Contact⟩

⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Tanahashi (+81-88-633-9004, tana@basic.med.tokushima-u.ac.jp)

Molecular Medicine Seminar 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Takuya Sasaki⟨Manager⟩ · Professor / Molecular Medicine, Proteomics

, Akio Adachi · Professor / Course of Microbiology and Immunology, Medical Science

Yutaka Nakahori ·, Toshio Matsumoto · Professor / Molecular Medicine, Proteomics , Kazuhito Rokutan · Professor / Molecular Medicine, Proteomics

Tomomi Kuwahara · Associate Professor / Course of Microbiology and Immunology, Medical Science Tsuneo Uchiyama · Associate Professor / Course of Microbiology and Immunology, Medical Science

Target⟩ ポストゲノム時代の医科学研究の中心となるプロテオミクス研究の基本理論と実験技術・戦術を，具

体的な研究テーマのもとに指導する． Outline⟩ 生体はその生命活動を維持するため，たえず分子間ネットワークに基づく生体制御系によって恒常

性の維持を図っており，疾病はこの制御系の破綻の結果である．プロテオミクス研究を基盤とした分子間ネットワークの研究から疾病の機序解明，さらにはその予防・治療法等の開発を行うことが生体制御医学講座の目標である．この目標達成に必須な種々の技法，解析法に習熟して研究遂行能力を高めるよう本実験実習で学生を指導する．また，実験成績に基づいた作業仮説の立案と実証，修正のプロセスを学ばせ，最終的に研究成果を学術論文にまとめる方法を指導する．

Notice⟩ 特になし Schedule⟩

大項目

1. {各分野においてリサーチカンファレンス，抄読会を行うと共に，研究指導，技術指導，さらには，研究成果の発表の仕方，論 \atop 文作成指導を随時に行う．}

Evaluation Criteria⟩ 受講状況と演習内容の理解の程度に応じて評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217590 Contact⟩ ⇒ Sasaki (+81-88-633-9223, sasaki@basic.med.tokushima-u.ac.jp) ⇒ Adachi (+81-88-633-7078, adachi@basic.med.tokushima-u.ac.jp) ⇒ Nakahori . ⇒ Matsumoto (+81-88-633-7119, toshimat@clin.med.tokushima-u.ac.jp) ⇒ Rokutan (+81-88-633-9007, rokutan@basic.med.tokushima-u.ac.jp) ⇒ Kuwahara (+81-88-633-9229, tomomi@basic.med.tokushima-u.ac.jp) ⇒ Uchiyama (+81-88-633-9232, uchiyama@basic.med.tokushima-u.ac.jp)

Note⟩ これからの医科学研究は，いかなる研究領域においても，プロテオミクス的な思考と実験技術を理解し，習得することなく，進めていくことはできない．したがって，将来何らかの形で医科学研究に携わることを希望する者は本実験実習を選択することが望ましい．

Proteomics in Enzymology 2 units (selection)

1st-year(1st semester), 2nd-year(1st semester), 3rd-year(1st semester)

Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 蛋白質と酵素の機能解析法を習得して，病気の原因や治療法の解明など医学応用できる実力を身につ

ける． Outline⟩ 代表的酵素の一つとして，蛋白質分解酵素と分子シャペロン蛋白質を取り上げ，酵素機能の解析法

と医学応用として，各種病態の解明，治療法への応用についての最新情報を学ぶ．酵素蛋白質の精製法，活性測定法，細胞内機能解析法などの講義を行う．

Notice⟩ 講義と演習を行う． e-learning は行っていない． Schedule⟩

大項目担当
1. 蛋白質・酵素機能の解析法 1  木戸博
2. 蛋白質・酵素機能の解析法 2  〃
3. 蛋白質・酵素機能の解析法 3  〃
4. 蛋白質・酵素機能の解析法 4  〃
5. 蛋白質・酵素機能の解析法 5  〃
6. 蛋白質・酵素機能の解析法 6  〃
7. インフルエンザウイルス感染，インフルエンザ脳症と酵素 1  〃
8. インフルエンザウイルス感染，インフルエンザ脳症と酵素 2  〃
9. インフルエンザウイルス感染，インフルエンザ脳症と酵素 3  〃
10. インフルエンザウイルス感染，インフルエンザ脳症と酵素 4  〃
11. 分子シャペロンとシャペロン病 1  矢野仁康
12. 分子シャペロンとシャペロン病 2  〃
13. 分子シャペロンとシャペロン病 3  〃
14. 分子シャペロンとシャペロン病 4  〃
15. 分子シャペロンとシャペロン病 5  〃

Evaluation Criteria⟩ 出席とレポート提出 (出席は 2/3 以上をもって合格とする． ) Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217573 Contact⟩

⇒ Kido (+81-88-633-7423, kido@ier.tokushima-u.ac.jp) (Office Hour: e-mail により時間調節を適宜おこ

ないます． ) ⇒ Yano (yano@ier.tokushima-u.ac.jp) (Office Hour: e-mail により時間調節を適宜おこないます． )

Practice in Proteomics Enzymology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hiroshi Kido⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Suehiro Sakaguchi · Professor / Course of Medicine for Biological Responses, Proteomics Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Outline⟩ プロテオミクス解析による新しい蛋白質機能の解析法と医学応用について演習を行う．さらにプロ

テオミクの解析技術についても解説する． Notice⟩ e-learning は行っていない．

Schedule⟩

大項目内容
1. 論文抄読会第 2， 3 週月曜日と金曜日 9:00∼ 10:00
2. リサーチカンファレンス第 1， 4 週月曜日 9:00∼ 12:00
3. 研究指導，技術指導随時

Evaluation Criteria⟩ 出席と抄読参加 (出席は 2/3 以上をもって合格とする． ) Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217574 Contact⟩

⇒ Kido (+81-88-633-7423, kido@ier.tokushima-u.ac.jp) (Office Hour: e-mail により時間調節を適宜おこ

ないます． ) ⇒ 他の教員についても， e-mail にて時間調節の上，面談して下さい．

Proteomics 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hisaaki Taniguchi⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics , Kazuko Fujiwara ·

, Hiroaki Konishi ·

Target⟩ 遺伝子産物である蛋白質がどのように相互作用して生命活動を制御するのかを理解する手段として，

プロテオミクス解析学の理論的基礎を習得する． Outline⟩ ゲノム情報を基に，多数の微量蛋白質の構造を網羅的に解析する方法は，現在急速に進歩している．

そこで，プロテオミクス解析の基盤となる方法論を中心に，微量蛋白質の網羅的分離，高感度質量分析計を駆使した一次構造解析，ゲノム情報と得られた蛋白質情報の総合的処理法について体系的に学習する．

Notice⟩ 印刷したレジュメは配布するが，パワーポイントなどのコンピュターによるプレゼンテーションが多

いので，ノートの取り方などに工夫する必要がある． Schedule⟩

大項目内容
1. リサーチカンファレンス各週 1 回
2. 抄読会〃
3. 研究指導・技術指導随時
4. 方法論の開発による将来的発展性についても学ぶ．

Evaluation Criteria⟩ 出席状況，授業内容への質疑応答などの受講態度など． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217569 Contact⟩

⇒ Taniguchi (+81-88-633-7426, hisatan@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 18:00 (e-

mail により時間調節を適宜おこないます)) ⇒ 他の教員についても， e-mail にて時間調節の上，面談して下さい．

Proteomics seminar 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Hisaaki Taniguchi⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics , Kazuko Fujiwara · , Hiroaki Konishi · Target⟩ 生体におけるシグナル伝達，蛋白質―蛋白質相互作用を研究するうえで必要なプロテオミクス解析学

の手法を実際に修得する． Outline⟩ 典型的な蛋白質の数種類を実際例として，二次元電気泳動，高速クロマトグラフィー，数種の質量分

析計を使用した構造解析，得られているゲノム情報と，新たに得られた蛋白質情報を総合的に解析する手法に関し，研究計画の立案の仕方およびその実証法について統括的に指導する．

Notice⟩ パワーポイントなどのプレゼンテーションを用いた講義，発表などを行うため，自ら習認することが

望ましい． Schedule⟩

大項目内容
1. リサーチカンファレンス各週 1 回
2. 抄読会〃
3. 研究指導・技術指導随時
4. {病因解明におけるプロテオミクス解析学の応用に \atop ついても言及し，臨床医との交流を活発に行なう．}

Evaluation Criteria⟩ 出席状況，授業内容への質疑応答などの受講態度など． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217570 Contact⟩

⇒ Taniguchi (+81-88-633-7426, hisatan@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 18:00 (e-

mail により時間調節を適宜おこないます)) ⇒ 他の教員についても， e-mail にて時間調節の上，面談して下さい．

Molecular immunology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuru Matsumoto⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 生体防御機構としての免疫システムの概念を学ぶ． Outline⟩ 生体防御機構としての免疫系は，リンパ球や樹状細胞などの免疫担当細胞による異物の認識とその後

の排除機構によって狙われている．また，これらの免疫担当細胞が効率よく機能するためには，リンパ節や膵臓といったリンパ組織が，免疫担当細胞に対して免疫応答の場を与えることも重要な要素である．こうした事実をふまえ，免疫系を高度に統御された生体システムとして理解するためには，免疫システムを構築する血球細胞やリンパ組織の発生や分化についての理解が必要である．

Notice⟩ 受講に当たっては事前に担当教官に電話等で確認すること．講義の一部については e-learning 化も

検討しているが，利用可能になった場合には e-learning 学習も出席として取り扱う． Schedule⟩

大項目内容担当
1. 論文抄読会毎週火曜日 9:00∼ 11:00
2. リサーチカンファレス毎週土曜日 9:00∼ 11:00
3. 研究指導・技術指導随時
4. 授業ガイダンス松本
5∼8. 免疫ゲノム序論〃
9∼13. 免疫ゲノム理論〃

14∼18.免疫ゲノム応用〃

Evaluation Criteria⟩ 受講状況と授業内容の理解の程度に応じて評価する．再試験は行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217581 Contact⟩

⇒ Matsumoto (Room 701, +81-88-633-7432, mitsuru@ier.tokushima-u.ac.jp) (Office Hour: 火曜日の16: 00∼ 18:00 (e-mail により時間調節を適宜，行います))

Practice for molecular immunology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuru Matsumoto⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ ゲノム情報に基づく生体防御機構としての免疫学研究方法を学ぶ． Outline⟩ 生体防御機能としての免疫系は，強力な病原体の侵入や原因不明の機能破綻により，種々の疾患を

もたらす．このような病態を明らかにするためには，ゲノム情報に基づいた最新のバイオ技術を駆使した医科学研究が必要である．

Notice⟩ 受講に当たっては事前に担当教官に電話等で確認すること．講義の一部については e-learning 化も

検討しているが，利用可能になった場合には e-learning 学習も出席として取り扱う． Schedule⟩

大項目内容担当
1. 論文抄読会毎週火曜日 9:00∼ 11:00
2. リサーチカンファレス毎週土曜日 9:00∼ 11:00
3. 研究指導・技術指導随時
4. 授業ガイダンス松本
5∼8. 免疫ゲノム機能序論〃
9∼13. 免疫ゲノム機能理論〃

14∼18.免疫ゲノム機能応用〃

Evaluation Criteria⟩ 受講状況と授業内容の理解の程度に応じて評価する．再試験は行わない． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217582 Contact⟩

⇒ Matsumoto (Room 701, +81-88-633-7432, mitsuru@ier.tokushima-u.ac.jp) (Office Hour: 火曜日の16: 00∼ 18:00 (e-mail により時間調節を適宜，行います))

Systems Biology for Molecular Signailng 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Suehiro Sakaguchi⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Keiji Uchiyama · Associate Professor / Course of Medicine for Biological Responses, Proteomics Tsuyoshi Mori · Assistant Professor / Institute for Enzyme Research

Target⟩ 生体における情報の統御システムの基礎知識を習得し，その統御システムの破綻と様々な疾病の病態

生理について学習する． Outline⟩ 分子間における情報伝達の仕組みについて体系的に講義するとともに，その仕組みの破綻が如何に

病気を招くのか，神経変性疾患を例にとりながら講義する． Notice⟩ なし Schedule⟩

大項目内容
1. リサーチカンファランス毎週水曜日 10:00∼ 12:00
2. 抄読会〃
3. 研究指導・技術指導随時

Evaluation Criteria⟩ 出席状況，実習態度，レポートなどにより評価する． Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217597 Contact⟩

⇒ Sakaguchi (sakaguch@ier.tokushima-u.ac.jp) ⇒ Uchiyama (+81-88-633-7448, ku200@ier.tokushima-u.ac.jp) ⇒ Mori (morit@ier.tokushima-u.ac.jp)

Practice on Systems Biology for Molecular Signailng 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Suehiro Sakaguchi⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Keiji Uchiyama · Associate Professor / Course of Medicine for Biological Responses, Proteomics Tsuyoshi Mori · Assistant Professor / Institute for Enzyme Research

Target⟩ 生体における情報の統御システム，及びその統御システムの破綻と様々な疾病の病態生理についての

基礎知識を習得する． Outline⟩ 分子間における情報伝達の仕組み，さらにその仕組みの破綻が如何に病気を招くのかにつてのいく

つかの代表的な学術論文を抄読する．また，これらの論文の抄読を通して，作業仮説や研究計画の立案及び実証の方法について統括的に指導する．

Notice⟩ なし Schedule⟩

大項目内容
1. 抄読会毎週水曜日 10:00∼ 12:00
2. 研究指導・技術指導随時

Evaluation Criteria⟩ 出席状況，実習態度，レポートなどにより評価する． Re-evaluation⟩ なし Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217598 Contact⟩

⇒ Sakaguchi (sakaguch@ier.tokushima-u.ac.jp) ⇒ Uchiyama (+81-88-633-7448, ku200@ier.tokushima-u.ac.jp) ⇒ Mori (morit@ier.tokushima-u.ac.jp)

Proteome Genetics 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yousuke Ebina⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Tomoyuki Yuasa · Associate Professor / Course of Medicine for Biological Responses, Proteomics Keiji Uchiyama · Associate Professor / Course of Medicine for Biological Responses, Proteomics Hisao Nagaya · Assistant Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 糖尿病を例に取り，プロテオミクス遺伝学を修得し，先端的研究を理解する． Outline⟩ インスリンホルモンを中心にプロテオミクス解析し，分子間相互作用の解析技術を用いた細胞内シ

グナル伝達のメカニズムの解析と，その破綻によって起こる糖尿病について最新の知見を含めて体系的に解決するとともに，発生工学を用いた糖尿病モデルマウスの構築とその解析についても講義する．

Notice⟩ e-learning 学習は扱っていない

Schedule⟩

大項目中項目担当
1. 授業ガイダンス+プロテオミクス遺伝学序論蛯名
2. 遺伝子工学論 1  遺伝子工学総論〃
3. 遺伝子工学論 2  遺伝子工学基礎技術論〃
4. 遺伝子工学論 3  遺伝子工学特論〃
5. 細胞生物学論 1  細胞内情報伝達総論〃
6. 細胞生物学論 2  インスリン情報伝達論〃
7. 細胞生物学論 3  インスリン情報伝達とインスリン作用〃
8. 分子生物学論 1  分子生物学総論〃
9. 分子生物学論 2  分子生物学技術論〃

10. 分子生物学論 3  分子生物学特論〃
11. プロテオミクス分子遺伝学解析法 1  〃
12. プロテオミクス分子遺伝学解析法 2  〃
13. プロテオミクス分子遺伝学解析法 3  〃
14. パイオインフォマティクス総論〃
15. バイオインフォマティクス各論〃

Evaluation Criteria⟩ 出席状況・実習態度・レポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217571 Contact⟩

⇒ Ebina (+81-88-633-7436, iden@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 17:30 (e-mailに

より時間調節を行います． )) ⇒ 他の教員についても,e-mailにて時間調節の上,面談してください.

Note⟩ 視聴覚設備を最大限利用し，わかりやすい講義を行う

Laboratory Practice of Proteome Genetics 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yousuke Ebina⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Keiji Uchiyama · Associate Professor / Course of Medicine for Biological Responses, Proteomics Tomoyuki Yuasa · Associate Professor / Course of Medicine for Biological Responses, Proteomics Hisao Nagaya · Assistant Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 糖尿病を例に取り，プロテオミクス遺伝学を修得し，先端的研究を推進する． Outline⟩ インスリンホルモンを中心に細胞内シグナル伝達のメカニズムの解析や，その破綻によって起こる

糖尿病について，特に必要ないくつかのこの分野での代表的学術論文を抄読させ，仮説や研究計画の立案の仕方や，その実証方法について統括的に指導する．

Notice⟩ e-learning 学習は扱っていない

Schedule⟩

大項目担当
1. 授業ガイダンス+プロテオミクス遺伝学序論蛯名
2. 遺伝子工学基礎実習 1  〃
3. 遺伝子工学基礎実習 2  〃
4. 遺伝子工学基礎実習 3  〃
5. 細胞生物学演習・講読 1  〃
6. 細胞生物学演習・講読 2  〃
7. 細胞生物学演習・講読 3  〃
8. ゲノムシークエンス実習〃
9. 組換え DNA 実習 1  〃
10. 組換え DNA 実習 2  〃
11. プロテオミクス分子遺伝学解析実習 1  〃
12. プロテオミクス分子遺伝学解析実習 2  〃
13. プロテオミクス分子遺伝学解析実習 3  〃
14. パイオインフォマティクス総論演習〃
15. バイオインフォマティクス各論演習〃

Evaluation Criteria⟩ 出席状況・実習態度・レポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217572 Contact⟩

⇒ Ebina (+81-88-633-7436, iden@ier.tokushima-u.ac.jp) (Office Hour: 水曜日の16:00∼ 17:30) Note⟩ 視聴覚設備を最大限利用し，わかりやすい講義を行う

Research Progress in Enzyme Pathophysiology 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics Takashi Sakai · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217538

Contact⟩ ⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 8:30 to 17:30) ⇒ Sakai (405 号室, +81-88-633-7430, sakai@ier.tokushima-u.ac.jp)

Selected Topics on Enzyme Pathophysiology 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics Takashi Sakai · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217539

Contact⟩ ⇒ Fukui (401, +81-88-633-7429, kiyo@ier.tokushima-u.ac.jp) (Office Hour: 8:30 to 17:30) ⇒ Sakai (405 号室, +81-88-633-7430, sakai@ier.tokushima-u.ac.jp)

Course on Genomic Information 2 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuo Itakura⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Hiroshi Inoue · Designated Professor / Institute for Genome Research

Target⟩ ヒトゲノムの多様性がどのように疾患感受性や薬剤応答性を決めるか，また，これらに関わる遺伝子の多型をヒトだけでなくマウス等のモデル生物を用いてどのように探索するかについて理解させる．

Outline⟩ ヒトゲノムの多様性は，蛋白質の多様性を介して，個人の差，病気の罹り易さ，薬剤に対する応答性等の体質を規定している．ヒトゲノム多様性マーカーを用いることにより，疾患関連遺伝子を同定するための遺伝統計学的を用いる筋道と，標的遺伝子の調節機能を遺伝子改変動物を用いてどのように解析するかについて体系的に講義を行う． Notice⟩ e-learning は取り扱っていない

Schedule⟩

大項目内容担当
1. リサーチカンファレンス毎週金曜日 17:00∼ 18:00  板倉，井上
2. 抄読会毎週月曜日 17:00∼ 18:00  〃
3. 研究指導・技術指導随時〃

Evaluation Criteria⟩ 日頃の実験に取り組む姿勢，出席率，レポート等により判断する．

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217560

Contact⟩

⇒ e-mailにて時間調節の上,面談してください. Note⟩ 特になし

Practice on Genomic Information 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Mitsuo Itakura⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics Hiroshi Inoue · Designated Professor / Institute for Genome Research Target⟩ ゲノムの一次配列情報のみならず，それぞれの個人毎に異なる多型の意味を理解させる． Outline⟩ 遺伝子多型と疾患関連遺伝子の同定を中心に，実際の代表的研究例を基に，学術論文を学生に抄読

させ，仮説や研究計画の立案の仕方およびその実証方法について統括的に指導する．

ゲノム機能研究センター遺伝情報分野に設置された機器を用いて指導する． Notice⟩ e-learning は取り扱っていない

Schedule⟩

大項目内容担当
1. リサーチカンファレンス毎週金曜日 17:00∼ 18:00  板倉，井上
2. 抄読会毎週月曜日 17:00∼ 18:00  〃
3. 研究指導・技術指導随時〃

Evaluation Criteria⟩ 日頃の実験に取り組む姿勢，出席率，レポート等により判断する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217561 Contact⟩

⇒ e-mailにて時間調節の上,面談してください. Note⟩ 特になし

Genome Medicine 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217490 Contact⟩

⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp)

Practice on Genome Medicine 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Toyomasa Katagiri · Professor / Course of Medicine for Biological Responses, Proteomics

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217491 Contact⟩

⇒ Katagiri (tkatagi@genome.tokushima-u.ac.jp)

Experimental practice of medicine for biological responses 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Yousuke Ebina⟨Manager⟩ · Professor / Course of Medicine for Biological Responses, Proteomics

Hiroshi Kido · Professor / Course of Medicine for Biological Responses, Proteomics , Hisaaki Taniguchi · Professor / Course of Medicine for Biological Responses, Proteomics Mitsuru Matsumoto · Professor / Course of Medicine for Biological Responses, Proteomics Suehiro Sakaguchi · Professor / Course of Medicine for Biological Responses, Proteomics

Kiyoshi Fukui · Professor / Course of Medicine for Biological Responses, Proteomics

, Mitsuo Itakura · Professor / Course of Medicine for Biological Responses, Proteomics

Eiji Hara ·, Akira Tanigami · Part-time Lecturer / Graduate School of Medical Sciences

Eiichi Takahashi · Part-time Lecturer / Graduate School of Medical Sciences

Mihiro Yano · Associate Professor / Course of Medicine for Biological Responses, Proteomics , Kazuko Fujiwara ·

, Isao Inoue ·

Takashi Sakai · Associate Professor / Course of Medicine for Biological Responses, Proteomics

Target⟩ 生体応答医学・生体対応医学大講座を構成する各分野の実験実習を通して，実験主義・技能・実験研

究法・論文作成法を習得する． Outline⟩ えられたゲノム情報を基にプロテオミクス研究を展開し，その結果を医学的応用するために必要な

基本的実験技能を習得させる．さらに，実験結果を基に目的を達成するための作業仮説の立案・実証・修正プロセスを学ばせ，最終的に研究成果を学術論文としてまとめる方法を指導する．具体的には，生命現象の動的バランスを支える生体制御の仕組みをプロテオミクスの観点から解説する実験実習と，この動的バランスの破綻に基づく病気の発症原因と治療法の解明のためのプロテオミクス研究の実験実習を行い，両研究を統括的に理解しえる能力を養う．

Notice⟩ e-learning 学習は扱っていない

Schedule⟩

大項目

1. 各研究室にて随時，技術指導をする．

Evaluation Criteria⟩ 出席状況・実習態度・レポート等により評価する． Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217591 Contact⟩

⇒ e-mailにて各教員と時間調節の上，面談してください． Note⟩ 生体応答医学・生体対応医学大講座の複数の教員がそれぞれの専門分野の実験実習を指導する．

ゲノム機能解析学 2 units (selection) 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217492

ゲノム機能解析学演習 4 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217493

生体対応医学実験実習 12 units (selection)

1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=217522

Graduate School of Medical Sciences (2011) ⟩ Course of Medicine (Doctor’s Degree)

Master Course of Medical Sciences (Master’s Degree)

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ............................................................................................................3 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) .......................................................................................................................... 4 Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka ・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) .......................................... 4 Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・ Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester) ......................................................5 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao ・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ........................................................ 5 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・ Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga ..........................................6

• 各専攻系間共通カリキュラム科目

Human Science (Basic human science from structure to function) … Rokutan ・ Kawai ・ Masuda / each-year(1st semester) ............................................................................................ 6

Introduction to Clinical Pharmacology … Tamaki ・ Takiguchi ・ Kiwada ・ Tsuchiya ・ Yamauchi ・ Tomita / each-year(2nd semester)............................................................................................7 Introduction to Proteomics … Taniguchi ・ Kido ・ Ebina ・ Noma ・ Miyamoto / each-year(1st semester) ..................................................................................................................... 7 Genomic Drug Discovery … Itou ・ Itakura ・ Tamaki ・ Sano ・ Part-time Lecturer / each-year(1st semester)...................................................................................................................8

• 専門科目

Human Science(Pathology) … Sakashita ・ Izumi ・ Kuwahara ・ Adachi ・ Yasutomo / 1st-year(1st semester, 2nd semester) ......................................................................................................8 General Medicine … Kaji ・ Ito ・ Soeki ・ Doi ・ Bando ・ Nakajo ・ Kawano ・ Takeda ・ Saito .................................................................................................................................9 Summary of Medical Genomic … Ebina ・ Itakura ・ Fukui / each-year(1st semester) ....................................................................................................................................9 Basic practice for medical science … Matsumoto ・ Fukui ・ Kido ・ Yano ・ Ishimura ・ Fukui ・ Sano ・ Izumi ・ Sasaki ・ Ebina ・ Yasutomo ・ Tamaki ・ Morita ・ Yoshizaki ・ Taniguchi ・ Itakura ・ Matsumoto ・ Takahama ・ Adachi / 1st-year(1st semester)9 Proteomics in Medicine … Siomi ・ Sakaguchi / 1st-year(1st semester) ............................................................................................................................................... 10

Applied Molecular Enzymology and Pathogenesis … Kido ・ Yano / 1st-year(1st semester) ..........................................................................................................................10

Medical Electronics … Yoshizaki ・ Sei ・ Moriguchi ・ Kinouchi / each-year(1st semester) .................................................................................................................................10 Medical Application of Nanotechnology … Doi ・ Miyoshi ・ Ishida ・ Ishikawa / 1st-year(1st semester)...................................................................................................................11 Clinical Genetics … Nakahori ・ Sato / 1st-year(1st semester) ........................................................................................................................................................ 11 Development, Differentiation, and Regeneration … Takahama ・ Matsumoto ・ Noji ・上野・ Hibi / each-year(1st semester) .................................................................................................11 Project Search Study Course … Fukui ・ Sone ・ Ishimura ・ Ebina ・ Yoshizaki ・ Itakura ・ Tamaki ・ Siomi ・ Fukui ・ Kido ・ Matsumoto ・ Taniguchi ・ Sasaki ・ Takahama ・ Sano ・ Izumi ・ Yasutomo ・ Adachi ・ Nakahori ・ Kubo ・ Moriguchi ・ Matsumoto ・

Doi ・ Nagahiro ・ Morita ・ Yano ・ Matsumoto ・ Tokunaga ・ Kagami ・ Rokutan ・ Ohmori ・ Takeda ・ Sei / 1st-year(whole year), 2nd-year(whole year) .................................................................................12

メディカルサイエンス特別研究 … 修士課程全教員 / 2nd-year .....................................................................................................................................................12

Course of Medicine (Doctor’s Degree)

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ...........................................................................................................17 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) ......................................................................................................................... 17 Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka ・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) .........................................18 Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・ Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester) .....................................................18 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao ・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ....................................................... 19 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・ Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga .........................................19

• 各専攻系間共通カリキュラム科目

Introduction to Clinical Medicine … Kaji ・ Ito ・ Soeki ・ Doi ・ Bando ・ Nakajo ・ Kawano ・ Takeda ・ Saito / each-year(1st semester) ...........................................................................................20 A Training of Microbiology and Immunology … Miyake ・ Adachi ・ Yasutomo ・ Ono ・ Kuwahara ・ Uchiyama ・ Maekawa ・ Hirota / 1st-year(2nd semester), 2nd-year(2nd semester) ................................................20

• 指定科目

Basic practice for physiological and pharmacological research … Tamaki ・ Yoshizaki ・ Sei ・ Sano ・ Minakuchi / 1st-year(1st semester) .................................................................................. 20 Introductory lab course for biochemical and molecular biology experiments … Takahama ・ Oyadomari / each-year(1st semester) ....................................................................................21

• 専門科目

Lecture for Birth Defects … Fukui ・ Sakata-Haga / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................................................21 Seminar for Birth Defects … Fukui ・ Sakata-Haga / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .....................................................................................................21 Developmental Pathophysiology … Kagami ・ Mori ・ Hayabuchi ・ Watanabe ・ Kondo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................... 22 Practical developmental pathophysiology … Kagami ・ Mori ・ Hayabuchi ・ Watanabe ・ Kondo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................22 Obstetrics and Gynecology … Irahara ・ Furumoto ・ Yasui ・ Matsuzaki ・ Maeda ・ Kuwahara ・ Yamamoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................... 22 Practice of Obstetrics and Gynecology … Irahara ・ Furumoto ・ Yasui ・ Matsuzaki ・ Maeda ・ Kuwahara ・ Yamamoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................. 22 発生発達医学実験実習 … Fukui ・ Kagami ・ Irahara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...................................................................................................... 23 Physiology … Yoshizaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................................................................................. 23 Practice of Physiology … Yoshizaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................................................23 Gastroenterology and Oncology … Takayama ・ Okahisa ・ Muguruma / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................... 23 Practice of Gastroenterology and Oncology … Takayama ・ Okahisa ・ Muguruma / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .........................................................................23 Immune system development … Takahama ・ Takada / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................... 24 Immune system development … Takahama ・ Takada / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................... 24 Immune Regulation … Okazaki ・ Okazaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................................. 24 Practice of Immune Regulation … Okazaki ・ Okazaki / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................. 24 Experimental practice for preventive medicine … Yasutomo ・ Yoshizaki ・ Takahama ・ Kishihara ・ Shimizu ・ Tomita ・ Honda / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).....................................24 Immunology … Yasutomo ・ Kishihara ・ Maekawa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .........................................................................................................24 Practice for Immunology … Yasutomo ・ Kishihara ・ Maekawa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................24 Microbiology … Adachi ・ Kido ・ Uchiyama / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..............................................................................................................25 Practice for Microbiology … Adachi ・ Uchiyama / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).......................................................................................................25 Practice for Microbiology and Immunology … Yasutomo ・ Adachi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................... 25 Preventive Medicine … Arisawa ・ Uemura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................................. 25 Practice of Preventive Medicine … Arisawa ・ Uemura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................. 26 Community Medicine … Tani / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................................... 26 Practice of Community Medicine … Tani / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................................... 26 Practice of Social and Environmental Medicine … Arisawa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................................... 26 Anatomy and Cell Biology … Ishimura ・ Nakamura / 1st-year(1st semester, 2nd semester), 2nd-year(1st semester, 2nd semester), 3rd-year(1st semester, 2nd semester) ...........................................................26 Practice of Anatomy and Cell Biology … Ishimura ・ Nakamura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................27 Integrative Physiology … Sei / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)........................................................................................................................27 Practice of Integrative Physiology … Sei ・ Shiuchi ・ Fujihara ・ Chikahisa / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................. 27 Psychiatry … Ohmori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................................................27 Practical Psychiatry … Ohmori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................................................................27 Brain Pathophysiology and Medicine … Nagahiro / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................... 27 Practice of Brain Pathophysiology and Medicine … Nagahiro / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................... 27 Strategic medical informatics … Moriguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................................................................................................... 27 Practice of strategic medical informatics … Moriguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...............................................................................................28 Molecular Biology … Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .....................................................................................................................28 Practice of Molecular Biology … Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .........................................................................................................28 Practice for Integrated Brain Sciences and Medical Information … Ishimura ・ Sei ・ Ohmori ・ Nagahiro ・ Moriguchi ・ Oyadomari / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................ 28 Pharmacology … Tamaki ・ Minakuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................................................28 Practice for Pharmacology … Tamaki ・ Minakuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .....................................................................................................28 Anesthesiology … Oshita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................................................................................29 Practice of Anesthesiology … Oshita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).................................................................................................................29 Nephrology … Doi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................................................................. 29 Seminar for Nephrology … Doi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .....................................................................................................................29 救急集中治療医学 … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).......................................................................................................................29 救急集中治療医学演習 … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................................................. 29 Experimental Practice for Biological Medicine … Tamaki ・ Oshita ・ Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........................................................................29 Ophthalmology … Mitamura ・ Eguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................................................29 Practice of Ophthalmology … Mitamura ・ Eguchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................................................29 Otorhinolaryngology … Takeda ・ Tamura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................................................................. 30 Experimental Practice for Otorhinolaryngology … Takeda ・ Tamura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...................................................................................30 Clinical Neuroscience … Kaji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................................... 30 Clinical Neuroscience Workshop … Kaji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................................ 30 Sensory Neuroscience Laboratory Experiment … Takeda ・ Shiota ・ Kaji ・ Tamura ・ Naito ・ Izumi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................30 Human Pathology … Sakashita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................................................................................................................31 Practice of Human Pathology … Sakashita / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...........................................................................................................31 Gastroenterological Surgery and Transplantation … Shimada ・ Utsunomiya ・ Kurita ・ Iwata ・ Nishioka ・ Morine ・ Imura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................... 31 Practice of Gastroenterological Surgery and Transplantation … Shimada ・ Utsunomiya ・ Kurita ・ Iwata ・ Nishioka ・ Morine ・ Imura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................... 31 Cardiovascular Surgery … Kitagawa ・ Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .........................................................................................................31 Practice of Cardiovascular Surgery … Kitagawa ・ Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).............................................................................................32 Urology … Kanayama ・ Fukumori ・ Takahashi ・ Nakatsuji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...................................................................................................32 Experimental Practice for Urology … Kanayama ・ Fukumori ・ Takahashi ・ Nakatsuji / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................... 32 Cardiovascular Medicine … Sata / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................................32 Cardiovascular Medicine Practice … Sata / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................................................................32 Laboratory Experiment Practice of Diseased Organ and Repair Medicine … Kitagawa ・ Sakashita ・ Shimada ・ Kanayama ・ Ishibashi ・ Kitaichi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...............33 Experimental Oncology … Izumi ・ Uehara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).............................................................................................................33 Practice for Experimental Oncology … Izumi ・ Uehara / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................... 33 Radiology … Harada ・ Otsuka ・ Ikushima ・ Ueno ・ Matsuzaki ・ Takeuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...................................................................................... 33

Graduate School of Medical Sciences (2011) ⟩ Course of Proteomics Medical Sciences (Doctor’s Degree)

Practice of Radiology … Harada ・ Ikushima ・ Otsuka ・ Matsuzaki ・ Takeuchi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).................................................................................33 Respiratory Medicine and Rheumatology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)..........................................................................34 Practice of Respiratory Medicine and Rheumatology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................. 34 Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................. 34 Practice of Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................... 34 Thoracic, endocrine surgery and Oncology … Tangoku ・ Sakiyama ・ Kondo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................................34 Surgical Oncology … Tangoku / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................................... 35 Experimental Practice for Bioregulation and Medical Oncology … Izumi ・ Nishitani ・ Sone ・ Tangoku / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................... 35 Reconstructional Pathology … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .............................................................................................................35 Practice of Reconstructional Pathology … Nishimura / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).................................................................................................35 Dermatology … Kubo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ............................................................................................................................... 35 Practice for Dermatology … Kubo / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................................................. 35 Orthopedic Surgery … Yasui ・ Katoh ・ Sairyo ・ Takahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year), 4th-year ........................................................................................35 Orthopedic Surgery Practice … Yasui / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...............................................................................................................35 Plastic Surgery and Reconstructive Surgery … Nakanishi ・ Hashimoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................36 Practice for Plastic Surgery and Reconstructive Surgery … Nakanishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...............................................................................36 Laboratory Experiment of Pathologic Medicine for Cutaneo-motoric System … Nishimura ・ Yasui ・ Nakanishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..........................................36

Course of Proteomics Medical Sciences (Doctor’s Degree)

• 全専攻系共通カリキュラム科目

Introduction to Biological Ethics … Katagiri ・ Kitamura ・ Terao ・ Nakajo ・ Minakuchi / each-year(1st semester) ...........................................................................................................38 Clinical Psychology … Sato ・ YAMAMOTO ・ Sakai ・ Fukumori ・ Mitome / 1st-year(1st semester) ......................................................................................................................... 38 Introduction to social medicine, epidemiology and biostatistics … Arisawa ・ Ito ・ Uemura ・ Yamauchi ・ Tanioka ・ Tokumura ・ Hinode ・ Moriguchi ・ Yanagawa / each-year(2nd semester) .........................................38 Introduction to How to Write Up Scientific Manuscript in English … Tomita ・ Kaji ・ Fukui ・ Nakaya ・ Okazaki ・ Nagashino ・ Fukui ・ Kalubi / each-year(2nd semester) .....................................................39 Psychosomatic health and environmental stress … Takeda ・ Ohmori ・ Takahama ・ Sei ・ Rokutan ・ Oyadomari ・ Terao ・ Miyamoto ・ Nakaya ・ Yamanishi ・ Sakaue ・ Hosaka ....................................................... 39 Basic methods in life sciences … Fukui ・ Mizuguchi ・ Nakaya ・ Haneji ・ Fukui ・ Seta ・ Itou ・ Matsumoto ・ Kataoka ・ Yamazaki ・ Okazaki ・ Nikawa ・ Yamamoto ・ Katagiri ・ Nagata ・ Yoshinaga .........................................40

• 各専攻系間共通カリキュラム科目

A Training of Microbiology and Immunology … Miyake ・ Adachi ・ Yasutomo ・ Ono ・ Kuwahara ・ Uchiyama ・ Maekawa ・ Hirota / 1st-year(2nd semester), 2nd-year(2nd semester) ................................................41

• 指定科目

Introductory lab course for protein experiments … Takahama ・ Oyadomari ・ Okazaki ・ Katagiri / each-year(1st semester) ............................................................................................... 41 Introductory practice of proteomic analysis … Sakaguchi ・ Taniguchi ・ Ebina ・ Uchiyama ・ Yuasa / 1st-year(2nd semester)................................................................................................41 Introductory practice for functional genomics … Matsumoto ・ Itakura ・ Kuwahara / 1st-year(1st semester) .............................................................................................................42 Application of proteomics to medicine … Matsumoto ・ Adachi ・ Nakahori ・ Rokutan / 1st-year(2nd semester)............................................................................................................42

• 専門科目

Graduate School of Medical Sciences (2011) ⟩ Course of Proteomics Medical Sciences (Doctor’s Degree)

Disease Proteomics … Sasaki ・ Sakane / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................................ 43 Disease Proteomics Seminar … Sasaki ・ Sakane / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................... 43 Human Genetics … Imoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...........................................................................................................................43 Human Genetics seminor … Imoto ・ Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................................................................................................44 Bioregulatory Sciences and Medicine … Matsumoto ・ Azuma ・ Abe ・ Inoue ・ Akaike / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................................................................44 Bioregulatory Sciences and Medicine … Matsumoto ・ Azuma ・ Abe ・ Inoue ・ Akaike / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .......................................................................44 Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................. 44 Practice of Clinical Oncology … Sone ・ Nishioka ・ Hanibuchi ・ Goto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................... 45 Stress Science … Rokutan ・ Tanahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................................................................45 Stress Science Seminar … Rokutan ・ Tanahashi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................................ 45 Molecular Medicine Seminar … Sasaki ・ Adachi ・ Nakahori ・ Matsumoto ・ Rokutan ・ Kuwahara ・ Uchiyama / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................... 45 Proteomics in Enzymology … Kido ・ Yano / 1st-year(1st semester), 2nd-year(1st semester), 3rd-year(1st semester) ........................................................................................................ 45 Practice in Proteomics Enzymology … Kido ・ Sakaguchi ・ Yano / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................ 46 Proteomics … Taniguchi ・ Fujiwara ・ Konishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).............................................................................................................46 Proteomics seminar … Taniguchi ・ Fujiwara ・ Konishi / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ....................................................................................................46 Molecular immunology … Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................................ 46 Practice for molecular immunology … Matsumoto / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................... 47 Systems Biology for Molecular Signailng … Sakaguchi ・ Uchiyama ・ Mori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...............................................................................47 Practice on Systems Biology for Molecular Signailng … Sakaguchi ・ Uchiyama ・ Mori / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ..................................................................47 Proteome Genetics … Ebina ・ Yuasa ・ Uchiyama ・ Nagaya / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................ 47 Laboratory Practice of Proteome Genetics … Ebina ・ Uchiyama ・ Yuasa ・ Nagaya / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................48 Research Progress in Enzyme Pathophysiology … Fukui ・ Sakai / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ......................................................................................48 Selected Topics on Enzyme Pathophysiology … Fukui ・ Sakai / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year).........................................................................................48 Course on Genomic Information … Itakura ・ Inoue / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ................................................................................................... 48 Practice on Genomic Information … Itakura ・ Inoue / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) .................................................................................................. 48 Genome Medicine … Katagiri / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................................................49 Practice on Genome Medicine … Katagiri / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ........................................................................................................... 49 Experimental practice of medicine for biological responses … Ebina ・ Kido ・ Taniguchi ・ Matsumoto ・ Sakaguchi ・ Fukui ・ Itakura ・ Hara ・ Tanigami ・ Takahashi ・ Yano ・ Fujiwara ・ Inoue ・ Sakai / 1st-year(whole year), 2nd-year(whole year),

3rd-year(whole year).........................................................................................................................................................................................49 ゲノム機能解析学 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year) ...............................................................................................................................49 ゲノム機能解析学演習 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...........................................................................................................................49 生体対応医学実験実習 … / 1st-year(whole year), 2nd-year(whole year), 3rd-year(whole year)...........................................................................................................................49